A Bibliography of Publications in Interval Computations and Reliable Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

and

R. Baker Kearfott University of Louisiana at Lafayette

Lafayette, LA 70504 USA and

Vladik Kreinovich University of Texas at El Paso Department of Computer Science

Chemistry & Computer Science Building Room 3.1002

500 West University Avenue

El Paso, TX 79968-0518

USA

E-mail: vladik@utep.edu
WWW URL: http://www.cs.utep.edu/vladik/

14 October 2017 Version 0.03

Title word cross-reference

2 [149]. 3 [149, 706]. [x] = [A][x] + [b] [604, 683]. ϵ^{2} [531]. ϵ^{2−δ} [531]. G [694]. H [594]. H_{∞} [309]. I(R), I(R^{n}) [86]. L^{2} [352]. l^{p} [713]. M [46, 130, 434, 518]. N [294, 504, 548]. n = 4 [606]. t [706]. t = 3 [706]. ∨ [461]. ∧ [461].

-Control [309]. -D [149, 706]. -Designs [706]. -Dimensional [294]. -intervals [694]. -Matrices [130, 434, 518, 594]. -matrix [46]. -Subdistributivity [461]. -Superdistributivity [461].

1 [569, 752]. 10th [500, 546]. 11th [608]. 1916-2001 [493]. 1990 [51]. 1991 [25–30]. 1992 [80, 118]. 1993 [31, 782]. 1996 [289]. 1998 [382]. 19th [466].

2 [753]. 2000 [464–467, 475, 483]. 2002 [555]. 2003 [562, 571–574, 584]. 2004 [608]. 2005 [654]. 2006 [692]. 2007 [697, 698]. 20th [633]. 22th [573].

360/370 [50].

4 [610]. 4th [392, 572].

5 [441].

60th [323]. 6th [465].

7-10 [159]. 77 [74, 98, 120].

8-vo [30]. 8th [29, 30].

’92 [80, 118, 781]. 93 [775]. 94g [175]. 98 [393].

= [366].

A. [51, 350]. Aberth [368]. absolute [235, 285]. Abstract [306]. Abstracts [52, 91, 782]. ACA’2000 [465]. Accelerated [437, 543, 561]. acceleration [258, 280]. Accumulation [311]. Accuracy [111, 180, 291, 453, 528, 531, 560, 712, 714]. Accurate [214, 409, 638]. ACM [332, 633]. ACRITH [88]. Actual [25, 26]. Adams [71, 180]. adaptive [276]. Add [437]. Addresses [35, 42, 95, 173, 215, 230, 242, 255]. Advanced [27]. AE [629]. AE-Solution [629]. Affine [532, 620, 648, 751]. Aided [75, 176, 682]. aims [99]. Aktyalnije [26]. Alamos [639]. Aleatory [766]. Alexandrov [350]. Algebra [72, 213, 214, 237, 465]. Algebraic [18, 56, 77, 108, 131, 155, 159, 194, 203, 245, 286, 297, 300, 374, 465, 470, 605, 629, 660]. algebraicheskim [78]. Algorithm [63, 68, 137, 142, 150, 179, 277, 296, 323, 360, 409, 416, 436, 462, 481, 510, 516, 517, 540, 583, 593, 597, 617, 620, 623, 625, 638, 641, 656, 681, 687, 688, 695, 726, 732, 756]. Algorithmic [50, 99, 602]. Algorithms [5, 73, 84, 185, 273, 276, 342, 437, 438, 486, 621, 654, 670, 700, 718, 770]. All-Union [12, 25, 26, 29, 30]. Allow [378]. Almost [380]. Also [356]. Alternative [673]. Alternatives [642]. ALU [719]. Ambiguity [525]. Amendments [243, 244]. American [466, 527, 573, 675, 698]. amplitude [63, 87]. Analiz [23]. Analyse [160, 161, 772]. Analysis [22, 23, 31, 65, 85, 106, 116, 315, 364, 422, 444, 445, 447, 448, 453, 479, 486, 488, 524, 525, 535, 548, 574, 581, 663, 664, 668, 674, 678, 695, 704, 709, 767, 768, 772, 782–784]. Analysts [467]. Analytic [141, 219, 455, 485]. Analytical [191]. Announcement [31, 97, 159, 239, 321, 359, 383]. Announcing [169]. Annual [382, 527, 675, 698]. ANSI [21]. ANSI/IEEE [21]. Antipolis [538]. Antwerp [383]. Anwendungen [39, 780]. Any [598]. apparata [10]. Applicability [360]. Application [10, 76, 114, 138, 142, 184, 217, 224, 245, 278, 346, 347, 365, 446, 474, 591, 696, 724, 782]. Applications [51, 53, 75, 206, 212, 251, 270, 274, 282, 299, 330, 331, 351, 398, 464, 465, 467, 470, 593, 607, 623, 633, 640, 647, 668, 672, 679, 737, 769]. Applied [25, 26, 89, 121, 163, 164, 192, 227, 625, 633]. Approach [49, 65, 103, 131, 193, 238, 245, 264, 300, 310, 326, 395, 498, 517, 629, 668, 669, 674, 704, 710, 768]. Approximants [450]. Approximate [110, 503, 549]. Approximation [100, 311, 362, 383, 453, 544, 580, 702, 765]. approximations [156]. April [697]. Arbitrarily [606]. Arbitrary [179, 416, 616, 770]. Arbitrary-Precision [416]. Archetypal [709]. Architectures [523]. arifmetiky [21]. Arithmetic [101, 119, 123, 135, 136, 141, 155, 168, 175, 196, 198, 199, 226, 245, 247, 252, 259, 261, 292, 309, 354, 393, 416, 419, 420, 426, 444, 461, 473, 475, 500, 523, 540, 542, 546, 581, 608, 616, 618, 634, 648, 651, 662, 715, 725, 735, 740, 751, 759]. Arithmetics [105, 324, 683]. Arrowhead [462]. ASIAS [22, 23]. Aspects [720]. Assisted [314, 559]. Association [465]. Asteroids [536]. astronomy [87]. Asymptotic [68, 112, 553]. Asymptotically [593]. Asynchronous [313, 484]. Atlanta [466]. Attacking [315]. August [165, 574]. Austria [168]. Author [337]. auto [100]. auto-correction [100]. Automata [708]. Automatic [22, 23, 31, 69, 75, 121, 219, 320, 457, 486, 687, 726, 782–784]. Automatically [132, 345]. Automation [464]. Autonomous [448]. avtomaticheskix [23]. Award [250, 335]. Awarded [609]. Axes [544].

B [24, 39, 589]. B-splines [589]. Back [332, 405]. Back-to-Back [332]. Baja [654]. balancing [207]. Band [39]. Bangkok [483]. Base [497]. Based [67, 124, 125, 142, 144, 211, 223, 291, 385, 412, 478, 540, 567, 622, 677, 722]. Basel [778]. Basic [535]. Basin [669]. Basis [621, 680, 761, 769]. Bauch [39]. Be [140, 379, 439, 558]. Bearing [759]. Because [378]. Become [378]. Behaviour [134]. Berlin [160, 161, 778]. Bernstein [137, 433, 446, 547, 569, 710, 722, 760–762, 764, 767, 769]. Best [229, 250, 335]. Better [139]. Between [355, 507, 750]. Bibliography [32, 54, 92, 170, 182, 183]. Bibliothek [39]. Bicentered [552]. Biochemical [664]. Birkhäuser [778]. birthday [323]. Bisection [142]. Bisectors [544]. Bishkek [29, 30]. Bivariate [339]. Bliek [396, 441, 689]. Block [179, 502, 719]. Board [35, 42, 95, 173, 215, 230, 242, 255]. Book [24, 51, 160, 161, 368, 381, 402, 778]. Boolean [303]. Boston [778]. Both [132]. Bound [436, 565, 621]. Boundary [144, 188, 316, 561, 626]. Boundary-Based [144]. Bounded [63, 285, 402, 474, 778, 779]. Bounding [5, 120, 296, 354, 477, 514, 722, 763]. Bounds [140, 192, 241, 246, 310, 339, 346, 397, 410, 413, 485, 490, 509, 528, 549, 564, 613, 659, 681, 686, 750]. Box [424, 503, 763]. Box-Plane [424]. Bracket [577]. Branch [436, 621]. Branch-and-Bound [621]. Brief [25, 26, 349]. bright [227]. BSB [39]. Bsesoyuznaya [26]. Bsesoyuznoe [12]. Bsesoyuznovo [30]. bubble [290]. Bulgaria [166, 571]. Bulgariya [167]. Buneman [179]. Burgers [267].

C [127, 368, 618, 711, 776]. C-XSC [711]. C/C [618]. caches [263]. CAE [160, 161, 772]. Calculating [681]. Calculation [146, 409]. Calculations [670]. Calculus [134, 357, 388]. California [654, 698]. Call [31, 97, 159, 169, 228, 239, 240, 249, 251, 321, 351, 383]. Cambridge [51]. Can [140, 424, 439, 516]. Canada [555]. Canadian [229]. Capabilities [717]. Carlo [264, 577, 672, 673]. Cartesian [328]. Case [301, 516, 535, 548, 551, 606, 638, 690, 706]. Catania [467]. Cauchy [497]. caused [19]. CCA [574]. CELL [712]. Centered [327]. Certifying [723]. Chains [508]. Challenges [264, 524]. Characteristics [599]. Characterization [223, 588, 630, 691]. characterizing [269]. Cheap [140, 490]. Chebyshev [258, 622]. Check [223, 558]. Checking [617]. Chemical [10, 142, 217, 667, 668]. Chiang [584]. Chicago [573]. Ch’in [687]. chislennix [11]. Choice [329, 736]. Cholesky [502]. Choose [516]. chosen [229]. Cincinnati [574]. Class [84, 286, 484, 625]. Classes [6, 297]. Classification [49, 567]. Claude [493]. CLINAID [124, 211]. closed [219]. Closing [706]. closure [87]. closure-amplitude [87]. closure-phase [87]. Clouds [586]. cluster [106]. CMMs [222]. co [102]. co-integration [102]. Cobb [344]. COCOS’02 [538]. Codes [389, 438]. Coefficients [433, 482, 485, 592, 741, 764]. Coherence [435]. Colleagues [1, 2, 16, 17, 38, 44, 58, 59, 81, 82, 152, 153, 209, 210, 231, 232, 283, 293, 305, 406, 442, 449, 476, 539, 554, 575, 585, 665]. Collision [371]. Column [304]. Column-Vague [304]. Combination [496]. Combinations [268, 432]. Combinatorial [762]. Combined [205, 569]. Combining [315, 625, 670]. Coming [738]. Comment [482]. Common [125, 454, 478]. Comparative [447, 746]. Comparing [407]. Comparison [60, 715]. compatibility [9]. Compatible [317]. Compensators [422]. Compiler [414]. Complementarity [594]. Complex [219, 225, 258, 322, 504, 618, 634, 636, 683, 707, 755]. Complexity [157, 178, 260, 318, 332, 349, 381, 574, 581, 602, 670, 761, 762, 777]. Complicated [505]. Components [599]. Componentwise [418, 549, 625, 749]. Composite [443]. Comput [175, 610]. Comput./Interval [175]. Computability [574]. Computation [11, 27, 88, 97, 115, 148, 165, 198, 233, 237, 238, 276, 292, 346, 391, 399, 408, 427, 433, 450, 454, 457, 511, 549, 560, 564, 568, 597, 674, 764]. Computational [178, 260, 315, 381, 391, 602, 659, 670, 697, 720, 724, 760, 777]. Computations [3, 4, 15, 32, 37, 43, 48–50, 54, 57, 80, 89, 92, 103, 109, 118, 122, 130, 166, 169, 170, 182, 183, 185, 212, 216, 227, 248, 253, 290, 291, 295, 334, 380, 381, 401, 439, 459, 466, 467, 506, 558, 574, 633, 635, 671, 714, 774, 777]. Compute [212, 328]. Computed [750]. Computer [50, 72, 75, 77, 159, 163, 164, 168, 176, 188, 252, 292, 310, 314, 393, 465, 475, 500, 546, 559, 608, 682]. Computer-Aided [75, 682]. Computer-Algebraic [77, 159]. Computers [19, 465]. Computing [20, 119, 168, 249, 251, 252, 269, 270, 282, 285, 291, 303, 308, 315, 330, 332, 334, 351, 393, 397, 413, 441, 466, 475, 481, 500, 501, 507, 513, 526, 546, 555, 571, 599, 608, 633, 638, 666, 670, 676, 700, 701]. concept [116]. Concerning [603]. Concurrent [48, 319]. Condition [111, 194]. Conditions [66, 70, 558, 695]. Conference [12, 25, 26, 31, 77, 80, 90, 97, 118, 159, 166, 333, 369, 383, 392, 465, 466, 475, 483, 526, 527, 555, 562, 572, 573, 654, 675–677, 692, 698, 781–784]. Conflict [770]. Congress [163, 164, 464, 467]. Conjecture [315, 401]. connected [102, 106]. Connections [314]. Connectives [147, 588]. Consistencies [407]. consistency [223]. Consistent [617]. Constant [275, 511]. Constrained [421, 538, 620]. Constraint [320, 429, 496, 503, 512, 538, 614, 621, 624, 717, 769]. Constraints [138, 184, 682, 725, 731, 741]. Constructing [191]. Construction [60, 435]. Constructions [650]. Constructive [352, 618]. Contents [36, 55, 79, 96, 117, 174, 337, 534]. Context [615]. Continuity [748]. Continuous [661]. Contracting [598]. Contractor [745]. Contribution [641]. contributions [240]. Control [60, 89, 111, 113, 224, 245, 278, 309, 331, 344, 347, 351, 385, 392, 443, 444, 446, 478, 479, 568, 623, 625, 637, 740, 772]. controlled [108]. Controllers [445, 663]. Controlling [536]. Convection [559]. Convergence [286, 434, 547, 615, 761]. Convex [268, 409, 434, 566]. Cooperating [319]. Coordinates [328]. coprocessors [247]. Correct [426]. correcting [262]. Correction [100, 175, 714]. Corrections [595]. Correctly [143]. Correlation [578]. Covering [510]. CRAY [221]. Credal [567]. Criteria [731]. Criterion [358, 541]. Criticism [618]. Crossing [436]. CSAM [163, 164]. CSAM-93 [163, 164]. Csets [579]. Cumulative [542]. curve [219]. Curves [544, 660, 688]. Cycle [358]. Cylindrical [503].

D [39, 149, 350, 706]. DAEs [743]. Data [19, 61, 64, 128, 138, 184, 288, 318, 381, 451, 564, 589, 596, 606, 613, 628, 630, 638, 658, 686, 709, 777]. database [288]. Dawning [423]. Dawson [62]. Days [405]. Dealing [646]. Dear [1, 2, 16, 17, 38, 44, 58, 59, 81, 82, 152, 153, 209, 210, 231, 232, 283, 293, 305, 406, 442, 449, 476, 539, 554, 575, 585, 665]. December [483, 584, 692]. Decidability [735]. Decision [114, 416, 731]. Decomposition [155, 502, 503]. Decreasing [408]. Dedicated [323]. defects [234]. Defined [429, 693]. Definite [61, 276, 598, 703]. Definiteness [699]. Definition [715]. Definitions [181].

Degeneracies [428]. Delaunay [426, 427]. Delay [553]. Dense [729]. Density [132]. DEnv [540]. Department [24, 89]. Dependable [513, 563]. Dependence [691]. Dependencies [489, 644]. Dependency [354, 385]. Dependent [188, 371, 482, 510, 740]. Depending [643]. Derivation [396, 441]. Derivative [116, 361, 460]. Derivatives [8, 212, 312, 325, 579]. Deriving [134]. Describe [349]. Described [550]. Describing [411]. Description [349]. Design [75, 192, 226, 237, 348, 444, 528, 556, 663, 667, 682, 772]. Designs [143, 706]. Destructive [487]. Detection [371, 602, 739, 770]. Determination [69, 219, 540]. Determining [699, 744]. Developing [621]. Development [542, 667, 695]. Developments [624]. Devoted [291]. diagnostics [10]. diagnostiki [10]. Diagrammatic [307, 535, 548]. Diego [698]. Different [181, 746]. Differentiable [294]. Differential [45, 69, 83, 102, 180, 374, 413]. Differentiation [121, 687]. Digital [719]. Digits [570]. Dimension [179]. Dimensional [60, 85, 294, 301, 386, 436, 535, 548]. Dimensions [548]. Diophantine [663]. Direct [130, 649]. Dirty [576]. Disconnected [748]. Discrete [157, 399, 419, 551, 596, 637]. Discrete-Event [399]. Discretized [650]. Disk [720]. Disks [156, 236, 504]. Distance [409]. Distance-Calculation [409]. Distribution [540, 542]. Distributions [345, 578]. Distributivity [471]. Division [263, 517]. dla [88]. dlya [10, 11, 20]. do [104]. Dobronets [24]. Domains [233, 505]. dopykskax [9]. Dot [390]. Double [290]. Douglas [344]. Dr [323]. due [361]. Duhamel [497]. Duplicates [593]. Dynamic [64, 99, 111, 207, 551, 626, 772]. Dynamical [395, 685, 708, 740, 767]. Dynamics [736]. Dynamischen [160, 161, 772].

Early [687]. earthquake [212]. earthquake-resistant [212]. easily [223]. Easy [558]. Easy-to-Check [558]. Ecological [673]. Economics [676]. Editor [653]. Editorial [35, 42, 95, 173, 215, 230, 242, 255, 468]. Effect [370, 536]. Effective [450, 495]. efficiency [287]. Efficient [140, 155, 385, 425, 512, 617, 645, 729, 756, 764]. Eigenpairs [703]. Eigenvalue [176, 258, 644, 703, 745]. Eigenvalues [177, 750]. Element [352, 669, 692]. Elementary [74, 120, 143]. Elements [397, 532, 680]. Eliminating [593]. Ellipsoidal [71, 83]. Elliptic [188, 316]. Embedded [308]. Embeddings [306]. Empirical [615, 628]. Enclosing [205, 266, 287, 497, 530, 594, 611]. Enclosure [373, 396, 439, 441, 457, 561]. Enclosures [179, 188, 278, 285]. Encyclopedia [51]. Endpoint [638]. Engineering [27, 77, 88, 90, 159, 212, 217, 274, 465, 475, 666, 668, 670, 672, 674, 680, 692, 781]. Enhanced [414, 714]. Entries [502]. Entropy [248, 340, 610, 700]. Entwurf [160, 161, 772]. Envelope [540]. Envelopes [474, 578]. Environment [199, 728]. Epistemic [388, 533, 766]. Epsilon [329, 375]. Epsilon-Inflation [329, 375]. Equality [512]. Equation [142, 177, 267, 287, 304, 314, 395, 413, 663]. Equation-Based [142]. Equations [45, 51, 69, 83, 102, 107, 131, 180, 187, 190, 197, 205, 235, 246, 266, 285, 300, 301, 352, 353, 396, 417, 431, 432, 441, 482, 484, 490, 510, 515, 522, 530, 535, 548– 550, 605, 606, 643, 652, 707, 729, 773]. Erdos [289]. Errata [610]. Erratum [441]. Error [71, 100, 241, 274, 310, 311, 327, 352, 391, 392, 410, 474, 478, 486, 549, 628, 659, 749, 750]. Error-Bounded [474]. Errors [19, 63, 149, 257]. estimate [19, 235]. Estimates [67, 69, 71, 83, 87, 139, 157, 274, 352, 400, 749]. Estimating [105, 149, 363, 542, 673, 718]. Estimation [114, 122, 203, 327, 391, 399, 415, 447, 480, 491, 531, 629, 690, 708]. Estimations [522]. Evaluation [312, 545, 591]. Even [342, 377, 606]. Event [134, 399]. events [281]. Every [439]. Exact [192, 397, 427, 504, 558, 565, 613]. Example [395, 492, 519]. Examples [288, 529]. Excel [543]. Exchange [658]. Exclude [377]. Exclusion [325]. Exclusive [367]. Existence [176, 313, 358, 495, 521]. Exists [617]. Expanded [747]. Expansion [446, 622, 767]. Expectations [565]. Experience [89, 659]. Experiences [541]. Experimental [63, 64, 135]. Experiments [671]. expert [110, 224]. Exploration [429]. Exponent [714]. Exponential [685, 690]. Expression [303, 390, 738, 754]. Expressions [155, 390]. Extend [502]. Extended [105, 199, 557]. Extension [116, 181, 648, 689, 752, 753]. extensions [6, 72]. Extrapolated [734]. Extrapolation [671]. Extrema [150, 273]. Extremal [186]. Extremely [714].

Factorable [162, 298]. Factorization [455, 728]. Factors [104, 478]. Fast [212, 274, 312, 408, 583, 612, 638, 703]. Fat [765]. Feasibility [381, 462, 580, 641, 777]. Feasible [531]. February [31, 782, 784]. Feedback [704]. Few [412]. Fields [459, 526]. filib [738]. Finance [676, 737]. finder [225]. Finding [186, 325, 432, 436, 550]. Finite [85, 291, 352, 389, 508, 576, 613, 630, 669, 680, 692]. finite-dimensional [85]. First [97, 239, 246, 254, 299, 321, 359, 383, 436, 530]. First-Zero-Crossing-Point [436]. Fischer [280]. Fitting [628]. Fixed [520]. Floating [195, 199, 416, 486, 719, 759]. Floating-Point [199, 759]. Flows [374]. Flowsheeting [142]. Fluents [388]. FNG’05 [654]. Forecasting [696]. Foreword [201, 338, 442]. Form [107, 327, 520, 547, 552, 569, 600, 634, 644, 648, 722]. Formal [515, 629, 770]. Formal-Algebraic [629]. Forms [537, 569]. Formulae [657]. Formulas [146, 202]. Formulation [678]. FORTRAN [74, 98, 120, 414]. FORTRAN-77 [74, 120]. Forward [486]. fractionally [278]. fractionally-linear [278]. Frames [678]. Framework [766]. France [538]. Fréchet [116]. Fredholm [530]. Free [459, 561, 607]. Frequencies [528]. Frequency [560]. Froissart [451]. Frontiers [572]. Fukuoka [608]. Function [74, 150, 219, 269, 273, 279, 370, 398, 411, 421, 438, 439, 453, 542, 547, 557, 569, 617, 763]. Functional [489, 591]. Functions [6, 98, 101, 105, 119, 120, 132, 139, 141, 143, 148, 162, 175, 198, 212, 277, 278, 280, 294, 296, 298, 325, 347, 373, 432, 436, 455, 485, 505, 521, 532, 545, 547, 550, 560, 578, 615, 661, 701, 702, 724, 758]. further [31]. future [216]. Fuzzy [147, 229, 347, 365, 466, 507, 527, 562, 573, 586–593, 654, 675, 676, 698, 718, 737]. fynktsiı [6].

G [39, 160, 161, 402, 499, 728]. GAMM [165, 241, 393, 475, 500, 546, 608]. Gasarch [402, 778]. Gauss [7, 276, 434, 472, 517]. Gaussian [146, 323, 360, 462, 620, 641]. Gell [349]. Gell-Mann [349]. General [112, 220, 281, 422, 516, 644, 648]. Generalization [548]. Generalized [119, 142, 415, 472, 541, 579, 680, 686, 703, 752, 753]. Generating [528]. Generation [320, 474]. Genetic [654]. Geodetic [328, 607]. Geometric [290, 334, 614, 650, 658, 769]. Geophysical [363]. Georgia [466, 778]. Geosciences [251]. Geospatial [593]. Gerhard [323]. German [772]. Germany [165, 252, 475]. Get [139, 279, 310]. Given [111, 599]. Global [73, 127, 133, 207, 325, 340, 353, 389, 394, 498, 538, 541, 543, 556, 569, 572, 610, 615, 621, 625, 717, 732, 733]. Globally [768]. GNU [414]. goda [167]. Good [405]. Gradients [545, 579]. Grand [524]. Granular [466]. Graphs [456]. gravitation [275]. Greece [572]. Gregory [499]. Group [526, 588]. Guarantee [111, 389]. Guaranteed [69, 122, 129, 180, 219, 259, 291, 528, 690, 708, 739]. Guaranteeing [693]. Guided [747].

H [39]. H. [39]. Half [327]. Hamiltonian [358, 739]. Handling [563, 667]. Hansen [140, 396, 441, 689]. Hard [154, 285, 297, 303, 377, 397, 531]. Harder [391]. Hardness [227, 372, 684]. Hardware [143, 145, 226, 400, 651]. Hausdorff [661]. Hawaii [464, 697]. Heat [559]. Heidelberg [160, 161]. Heindl [323]. Help [290, 348, 349, 401]. Helpful [379]. Hermite [413]. Hermitian [258]. Heteroclinic [314]. heuristic [155]. heuristics [227]. High [206, 246, 261, 374, 386, 489, 495, 756]. High-Dimensional [386]. High-Level [756]. High-Order [374, 489, 495]. Higher [322, 547, 569, 721]. Highlights [382]. Hilbert [380]. Hill [177]. Homotopy [693]. Honolulu [697]. Horner [687]. Householder [636]. Huang [610]. Hull [130, 376, 411, 431]. Hurwitz [191]. Hybrid [453, 708, 732]. Hydrologic [364]. Hyper [379]. Hyper-Spectral [379]. Hyperbolic [197].

IBM [50]. ICRA99 [383]. Idea [342]. Idempotent [488]. Identification [64, 245]. Identity [497]. IEEE [21, 199, 562, 697]. If [279, 377, 378, 617]. IFIS [208]. II [54, 292, 548]. III [92]. Illinois [573]. IMACS [165, 241, 393, 465, 475, 500, 546, 608]. IMACS/GAMM [393]. Image [366]. Images [379, 748]. Imaging [87, 348]. Implementation [127, 188, 207, 720, 756]. Implicit [425]. Implicitly [429]. Imprecise [19, 566, 646]. Imprecision [145]. impressions [248]. Improper [471]. Improve [578]. Improved [394]. Improvement [649]. Improving [287, 674]. inaccurate [61]. Inclusion [46, 156, 177, 236, 294, 317, 322, 547, 569, 595, 612, 615, 701, 721, 755, 758]. Inclusions [62, 328, 411, 458]. Incomplete [564, 686]. Independence [684]. Index [287, 337]. Index/Volume [337]. Indirect [149, 274, 391]. industry [10]. Inequalities [85, 587, 693]. Inexact [426, 630]. Inference [124]. Infinite [576]. Inflation [329, 375]. Informal [206, 526]. Information [28, 31, 66, 139, 466, 527, 573, 675, 677, 698]. Initial [66, 398, 413, 601]. Injectivity [695]. Inner [203, 327, 491, 629, 702]. Input [326, 591]. Inputs [303, 426]. Institute [526]. instrument [3]. INTBIS [221]. InTech’03 [584]. InTech’2000 [483]. Integral [62, 233, 246, 580]. integralnix [23]. Integrals [148, 276]. Integrated [728]. Integration [102, 374, 430, 726]. Intelligence [697]. Intelligent [224, 278, 483]. Interactive [429, 590]. interface [226]. interior [219]. International [77, 90, 97, 159, 163–166, 168, 169, 241, 249, 251, 252, 291, 299, 369, 383, 392, 393, 465, 466, 475, 483, 500, 546, 562, 572, 573, 608, 654, 676, 677, 692, 781–784]. Interplay [728]. Interpolants [450]. Interpolation [128, 180, 224, 341, 451]. Interrupted [342]. Intersection [376, 425]. Intersections [424]. Intersplines [768]. Interval [15, 28, 37, 43, 57, 62, 67, 77, 80, 90, 118, 119, 122, 130, 133, 135, 136, 159, 194, 196, 292, 295, 303, 312, 313, 317, 328, 340, 341, 346, 347, 362, 364, 366, 377, 381, 382, 388, 397, 399, 401, 411, 421, 429, 431, 436, 447, 467, 475, 482, 483, 489, 494, 496, 509, 512, 516, 540, 541, 556, 572, 583, 592, 598, 604, 611, 613, 620–622, 632, 634– 636, 646, 647, 652, 654–656, 661, 673, 683–685, 691, 696– 698, 707, 709, 715, 716, 719, 752, 753]. Interval [22, 23, 53, 75, 87, 98, 102, 123, 124, 127, 128, 162, 169, 178, 181, 185– 187, 193, 198, 208, 253, 254, 259, 290, 298, 299, 306, 307, 324, 332, 343, 356, 360, 373, 376, 379, 380, 391, 396, 400, 413, 422, 439, 441, 448, 457, 459, 461, 465, 466, 488, 495, 507, 508, 517, 524, 527, 532, 535, 543, 548, 557, 558, 573, 581, 582, 584, 588, 596, 602, 605, 625, 640, 641, 644, 651, 663, 668, 670, 672, 675, 677, 680, 692, 695, 699, 722, 729, 730, 738, 744, 773, 777]. Interval [4–7, 12, 18, 29, 30, 32, 47, 60, 64, 65, 68, 72, 76, 110, 141, 142, 144, 147, 150, 157, 158, 179, 189, 191, 199, 206, 207, 211, 217, 220, 264, 265, 273, 296, 300, 309, 318, 323, 326, 333, 355, 365, 387, 414, 415, 419, 430, 443, 444, 462, 464, 472, 479, 484, 487, 491, 498, 502, 511, 515, 522, 523, 525, 551, 553, 562, 570, 593, 599, 601, 606, 607, 610, 612, 616, 626, 628, 637– 639, 642, 643, 664, 669, 676, 682, 700, 704, 718, 725, 735, 740–742, 768]. interval [3, 10, 11, 49, 51, 54, 56, 84–86, 89, 92, 99, 101, 105–108, 112–114, 116, 154, 170, 175, 182, 183, 202, 203, 212, 216, 226, 227, 229, 235, 245, 247, 248, 266, 277– 280, 285, 286, 288, 404, 574, 774, 781]. INTERVAL-92 [90]. Interval-Affine [620]. Interval-Based [211, 540]. Interval-Computations [439]. Interval-Enhanced [414]. Interval-Related [299, 333, 527, 562, 572, 573, 584, 640, 646, 654, 675, 676, 692, 697, 698]. interval-segment [116]. Interval-Valued [124, 347, 365, 388, 508, 599]. INTERVAL’94 [77, 159]. Interval’98 [369]. Intervallmathematik [780]. intervally [111]. Intervalmathematik [39]. Intervalnije [11]. intervalnix [6]. Intervalnoi [10, 12, 30]. Interval’nye [3, 15, 37, 43, 57, 80, 118]. intervalnym [78]. Intervals [8, 104, 126, 129, 132, 134, 195, 204, 223, 224, 256, 281, 299, 308, 345, 348– 350, 354, 356, 357, 367, 378, 420, 435, 445, 456, 463, 469– 471, 474, 518, 576, 579, 586, 591, 657, 658, 694, 737, 746, 752–754]. INTLIB [74]. Intractable [154, 378]. Introduction [185, 206, 424]. Introductory [666]. Invariant [551, 591]. Inverse [176, 397, 518]. inverses [5]. Inversion [489, 623]. Inversions [720]. Invertibility [420, 505]. Investigation [615]. Invited [464]. Involving [128, 265]. IPMU’2006 [677]. IR [435]. Irrational [162, 298]. ISIPTA’05 [646]. isledovaniĭ [3]. ISO [658]. Isolation [395]. Isomorphic [306]. Issue [80, 118, 169, 251, 334, 351, 424, 507, 513, 555]. Italy [467]. Iteration [155, 313, 418]. Iterations [434]. Iterative [107, 220, 286, 749]. Iteratively [611]. IV [170]. IVP [495, 743].

J [381]. Jacobi [434]. Jacobians [545]. Jahn [39]. Japan [608]. Journal [169, 249, 251, 291]. July [382, 466, 467, 573]. June [27, 383, 464, 465, 571, 572, 698]. Justification [259, 329]. Justified [342].

Kahl [381]. Kantorovich [603]. Karl [403, 705]. Karlsruhe [475]. Kaucher [245, 259, 461]. Kearfott [396, 441]. Kepler [401]. Kernel [458, 530]. Kernel-Splitting [530]. Kharitonov [443]. Kharitonov-Like [443]. Kind [246, 530]. klassov [6]. Knowledge [124, 211, 365, 388, 677]. Knowledge-Based [124, 211]. Known [565]. kollegi [1, 17, 44, 59, 82, 153, 210, 232]. Kolmogorov [129]. Komplex [23]. kompyuternim [164]. kompyuterno [78]. kompyuterno-algebraicheskim [78]. Konferentsiya [26, 78, 167]. Kongress [164]. Krawczyk [418]. Kreinovich [381]. Kuramoto [314, 395].

L [705]. Laboratory [639]. Lafayette [31, 782, 784]. Lagrange [341]. Lakeyev [381]. Lanczos [410, 452]. Lanczos-Type [452]. Landen [311]. Language [20, 50, 88, 115, 206, 756]. languages [27]. Large [46, 99, 138, 184, 192, 258, 267, 484, 505, 679, 729]. large-scale [99]. Largest [661]. latency [263]. Later [380]. Lead [684]. Leads [224, 367]. Learn [424]. Least [454, 607]. Least-Squares [607]. Lecture [349]. Legendre [276]. Leipzig [39]. Leningrad [27]. Letter [653]. Level [206, 756]. libavi.a [257]. Libraries [125]. Library [74, 98, 213, 237, 257, 616, 622, 659, 751]. Lifetime [576]. Like [443, 595, 755]. limit [204]. Limitation [473]. Limits [537]. line [261]. Linear [9, 18, 46, 56, 84, 111, 125, 150, 154, 184, 187, 189, 194, 203, 213, 214, 220, 235, 237, 245, 260, 266, 278, 285, 286, 297, 300, 304, 318, 320, 353, 377, 392, 394, 396, 410, 415, 431, 432, 441, 457, 472, 482, 491, 494, 512, 515, 522, 532, 535, 548, 549, 551, 582, 590, 591, 594, 605, 606, 626, 630, 636, 643, 649, 652, 659, 661, 679, 686, 689, 691, 693, 700, 707, 712, 723, 729, 741, 742]. Linear-Time [150, 700]. Linearization [682]. Linearly [643]. lineinoi [9]. Linkages [507]. List [52, 91]. load [207]. Local [133, 150, 186, 273, 347]. Localization [151, 448]. localizational [49]. locate [273]. Locates [150]. Locating [269]. Logarithm [576, 597]. Logic [512, 588]. long [192, 290]. long-standing [290]. Loosing [598]. Louis [562]. Louisiana [31]. Ludyk [160, 161]. machines [149]. Made [140]. Mai [584]. Making [114, 731]. Management [677]. Mann [349]. manufacture [10]. manuscript [13, 14, 33, 34, 40, 41, 93, 94, 171, 172]. Many [565]. Map [461]. Maple [135]. Mapped [757]. Mappings [748]. Maps [376]. March [31, 77, 159, 782]. Margin [699]. Market [696]. Markov [508]. Martin [402, 778]. mashiniyu [21]. Matematicheskomy [167]. Matematika [52, 91]. matematike [12, 30, 164]. matematiki [10, 26]. Material [487]. Mathematica [123]. Mathematical [97, 166, 291, 315, 365, 438, 622]. Mathematics [10, 12, 25, 26, 29, 30, 51, 53, 76, 89, 114, 163, 164, 206, 465, 507, 782]. Mathematisch [39]. Mathematisch-Naturwissenschaftliche [39]. Matrices [130, 268, 309, 434, 462, 502, 518, 583, 594, 632, 642, 683, 699, 711, 744]. Matrix [5, 46, 60, 68, 176, 233, 235, 397, 598, 612, 745, 750, 764]. Maui [464]. Maximal [134, 491]. Maximum [248, 325]. May [25, 26, 349, 379, 555, 562, 606]. Maybe [342]. Mean [565, 753]. means [85, 107]. measure [279]. Measurement [149, 150, 273]. Measurements [274, 391, 617, 638]. Measures [737]. Mechanics [680, 724]. Mechanising [256]. Medial [544]. Medical [124, 211]. Meeting [382]. Members [35, 42, 95, 173, 215, 230, 242, 255]. Menshikov [499]. meshes [149]. Method [7, 69, 71, 107, 144, 158, 180, 187, 190, 191, 205, 207, 217, 235, 280, 286, 316, 340, 353, 394, 413, 417, 472, 495, 582, 595, 610, 625, 626, 636, 643, 649, 655, 689, 722, 744, 755, 764]. Methodologies [126]. Methods [11, 24, 45, 51, 67, 77, 90, 106, 159, 186, 194, 220, 234, 236, 241, 259, 280, 315, 322, 343, 368, 374, 379, 385–387, 412, 447, 452, 464, 465, 475, 484, 487, 489, 543, 561, 621, 647, 692, 696, 721, 728, 730, 749, 771, 773, 776, 781]. metodam [78]. metodi [11]. Metric [317]. metrization [86]. Mexico [654]. Mezhdunarodnaya [78, 167]. Mezhdunarodnyı [164]. Microsoft [543]. Midpoint [606, 754]. Mini [253]. Mini-Symposium [253]. Minima [325, 347]. minimal [103]. Minimax [419, 480]. Minimization [421, 761]. minimizes [290]. Minimum [133]. Minisymposium [392, 467]. Minkowski [504]. Miranda [603]. MISC’99 [404]. Missouri [562]. MMSC [97, 167]. MMSC-93 [97, 167]. MMSC0 [166]. MMSC0-93 [166]. Modal [326, 445, 469, 474, 752, 753]. Model [288, 344, 385, 734]. Modeling [138, 184, 589, 668, 669, 724]. Modelirovanioyu [167]. Modelling [97, 166, 658]. Models [113, 157, 374, 529, 680, 727]. modifications [107]. Moduli [260, 478]. Monotone [101, 175, 708]. Monotonic [134, 617]. Monotonicity [139, 317]. Monotonicity-Type [139]. Monte [264, 577, 672, 673]. Monte-Carlo-Type [672]. Moore [418, 603, 609]. Morphological [355]. Morphology [365]. Moscow [80, 118, 781]. Motivations [616]. MP [221]. MPFI [616]. MR1253135 [175]. MR1617525 [610]. Multi [60, 662, 731, 748]. Multi-Criteria [731]. multi-dimensional [60]. Multi-Precision [662]. Multi-Valued [748]. Multiaspectness [151]. Multidimensional [547, 655]. Multigrid [316]. Multimedia [523]. multiobjectivity [157]. Multiple [198, 454, 721]. Multiple-Precision [198]. Multiplication [471, 583, 612, 657, 694].

Multivariate [768]. Mystery [463].

NAFIPS [208, 466]. NAFIPS/IFIS/NASA’94 [208]. NAFIPS’02 [527]. NAFIPS’03 [573]. NAFIPS’05 [640]. NAFIPS’06 [675]. NAFIPS’07 [698]. NaNs [265]. Narrow [606]. Narrowest [224, 367]. Narrowing [8]. NASA [333]. NASA’94 [208]. National [639]. Natural [104, 597, 752]. Naturwissenschaftliche [39]. nauchnix [20]. nauchno [88]. nauchno-texnicheckix [88]. Nauka [24]. Navier [417]. Navigation [343]. Naychnym [167]. nayke [78]. NE-factors [104]. nearly [149]. Necessary [281, 558]. necessity [223]. nekotorix [6]. nepreryvnovo [10]. Netlib [659]. Networks [66, 129, 145, 213, 362, 567, 654, 664]. Neumaier [51]. Neural [129, 145, 362, 654]. Neurons [145]. Newton [142, 144, 158, 190, 217, 220, 275, 484, 489, 625, 655, 722]. Newton/Generalized [142]. Nickel [403, 705]. Ning [396, 441]. NMA’98 [392]. No [158, 175, 516]. Nodes [146]. Non [10, 130, 134, 184, 258, 320, 353, 394, 487, 515, 521, 522, 693]. Non- [130]. Non-Destructive [487]. Non-Linear [184, 320, 353, 394, 515, 522, 693]. Non-monotonic [134]. Non-Smooth [521]. non-stop [10]. Nonconvex [444]. Nonlinear [64, 107, 121, 122, 138, 155, 188, 190, 192, 197, 205, 301, 340, 372, 421, 432, 467, 484, 490, 492, 510, 514, 568, 610, 645, 706, 730, 733, 747]. Nonlinearity [685]. Nonnegative [742]. Nonrational [560]. Normal [600]. Norms [713]. North [466, 527, 573, 675, 698]. Notation [570]. Note [280, 375, 460, 726]. notes [248]. Notions [535, 748]. Novel [131]. Novosibirsk [24]. novyı [20]. Nowhere [294]. NP [154, 227, 285, 297, 303, 372, 377, 397, 531, 684]. NP-Hard [154, 285, 297, 303, 377, 397, 531]. NP-Hardness [227, 372, 684]. NP-problems [227]. Number [219, 267, 279, 459, 581]. Numbers [194]. Numerical [11, 24, 31, 197, 234, 241, 267, 275, 299, 301, 320, 368, 395, 417, 454, 541, 695, 756, 771, 776, 782–784]. Numerics [168, 252, 393, 475, 500, 546, 571, 608].

O [9, 368]. OBJ3 [256]. Objective [716]. Objects [64, 658]. Obreschkoff [413]. Obtained [749]. occasion [323]. October [29, 30, 538, 608, 654]. ODE [398, 495, 497]. ODEs [374, 536, 743]. Oelschägel [39]. Oettli [600]. Offsets [544, 660]. oformleniyu [13, 34, 41, 94, 172]. Ohio [574]. Oil [254]. Old [405]. Oldenburg [165]. on-line [261]. One [150, 245, 273, 296, 301, 358, 424, 436, 535]. One-Dimensional [436, 535]. One-Space [301]. Only [420]. Operation [367]. Operations [98, 265, 345, 354, 355, 420, 715]. operator [279]. opinion [610]. Optimal [18, 48, 56, 84, 113, 149, 154, 156, 278, 342, 344, 366, 373, 552, 593, 605, 630, 662, 694, 740, 768]. Optimality [760]. Optimally [598]. Optimization [63, 70, 73, 121, 127, 133, 207, 340, 389, 488, 498, 526, 538, 541, 556, 569, 572, 596, 610, 615, 621, 625, 668, 710, 717, 732, 733]. Optimizing [221]. Order [63, 65, 322, 374, 422, 489, 495, 547, 569, 673, 721, 724, 746]. Order-Preserving [724]. order-theoretic [65]. Ordering [281]. Ordinary [45, 69, 180, 413, 561]. Oriented [310]. Origins [357]. orthogonal [238]. Ostrowski [595, 755]. Ostrowski-Like [595, 755]. Other [378]. Outcome [686]. Outer [300, 415, 532, 582, 629, 643, 644, 649]. Outlier [602]. output [112]. Ov [6]. Overcome [517]. Overdetermined [266, 652, 713]. Overestimation [285, 606, 607, 739]. Overestimation-Free [607]. Overestimations [323]. Overflow [377].

P [381]. p. [160, 161]. Package [22, 23, 135, 136]. Pages [576]. Paper [229, 250, 335]. Papers [31, 97, 159, 169, 228, 239, 249, 251, 321, 351, 383]. Parabolic [301]. Parallel [19, 73, 127, 133, 142, 143, 149, 185–188, 207, 211, 214, 225, 237, 645, 718, 729]. parallelization [220]. Parallelized [510]. Parameter [114, 122, 188, 447, 480, 510, 664, 690, 708]. Parameter-Dependent [188, 510]. Parameters [99, 111, 112, 491, 532, 564, 590, 643]. Parametric [371, 582, 644, 649]. Pareto [596]. Part [32, 162, 298, 535, 548, 569, 752, 753]. Partial [8, 312, 407]. PASCAL [20, 109, 329]. Paso [254, 692]. Path [456]. pathologies [158]. Patrick [336]. Paul [289]. pavement [274]. Pavings [757]. Pawlak [335]. Pearson [578]. Penalty [340, 610]. perovod [21]. Perspective [639]. Perturbation [452]. Perturbations [514]. Perturbed [490, 601]. Perturbing [428]. Pessimism [701]. Petersburg [77, 89, 159, 465]. phase [87]. Phenomenon [451]. Physics [315, 716]. Physiological [727]. Piecewise [550]. Piecewise-Trapezoidal [550]. Pitfalls [624]. Planar [660]. Plane [219, 424, 544]. Planning [456]. Plant [63, 111, 112, 326]. Plants [70, 422]. Plots [560]. Plotting [429, 660]. Point [195, 199, 416, 436, 486, 520, 710, 719, 759]. Points [186, 269]. Polar [634, 660]. polesnijĭ [3]. polnyı [21]. Polygon [444]. Polyhedra [409]. Polyhedron [411]. Polynomial [205, 233, 238, 262, 309, 339, 341, 416, 509, 595, 710, 722, 755, 758, 761, 766, 767]. Polynomials [191, 322, 346, 408, 454, 721]. Polytopic [434]. Population [564, 656, 700]. Populations [613]. portable [74, 109]. Portfolio [592]. Positive [598, 699, 703]. Possibility [587, 716]. Possible [129]. Possibly [748]. posteriori [67]. Potential [672]. Power [427, 557]. Powers [683]. pp. [24]. Practicalities [624]. practice [98]. Prager [600]. pragmatics [104]. Precise [141, 368, 776]. Precision [136, 198, 226, 247, 416, 616, 662]. Preconditioned [431]. preconditioner [155]. preconditioners [7]. Preconditioning [529, 606, 629]. Predictive [445]. Predictors [686]. Predislovie [201]. predmet [3]. Preface [200, 271, 272, 284, 384, 619]. Preference [731]. Preliminary [31, 497]. preparation [13, 14, 33, 34, 40, 41, 93, 94, 171, 172]. Prescribed [560]. Presence [426]. presentation [101, 175]. Presented [299]. Preserve [420]. Preserving [724]. Press [51]. Previsions [566]. prikladnoı [26, 164]. Primenenie [10]. Prize [609]. Probabilistic [391, 587, 670]. Probabilities [576, 646]. Probability [132, 345, 586, 587, 684]. Probability-Possibility [587].

Problem [9, 112, 176, 178, 189, 300, 341, 385, 391, 398, 413, 428, 494, 592, 598, 644, 728, 762]. Problem-Solving [728]. problemij [26]. Problems [25, 26, 61, 70, 128, 181, 188, 217, 227, 245, 258, 290, 316, 318, 378, 380, 419, 421, 446, 488, 559, 561, 581, 594, 596, 601, 614, 630, 684, 703, 712, 724, 731, 732, 735]. Procedure [528]. Proceedings [80, 118, 555, 783]. Process [142, 410, 629]. Processes [60, 89, 667]. Processing [257, 366, 381, 400, 466, 527, 573, 672, 675, 677, 698, 719, 777]. Processor [712]. Produce [342, 606]. Produced [323]. Product [356, 390, 452, 565, 754]. Production [344]. Products [202, 504]. Prof. [323]. professor [229]. Program [22, 23]. Programming [27, 49, 50, 88, 98, 109, 115, 494, 512, 590, 659, 717, 756]. programmirovaniya [88]. Programs [262, 733]. proisvodstva [10]. projection [238]. promyshlennosti [10]. Proof [85, 176, 314]. Proofs [559, 603]. Propagation [621, 727]. Proper [471]. Properties [443, 470]. Proposal [618]. Proving [307, 401]. provision [99]. Purpose [400]. Purposes [760].

QFT [528]. Qi [361, 460]. Quadratic [277, 373, 648, 685]. quadratic-time [277]. Quadrature [146, 275, 386]. quality [246]. Quantification [766]. Quantified [503, 623, 689]. Quasi [48, 317, 713]. quasi-concurrent [48]. Quasi-Metric [317]. Quasi-Norms [713]. Queries [402, 778, 779]. questions [114]. Queueing [625, 728].

R [419, 609]. radio [87]. Radius [238, 744, 754]. radix [261]. Random [299, 354, 540, 542, 565, 577, 650]. Range [105, 120, 139, 327, 339, 439, 506, 531, 671, 702, 714, 736, 763]. Ranges [101, 175, 296]. Rank [268]. rasreshimosti [9]. rasshireniyax [6]. Rational [100, 383, 435, 450, 451, 453, 763]. Ray [425]. Rays [580]. razike [21]. RCA [633]. Reachability [767]. Real [133, 139, 296, 349, 416, 522, 581, 640, 765]. Real-Valued [133, 139, 296, 522]. realistic [149]. Realizations [626]. Reals [319]. Reasoning [110, 132, 211, 513]. reciprocal [263]. Reconstruction [664]. Recursion [402, 778, 779]. Reducing [263]. Reduction [327, 739]. Redundancies [320]. Referativnyi [52, 91]. Refined [581]. Regelungssystem [161]. Regelungssystemen [160, 772]. Regular [757]. Regularity [631, 632, 642, 744]. Regulator [326]. Related [299, 333, 527, 562, 572, 573, 584, 640, 646, 654, 675, 676, 692, 697, 698]. Relation [129, 355, 750]. relational [288]. Relations [307, 429, 715, 746]. relative [285]. Reliab [610]. Reliability [99, 452, 599, 674, 681]. Reliable [103, 213, 249, 251, 270, 282, 291, 310, 330, 334, 344, 351, 371, 441, 480, 501, 507, 513, 560, 571, 597, 633, 648, 666, 668, 678]. Remainder [233, 346]. Remark [273, 302, 365, 662]. Remarks [666]. Reminiscences [405, 440]. Removing [428]. Report [25–27, 165]. represent [104]. Representation [103, 106, 291, 307, 365, 411]. Representations [760]. Representative [516]. Representing [388]. required [111]. Requirements [13, 14, 33, 34, 40, 41, 93, 94, 171, 172, 674]. research [3, 4]. Reshenie [30]. resheniya [10]. resistant [212]. Resolution [29, 30]. Respect [376, 461]. Response [560, 591]. Result [31, 121, 140, 372, 645, 706, 747, 782–784]. Result-Verifying [645, 706, 747]. resulting [259]. Results [150, 273, 299, 354, 412, 577]. Retsinsija [39, 161]. Review [7, 24, 39, 51, 160, 161, 368, 381, 402, 778]. Reviews [270, 282]. Revisited [519, 752, 753]. Rewriting [625]. Reynolds [267]. Right [629, 689]. Right-Preconditioning [629]. Right-Quantified [689]. Rigorous [395, 413, 549, 659]. Rings [192]. Risk [673]. Robot [343, 448]. Robust [273, 425, 444, 448, 479, 544, 637, 660, 663, 688, 704, 740]. Robustly [422]. Robustness [445]. Rohn [381, 396, 441]. Root [287, 481, 765]. Roots [219, 322, 509]. Rounded [143]. Rounding [195]. Roundoff [459]. Roundoff-Free [459]. routings [149]. Rump [519]. Runge [107]. Russia [77, 159, 465]. Russian [1, 3, 6, 9–13, 17, 20, 21, 23, 26, 30, 34, 39, 41, 44, 59, 78, 82, 88, 94, 153, 161, 164, 172, 201, 210, 232]. rykopici [13, 34, 41, 94, 172]. ryskom [21].

S [24, 39]. SAC’2005 [633]. Safe [717]. Saint [465]. Sample [638]. Sampling [222, 673]. San [698]. Santorini [572]. Saratov [25, 26]. Satellite [379]. Satisfaction [538]. Scalar [326]. Scale [99, 637]. SCAN [28, 168, 393, 475, 775]. SCAN-1991 [28]. SCAN-93 [168, 775]. SCAN-98 [393]. SCAN’2002 [546]. SCAN’2004 [608, 609]. SCAN’95 [252]. Scandinavian [647]. Scheduling [645]. Schemes [601]. Schur [238]. Schwarz [484]. Science [77, 90, 159, 465, 475, 692, 781]. Scientific [20, 27, 88, 97, 109, 115, 166, 168, 252, 270, 282, 315, 330, 393, 475, 500, 524, 546, 608]. scientific/engineering [27, 88]. Search [543, 681]. Second [159, 361, 460, 609, 647, 673, 676]. Second-Derivative [460]. Second-Order [673]. see [175]. segment [116]. Seidel [7, 434, 472]. Selecting [710]. Selection [155, 541, 592, 747]. Self [48, 262]. Self-correcting [262]. self-validation [48]. Semi [598]. Semi-Definite [598]. Semigradients [579]. seminar [27]. Sensitive [145]. sentyabrya [167]. September [80, 97, 118, 165, 166, 168, 248, 252, 475, 781]. Sequences [233, 308]. Series [50, 622, 697]. Session [465, 466, 571]. Sessions [208, 464]. Set [108, 203, 223, 299, 454, 482, 490, 507, 550, 568, 571, 577, 596, 623, 661, 690, 691, 693, 708, 741]. Set-Valued [550, 571]. Sets [86, 106, 147, 281, 299, 378, 415, 472, 491, 522, 550, 586, 587, 627, 629, 676, 707]. Shannon [493]. Shape [482]. Sharp [310, 509, 549]. Sharpening [635]. sharper [280]. Sharpness [295]. Shary [342]. Shaydurov [24]. Shell [254]. Shift [437]. Shift-and-Add [437]. SIAM [382, 526]. Siberian [24]. Sicily [467]. side [227]. sided [24, 45, 69, 316, 771]. Sign [304]. Sign-Stable [304]. Signal [400, 719]. Signals [351]. Significance [570]. Similar [204, 391]. Simple [204, 219, 396, 441]. Simplest [516]. Simplicial [761]. Simulation [160, 161, 264, 465, 551, 743, 772]. Simulations [577, 739]. Single [755]. Singular [611, 733]. Singularity [492]. Singularly [601]. sintez [23].

Sinusoidal [447]. Sistem [23]. sistemam [164]. Situation [388]. Sivashinsky [314, 395]. SIVIA [756]. Slant [579]. Slightly [391]. Slope [280, 322, 579]. Slopes [162, 298, 312, 411]. Smooth [521]. smoothing [277]. SMP [728]. SMP/G/1 [728]. Sobeshchanie [12]. Society [466, 527, 573, 675, 698]. Soft [676]. Software [125, 136, 193, 226, 261, 345, 438, 782]. Solar [302]. Solution [18, 46, 56, 83, 84, 108, 111, 125, 154, 203, 235, 260, 341, 342, 353, 394, 413, 415, 446, 472, 482, 490, 491, 495, 515, 522, 530, 532, 536, 550, 582, 605, 627, 629, 643– 645, 649, 691, 707, 712, 723, 727, 741, 742]. Solutions [188, 197, 205, 266, 267, 301, 304, 352, 417, 432, 477, 497, 549, 559, 592, 594, 596, 604, 611, 749]. Solvability [178, 602]. Solve [731]. Solvent [556]. Solver [220, 543, 706]. Solvers [319, 496, 729, 747]. Solving [10, 45, 61, 69, 107, 131, 138, 180, 184, 189, 290, 320, 377, 387, 392, 421, 503, 510, 512, 624, 652, 682, 713, 723, 728, 730, 735, 769]. Some [6, 45, 107, 114, 120, 128, 156, 288, 318, 443, 470, 529, 559, 581, 603, 607]. SONIC’95 [253]. Sophia [538]. Sophia-Antipolis [538]. soveshchania [30]. Soviet [32, 54, 92, 170, 182, 183]. Sozopol [97, 166, 167, 571]. Space [301, 307, 317, 350, 435, 660, 661]. Spanning [178]. Sparse [130, 220, 520, 589, 711]. Special [80, 107, 118, 169, 251, 334, 351, 400, 424, 466, 507, 513, 555, 571]. Specific [691]. Spectral [379]. Speed [286, 615]. Spheres [765]. Spherical [706]. Spigot [481, 597]. Splines [589, 768]. Splitting [530]. Springer [160, 161]. Springer-Verlag [160, 161]. Squares [607]. St [77, 89, 159, 562]. Stability [68, 192, 238, 553, 685]. Stabilizing [422]. Stable [302, 304]. Staggered [714]. Standard [198, 658, 738]. Standards [524]. standartov [21]. standing [290]. State [60, 704, 740]. State-Dependent [740]. Statements [603]. static [70, 245]. Stationary [269, 749]. Statistics [47, 204, 670]. Statool [540]. steady [112]. steady-output [112]. step [286, 658]. Still [342]. STOC’97 [332]. Stochastic [90, 451, 781]. Stock [696]. Stokes [352, 417]. stop [10]. Storage [192]. Straightforward [558]. Strassen [583]. Strategy [506, 710]. Strong [443, 631]. Structural [678]. Structure [691]. Structures [66, 487, 679]. Student [169, 250, 253]. Study [222, 746]. sub [61]. sub-definite [61]. Subdistributivity [376, 461]. Subdivided [433]. Subdivision [506, 710, 761]. Subinterval [541]. Subject [3, 4, 28, 766]. Subjective [576]. Suboptimal [179]. subresultant [233]. subroutines [214]. substitution [102]. Success [747]. Success-Guided [747]. Such [438]. Sufficient [281, 558, 695]. Sufficiently [638]. summarizing [27]. summer [248]. Sums [690]. Superdistributivity [376, 461]. Superinterval [439]. Suppes [336]. Supplement [600]. Support [50, 115, 651]. Surfaces [371, 425]. Surprising [498]. Süsse [39]. sweep [235]. Switchings [740]. Symbolic [292, 529, 625]. Symmetric [703, 745]. Symposium [28, 168, 241, 252–254, 332, 393, 475, 500, 546, 608, 633, 697]. Synchronous [484]. Synthesis [22, 23, 66, 113, 326, 422]. System [60, 69, 83, 98, 109, 124, 203, 211, 235, 302, 348, 599, 604, 605, 723]. Systematic [674]. Systemen [160, 161, 772]. Systems [18, 22, 23, 45, 51, 72, 75, 99, 125, 131, 138, 154, 155, 163, 164, 180, 184, 189, 192, 194, 205, 224, 260, 266, 286, 297, 304, 306, 351, 353, 364, 372, 377, 387, 392, 394, 395, 399, 410, 415, 434, 443, 444, 472, 477, 482, 484, 490–492, 510, 514, 515, 520, 522, 525, 532, 549, 551, 553, 562, 582, 588, 591, 611, 620, 625, 626, 636, 637, 643, 645, 649, 652, 654, 677, 679, 683, 685, 689, 691, 704, 707, 708, 713, 722, 728– 730, 739–742, 747, 766, 767, 769, 772, 773]. systems [46, 56, 84, 107, 108, 110].

Table [412]. Table-Based [412]. Tables [576]. Talks [254, 333, 483, 527, 562, 572, 573, 584, 640, 646, 654, 675–677, 692, 697, 698]. Task [645]. tasks [10]. Taylor [327, 346, 374, 385, 485, 529, 537, 547, 569, 622, 734]. Teaching [89, 181]. Technical [10, 633]. Technique [48, 525, 530]. Techniques [222, 258, 494, 672, 733]. Technologies [483]. technology [229]. Templates [390, 738]. Teorie [39]. Term [192]. Termination [389]. Terms [690]. Test [313, 361, 460, 695]. Testing [68, 193, 487]. Teubner [39]. Texas [692]. texnicheckix [88]. texnicheskoı [10]. texnike [78]. Thailand [483, 584]. Their [129, 278, 464, 633, 647, 670, 672, 728, 742]. Theorem [129, 204, 362, 460, 600, 632]. Theorems [204, 307, 603, 642]. theoretic [65]. Theoretical [329]. Theorie [780]. Theory [75, 188, 229, 256, 299, 332, 402, 452, 478, 507, 577, 581, 778, 779]. There [558, 617]. Thermal [669]. Thin [130]. Third [467]. Three [690]. Three-Terms [690]. Tight [140, 490]. Tijuana [654]. Time [134, 150, 277, 350, 371, 551, 553, 591, 637, 700]. Time-Dependent [371]. Time-Invariant [551, 591]. Tolerable [627, 741]. Tolerance [9, 126, 189, 245, 411, 494]. Tolerances [614, 658]. Tomography [363]. Tool [3, 4, 61, 345, 444, 540]. Topology [317]. Toronto [382, 555]. total [286]. Trace [688]. Track [633]. tracking [112]. Tractable [378, 598]. Trajectories [770]. Transfer [560]. Transform [311]. Transformations [587, 588]. Transparent [196]. transport [66]. transputer [213]. Trapezoidal [550]. Trebovaniya [13, 34, 41, 94, 172]. Tree [567]. Trees [178]. Triangle [339]. Triangles [433]. Triangular [587]. Triangulation [426]. Triangulations [427]. tridiagonal [235]. Truss [679]. turned [403, 499]. Twin [324, 473]. Two [5, 24, 45, 69, 276, 316, 502, 548, 637, 750, 754, 771]. Two-Dimensional [548]. Two-sided [24, 45, 69, 316, 771]. Two-Time-Scale [637]. Type [107, 139, 190, 276, 288, 452, 672, 693]. Types [750].

U [39]. UIA [383]. Unbounded [627]. Uncertain [66, 148, 589, 590, 740]. Uncertainties [297, 363, 679]. Uncertainty [70, 113, 366, 513, 516, 525, 533, 563, 593, 602, 639, 656, 667, 669, 670, 672, 677, 680, 700, 716, 718, 724, 727, 759, 766]. Uncertainty-Bearing [759]. Unconstrained [121]. Underdetermined [387, 477, 730]. underestimate [259]. Underestimation [473]. Underflow [377]. UniCalc [61, 131]. Unifying [766]. Unimodality [684]. Union [12, 25, 26, 29, 30]. Uniqueness [176, 460, 495, 521]. united [203]. Universal [362]. Universe [576]. Universität [165]. University [51, 89, 383]. Unknown [354]. Upper [565, 638]. URC [333]. USA [573, 574]. Use [133, 139, 162, 298, 364, 537, 577, 662]. Useful [3, 4, 438]. User [196]. Using [64, 98, 101, 105, 106, 119, 141, 175, 180, 186, 207, 234, 256, 288, 309, 325, 354, 368, 374, 388, 399, 416, 419, 422, 432, 448, 456, 479, 578, 581, 607, 634, 663, 664, 671, 706, 708, 731, 740, 765, 767, 768, 776]. usually [281]. Uvaszaemije [1, 17, 44, 59, 82, 153]. Uvazhaemye [210, 232].

V [10, 24, 39, 78, 182, 381]. Vague [304]. ValEncIA [743]. ValEncIA-IVP [743]. Validated [165, 168, 190, 252, 393, 475, 485, 500, 526, 546, 555, 608, 624, 681]. Validating [50, 495, 542]. validation [48]. Value [188, 235, 303, 316, 398, 413, 601, 753]. Valued [124, 133, 139, 147, 296, 347, 365, 388, 421, 508, 522, 550, 571, 599, 702, 748]. Values [105, 411]. Variable [102, 136, 150, 226, 247, 273, 296]. Variable-precision [226, 247]. Variables [354, 540, 542, 565, 650]. Variance [638, 656, 700, 718]. Vector [148, 257, 702]. Vector-Valued [702]. Vectors [711]. Verifed [728]. Verification [31, 121, 236, 417, 438, 455, 505, 521, 580, 703, 770, 782–784]. Verifications [197, 301]. Verified [106, 125, 132, 146, 177, 237, 275, 276, 345, 374, 386, 408, 410, 489, 726, 727, 743]. Verifying [389, 645, 706, 747]. Verlag [160, 161]. Verlagsgesellschaft [39]. Version [510, 583]. VI [183]. Via [122, 434, 445, 473, 695]. Vienna [168]. View [574]. Viewpoint [380]. VII [12]. Visual [429]. Viswanath [511]. vo [30]. Vol [441]. Volterra [246]. Volume [337, 534]. Vpernye [21]. VPI [136]. Vychisl [175]. vychisleniı [11, 20, 88]. vychisleniia [3, 15, 37, 43, 57, 80, 118]. Vychisleniiam [167].

W [402]. WAC’2000 [464]. WAI’96 [292]. Wave [48]. Wavelet [366]. Wavelets [366]. Ways [502]. Weak [596]. Weakly [192]. Web [438]. Weights [146]. Whose [532]. Wide [714]. width [202, 235]. Wiebigke [39]. William [778]. within [388, 392]. without [325, 598]. Working [526]. Works [32, 54, 92, 170, 182, 183]. Workshop [29, 30, 53, 165, 292, 299, 404, 533, 538, 639, 647]. Workshops [332]. World [349, 428, 464, 467, 640]. Worst [551]. Worst-Case [551]. Wrapping [370, 398, 536]. Wuppertal [252].

ximicheskoı [10]. XSC [20, 88, 109, 329, 711].

Y-MP [221]. YalAA [751]. Yazik [20, 88]. Years [380].

zadach [10]. zadachi [9]. Zdzislaw [335]. Zero [63, 225, 436, 517, 755]. zeroorder [63]. Zeros [119, 141, 294, 514, 595, 721, 722]. Zhengyu [610]. Zhurnal [52, 91]. Zonotopes [607].

References

Nesterov:1991:UKRa

[1] V. M. Nesterov. Uvaszaemije kollegi! (Russian) [Dear colleagues]. Interval Computations = Interval’nye vychisleniia, 1(1):2–5, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-1-pp-2-5.pdf.

Nesterov:1991:DCa

[2] V. M. Nesterov. Dear colleagues. Interval Computations = Interval’nye vychisleniia, 1(1):6–9, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-1-pp-6-9.pdf.

Yakovlev:1991:IVP

[3] A. G. Yakovlev. Interval’nye vychisleniia – predmet isledovaniĭ i polesnijĭ instrument. (Russian) [Interval computations — subject of research and useful tool]. Interval Computations = Interval’nye vychisleniia, 1(1):10– 26, 1991. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1991/interval-computations-
1991-1-pp-10-26.pdf.

Yakovlev:1991:ICS

[4] A. G. Yakovlev. Interval computations — subject of research and useful tool. Interval Computations = Interval’nye vychisleniia, 1(1):27–43, 1991.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1991/interval-computations-1991-1-
pp-27-43.pdf.

Herzberger:1991:TAB

[5] J. Herzberger and D. Bethke. On two algorithms for bounding the inverses of an interval matrix. Interval Computations = Interval’nye vychisleniia, 1(1):44–53, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-1-pp-44-53.pdf.

Dobronets:1991:OIR

[6] B. S. Dobronets and V. I. Senashov. Ov intervalnix rasshireniyax nekotorix klassov fynktsiĭ. (Russian) [on interval extensions of some classes of functions]. Interval Computations = Interval’nye vychisleniia, 1(1):

54–58, 1991. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1991/interval-computations-
1991-1-pp-54-58.pdf.

Kearfott:1991:RPI

[7] R. Baker Kearfott, Chen-Yi Hu, and Manuel Novoa III. A review of preconditioners for the interval Gauss–Seidel method. Interval Computations = Interval’nye vychisleniia, 1(1):59–85, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-1-pp-59-85.pdf.

Musaev:1991:NIP

[8] E. A. Musaev. Narrowing of intervals by partial derivatives. Interval Computations = Interval’nye vychisleniia, 1(1):86–91, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-1-pp-86-91.
pdf.

Sharyi:1991:RLZ

[9] S. P. Sharyĭ. O rasreshimosti lineinoi zadachi o dopykskax. (Russian) [on compatibility of linear tolerance problem]. Interval Computations = Interval’nye vychisleniia, 1(1):92–98, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-1-pp-92-98.pdf.

Paluh:1991:PAI

[10] B. V. Paluh, B. V. Vasilyov, and V. L. Perov. Primenenie apparata intervalnoĭ matematiki dlya resheniya zadach texnicheskoĭ diagnostiki nepreryvnovo proisvodstva v ximicheskoĭ promyshlennosti. (Russian) [application of interval mathematics for solving technical diagnostics tasks of non-stop manufacture in chemical industry]. Interval Computations = Interval’nye vychisleniia, 1(1):99–104, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-1-pp-99-104.pdf.

Musaev:1991:IMD

[11] Eldar A. Musaev. Intervalnije metodi dlya chislennix vychisleniĭ. (Russian) [Interval methods for numerical computation]. Interval Computations = Interval’nye vychisleniia, 1(1):108–110, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-1-pp-108-110.
pdf.

Sharyi:1991:VBS

[12] S. P. Sharyĭ. VII Bsesoyuznoe Sobeshchanie po Intervalnoi Matematike. (Russian) [VII All-Union Conference on Interval Mathematics]. Interval Computations = Interval’nye vychisleniia, 1(1):111–112, 1991. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1991/interval-computations-1991-1-
pp-111-112.pdf.

Anonymous:1991:TKOa

[13] Anonymous. Trebovaniya k oformleniyu rykopici. (Russian) [requirements for manuscript preparation]. Interval Computations = Interval’nye vychisleniia, 1(1):113, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-1-p-113.pdf.

Anonymous:1991:RMPa

[14] Anonymous. Requirements for manuscript preparation. Interval Computations = Interval’nye vychisleniia, 1(1):114, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-1-p-114.pdf.

Anonymous:1991:ICIa

[15] Anonymous. Interval computations — interval’nye vychisleniia. Interval Computations = Interval’nye vychisleniia, 1(1):cover, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-1-p-1.pdf.

Nesterov:1991:DCb

[16] V. M. Nesterov. Dear colleagues! Interval Computations = Interval’nye vychisleniia, 1(2):2–3, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-2-3.pdf.

Nesterov:1991:UKRb

[17] V. M. Nesterov. Uvaszaemije kollegi! (Russian) [Dear colleagues]. Interval Computations = Interval’nye vychisleniia, 1(2):4–6, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-pp-4-6.pdf.

Shary:1991:OSI

[18] Sergey P. Shary. Optimal solution of interval linear algebraic systems. I. Interval Computations = Interval’nye vychisleniia, 1(2):7–30, 1991.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1991/interval-computations-1991-2-
pp-7-30.pdf.

Kreinovich:1991:PCE

[19] Vladik Kreinovich, Andrew Bernat, Elsa Villa, and Yvonne Mariscal. Parallel computers. Estimate errors caused by imprecise data. Interval Computations = Interval’nye vychisleniia, 1(2):31–46, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-pp-31-46.
pdf.

Hammer:1991:PXN

[20] R. Hammer, M. Neaga, D. Ratz, and D. Shiryaev. PASCAL–XSC: novyĭ yazik dlya nauchnix vychisleniĭ. (Russian) [PASCAL–XSC: a new language for scientific computing]. Interval Computations = Interval’nye vychisleniia, 1(2):47–81, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-47-81.pdf.

Anonymous:1991:VRR

[21] Anonymous. Vpernye na ryskom razike: polnyĭ perovod standartov ANSI/ IEEE na mashiniyu arifmetiky. (Russian) []. Interval Computations = Interval’nye vychisleniia, 1(2):82, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-p-82.pdf.

Khlebalin:1991:PPA

[22] N. A. Khlebalin. Program package ASIAS: Analysis and synthesis of interval automatic systems. Interval Computations = Interval’nye vychisleniia, 1(2):83, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-83-84.pdf.

Khlebalin:1991:KPA

[23] N. A. Khlebalin. Komplex program ASIAS: Analiz i sintez integralnix avtomaticheskix sistem. (Russian) [program package ASIAS: Analysis and synthesis of interval automatic systems]. Interval Computations = Interval’nye vychisleniia, 1(2):84, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-83-84.pdf.

Pankov:1991:RBD

[24] P. S. Pankov. A review of the book: Dobronets B. S., Shaydurov V. V., Two-sided Numerical Methods, Nauka (Siberian Department), Novosibirsk, 1990, 208 pp. Interval Computations = Interval’nye vychisleniia, 1(2):85–86, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-85-86.pdf.

Zakharov:1991:APAa

[25] A. V. Zakharov. Actual problems of applied mathematics. All-Union conference. Saratov, May 20–22, 1991. A brief report. Interval Computations = Interval’nye vychisleniia, 1(2):87–90, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-87-90.pdf.

Zakharov:1991:APP

[26] A. V. Zakharov. Aktyalnije problemij prikladnoĭ matematiki. bsesoyuznaya konferentsiya. (Russian) [actual problems of applied mathematics. All-Union conference. Saratov, May 20–22, 1991. A brief report]. Interval Computations = Interval’nye vychisleniia, 1(2):91–95, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-pp-91-95.
pdf.

Davidenkoff:1991:ASP

[27] Alexander Davidenkoff. Advanced seminar on programming languages for scientific/engineering computation, Leningrad, June 3– 6, 1991. A summarizing report. Interval Computations = Interval’nye vychisleniia, 1(2):96–106, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-96-106.pdf.

Herzberger:1991:ISS

[28] J. Herzberger. Interval subject on the SCAN-1991 symposium. information. Interval Computations = Interval’nye vychisleniia, 1(2):107, 1991.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1991/interval-computations-1991-2-
p-107.pdf.

Anonymous:1991:RAU

[29] Anonymous. Resolution of 8th All-Union Workshop on interval mathematics. Bishkek, October 1–3, 1991. Interval Computations = Interval’nye vychisleniia, 1(2):108, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-p-108.pdf.

Anonymous:1991:RVB

[30] Anonymous. Reshenie 8-vo bsesoyuznovo soveshchania po intervalnoĭ matematike. (Russian) [resolution of 8th All-Union Workshop on interval mathematics. Bishkek, October 1–3, 1991]. Interval Computations = Interval’nye vychisleniia, 1(2):109, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-p-109.pdf.

Anonymous:1991:PAC

[31] Anonymous. Preliminary announcement, call for papers and further information for a conference on numerical analysis with automatic result verification. Lafayette, Louisiana, February 25–March 1, 1993. Interval Computations = Interval’nye vychisleniia, 1(2):110–114, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-pp-110-114.
pdf.

Yakovlev:1991:BSW

[32] Alexander G. Yakovlev and R. Baker Kearfott. Bibliography of Soviet works on interval computations, Part I. Interval Computations = Interval’nye vychisleniia, 1(2):115–122, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-115-122.pdf.

Anonymous:1991:RMPb

[33] Anonymous. Requirements for manuscript preparation. Interval Computations = Interval’nye vychisleniia, 1(2):123–124, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-pp-123-124.
pdf.

Anonymous:1991:TKOb

[34] Anonymous. Trebovaniya k oformleniyu rykopici. (Russian) [requirements for manuscript preparation]. Interval Computations = Interval’nye vychisleniia, 1(2):125–126, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1991/interval-computations-1991-2-pp-125-126.pdf.

Anonymous:1991:AEBa

[35] Anonymous. Addresses of the Editorial Board Members. Interval Computations = Interval’nye vychisleniia, 1(2):127–131, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-pp-127-131.
pdf.

Anonymous:1991:C

[36] Anonymous. Contents. Interval Computations = Interval’nye vychisleniia, 1(2):132–134, 1991. CODEN ???? ISSN 0135-4868. URL http://

interval.louisiana.edu/reliable-computing-journal/1991/interval-
computations-1991-2-pp-132-134.pdf.

Anonymous:1991:ICIb

[37] Anonymous. Interval computations — interval’nye vychisleniia. Interval Computations = Interval’nye vychisleniia, 1(2):cover, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1991/interval-computations-1991-2-p-1.pdf.

Nesterov:1991:DCc

[38] V. M. Nesterov. Dear colleagues! Interval Computations = Interval’nye vychisleniia, 1(3):2–3, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-2-3.pdf.

Zjuzin:1991:RRR

[39] V. S. Zjuzin. Retsinsija. (Russian) [Review of H. Bauch, K. U. Jahn, D. Oelschägel, H. Süsse, V. Wiebigke, Intervalmathematik (Teorie und Anwendungen), Leipzig, BSB B. G. Teubner Verlagsgesellschaft, 1987, 260 s. (Mathematisch-Naturwissenschaftliche Bibliothek, Band 72)]. Interval Computations = Interval’nye vychisleniia, 1(3):105–107, 1991. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1991/interval-computations-1991-1-

pp-105-107.pdf. D. Oelschlaegel, H. Suesse, V. Wiebigke Intervallmathematik (Teorie und Anwendungen). - Leipzig, BSB B. G. Teubner Verlagsgesellschaft, 1987, 260 S. (Mathematisch-Naturwissenschaftliche Bibliothek, Band 72).

Anonymous:1991:RMPc

[40] Anonymous. Requirements for manuscript preparation. Interval Computations = Interval’nye vychisleniia, 1(3):112–113, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-1-pp-112-113.
pdf.

Anonymous:1991:TKOc

[41] Anonymous. Trebovaniya k oformleniyu rykopici. (Russian) [requirements for manuscript preparation]. Interval Computations = Interval’nye vychisleniia, 1(3):114–115, 1991. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-114-115.pdf.

Anonymous:1991:AEBb

[42] Anonymous. Addresses of the Editorial Board Members. Interval Computations = Interval’nye vychisleniia, 1(3):116–118, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-1-pp-116-118.
pdf.

Anonymous:1991:ICIc

[43] Anonymous. Interval computations — interval’nye vychisleniia. Interval Computations = Interval’nye vychisleniia, 1(3):cover, 1991. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-1-p-1.pdf.

Nesterov:1992:UKRa

[44] V. M. Nesterov. Uvaszaemije kollegi! (Russian) [Dear colleagues]. Interval Computations = Interval’nye vychisleniia, 1(3):4–5, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-1-pp-4-5.pdf.

Dobronets:1992:STS

[45] Boris S. Dobronets. On some two-sided methods for solving systems of ordinary differential equations. Interval Computations = Interval’nye vychisleniia, 1(3):6–21, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-6-21.pdf.

Rump:1992:ISL

[46] Siegfried M. Rump. Inclusion of the solution for large linear systems with M-matrix. Interval Computations = Interval’nye vychisleniia, 1(3): 22–43, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-1-pp-22-43.pdf.

Orlov:1992:IS

[47] Alexander I. Orlov. Interval statistics. Interval Computations = Interval’nye vychisleniia, 1(3):44–52, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-44-52.pdf.

Musaev:1992:WCT

[48] Eldar A. Musaev. Wave computations. A technique for optimal quasi-concurrent self-validation. Interval Computations = Interval’nye vychisleniia, 1(3):53–60, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-53-60.pdf.

Yakovlev:1992:CAP

[49] Alexander G. Yakovlev. Classification approach to programming of localizational (interval) computations. Interval Computations = Interval’nye vychisleniia, 1(3):61–84, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-61-84.pdf.

Pankova:1992:PSA

[50] Galina D. Pankova. Programming support and algorithmic language for validating computations on ES (IBM 360/370) series computer. Interval Computations = Interval’nye vychisleniia, 1(3):86–95, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-1-pp-86-95.
pdf.

Alefeld:1992:RBN

[51] G. Alefeld. A review of the book: A. Neumaier, Interval Methods for Systems of Equations, Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, 1990. Interval Computations = Interval’nye vychisleniia, 1(3):96–97, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-96-97.pdf.

Anonymous:1992:LAR

[52] Anonymous. List of abstracts from “Referativnyi Zhurnal ‘Matematika’ for 1991”. Interval Computations = Interval’nye vychisleniia, 1(3):98– 99, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-1-pp-98-99.pdf.

Kozina:1992:IMA

[53] Galina L. Kozina. Interval mathematics and its applications: Workshop. Interval Computations = Interval’nye vychisleniia, 1(3):100–101, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-1-
pp-100-101.pdf.

Yakovlev:1992:BSWb

[54] Alexander G. Yakovlev and R. Baker Kearfott. Bibliography of Soviet works on interval computations. II. Interval Computations = Interval’nye vychisleniia, 1(3):104–111, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-1-pp-104-111.pdf.

Anonymous:1992:Ca

[55] Anonymous. Contents. Interval Computations = Interval’nye vychisleniia, 1(4):119–121, 1992. CODEN ???? ISSN 0135-4868. URL http://

interval.louisiana.edu/reliable-computing-journal/1992/interval-
computations-1992-1-pp-119-121.pdf.

Sharyi:1992:OSI

[56] Sergey P. Sharyĭ. Optimal solution of interval linear algebraic systems. I. Interval Computations = Interval’nye vychisleniia, 2(2):7–30, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1991/interval-computations-1991-2-
pp-7-30.pdf.

Anonymous:1992:ICIc

[57] Anonymous. Interval computations — interval’nye vychisleniia. Interval Computations = Interval’nye vychisleniia, 2(2(4)):cover, 1992. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-2-
p-1.pdf.

vonGudenberg:1992:DC

[58] J. Wolff von Gudenberg. Dear colleagues! Interval Computations = Interval’nye vychisleniia, 2(3(5)):6, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-p-6.pdf.

vonGudenberg:1992:UKR

[59] J. Wolff von Gudenberg. Uvaszaemije kollegi! (Russian) [Dear colleagues]. Interval Computations = Interval’nye vychisleniia, 2(3(5)):7, 1992. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-
p-7.pdf.

Akunova:1992:CCS

[60] Akylay Akunova, Taalaybek A. Akunov, and Anatoly V. Ushakov. Construction of a comparison system for multi-dimensional control processes with interval state matrix. Interval Computations = Interval’nye vychisleniia, 2(3(5)):8–12, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-pp-8-12.pdf. Interval

’92 (Moscow, 1992).

Babichev:1992:UTS

[61] A. B. Babichev, O. B. Kadyrova, T. P. Kashevarova, and Alexander L. Semenov. UniCalc as a tool for solving problems with inaccurate and sub-definite data. Interval Computations = Interval’nye vychisleniia, 2(3(5)):13–16, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-pp-13-16.pdf. Inter-

val ’92 (Moscow, 1992).

Blomquist:1992:IID

[62] Frithjof Blomquist. Interval inclusions for Dawson’s integral. Interval Computations = Interval’nye vychisleniia, 2(3(5)):17–26, 1992. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-17-26.pdf. Interval ’92 (Moscow, 1992).

Bochkov:1992:AEZ

[63] Alexander F. Bochkov and Lubov A. Yakovleva. Algorithm for experimental zero-order optimization for plant with bounded amplitude errors. Interval Computations = Interval’nye vychisleniia, 2(3(5)):27–30, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-27-30.pdf. Interval ’92 (Moscow, 1992).

Bochkov:1992:IND

[64] Alexandr F. Bochkov and Nguen Viet Zung. Identification of nonlinear dynamic objects using interval experimental data. Interval Computations = Interval’nye vychisleniia, 2(3(5)):31–37, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-

journal/1992/interval-computations-1992-3-pp-31-37.pdf. Inter-

val ’92 (Moscow, 1992).

Claudio:1992:OTA

[65] Dalcidio M. Claudio, Martín H. Escardó, and Beatriz R. T. Franciosi. An order-theoretic approach to interval analysis. Interval Computations = Interval’nye vychisleniia, 2(3(5)):38–45, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-pp-38-45.pdf. Inter-

val ’92 (Moscow, 1992).

Demchenko:1992:STN

[66] A. I. Demchenko, B. V. Pel′tsverger, and O. V. Khavronin. Synthesis of transport networks structures under conditions of uncertain initial information. Interval Computations = Interval’nye vychisleniia, 2(3(5)): 46–49, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-

1992-3-pp-46-49.pdf. Interval ’92 (Moscow, 1992).

Dobronets:1992:IMB

[67] Boris S. Dobronets. Interval methods based on a posteriori estimates. Interval Computations = Interval’nye vychisleniia, 2(3(5)):50–55, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-50-55.pdf. Interval ’92 (Moscow, 1992).

Dugarova:1992:AIM

[68] Irina V. Dugarova. An algorithm of interval matrix asymptotic stability testing. Interval Computations = Interval’nye vychisleniia, 2(3(5)): 56–62, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-

1992-3-pp-56-62.pdf. Interval ’92 (Moscow, 1992).

Ermakov:1992:TSM

[69] Oleg B. Ermakov. Two-sided method for solving system of ordinary differential equations with automatic determination of guaranteed estimates. Interval Computations = Interval’nye vychisleniia, 2(3(5)):63–69, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-63-69.pdf. Interval ’92 (Moscow, 1992).

Evtushenko:1992:OPS

[70] Tatjana V. Evtushenko. Optimization problems for static plants under uncertainty conditions. Interval Computations = Interval’nye vychisleniia, 2(3(5)):70–74, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-pp-70-74.pdf. Inter-

val ’92 (Moscow, 1992).

Filippov:1992:EEE

[71] Alexey F. Filippov. Ellipsoidal error estimates for Adams method. Interval Computations = Interval’nye vychisleniia, 2(3(5)):75–79, 1992. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-75-79.pdf. Interval ’92 (Moscow, 1992).

Glazunov:1992:IEC

[72] Nikolay M. Glazunov. On interval extensions of computer algebra systems. Interval Computations = Interval’nye vychisleniia, 2(3(5)):80–87, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-80-87.pdf. Interval ’92 (Moscow, 1992).

Henriksen:1992:PAG

[73] Tom Henriksen and Kaj Madsen. Parallel algorithms for global optimization. Interval Computations = Interval’nye vychisleniia, 2(3(5)): 88–95, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-

1992-3-pp-88-95.pdf. Interval ’92 (Moscow, 1992).

Kearfott:1992:IPF

[74] R. Baker Kearfott, Milind Dawande, Kaishen Du, and Chenyi Hu. INTLIB: a portable Fortran-77 elementary function library. Interval Computations = Interval’nye vychisleniia, 2(3(5)):96–105, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-3-pp-96-105.

pdf. Interval ’92 (Moscow, 1992).

Khlebalin:1992:IAS

[75] Nikolay A. Khlebalin. Interval automatic systems — theory, computeraided design and applications. Interval Computations = Interval’nye vychisleniia, 2(3(5)):106–115, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-

journal/1992/interval-computations-1992-3-pp-106-115.pdf. In-

terval ’92 (Moscow, 1992).

Korlyukov:1992:NAI

[76] Alexander V. Korlyukov. A new application of interval mathematics. Interval Computations = Interval’nye vychisleniia, 2(3(5)):116–121, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-3-

pp-116-121.pdf. Interval ’92 (Moscow, 1992).

Anonymous:1992:ICId

[77] Anonymous. International Conference on Interval and Computer-Algebraic Methods in Science and Engineering (INTERVAL’94): March 6–11, 1994, St. Petersburg, Russia. Interval Computations = Interval’nye vychisleniia, 2(3(5)):122–123, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-pp-122-123.pdf.

Anonymous:1992:MKP

[78] Anonymous. Mezhdunarodnaya konferentsiya po intervalnym i kompyuternoalgebraicheskim metodam v nayke i texnike. (Russian) []. Interval Computations = Interval’nye vychisleniia, 2(3(5)):124–125, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-3-pp-124-125.
pdf.

Anonymous:1992:Cb

[79] Anonymous. Contents. Interval Computations = Interval’nye vychisleniia, 2(3(5)):126–128, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-3-pp-126-128.pdf.

Anonymous:1992:ICIa

[80] Anonymous. Interval computations — interval’nye vychisleniia: Special issue: Proceedings of the conference “Interval ’92”, Moscow, September 22–25, 1992. Interval Computations = Interval’nye vychisleniia, 2(3(5)): cover, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-1-p-1.pdf.

Nesterov:1992:DC

[81] Vyacheslav M. Nesterov. Dear colleagues! Interval Computations = Interval’nye vychisleniia, 2(4):2–3, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-2-pp-2-3.pdf.

Nesterov:1992:UKRb

[82] V. M. Nesterov. Uvaszaemije kollegi! (Russian) [Dear colleagues]. Interval Computations = Interval’nye vychisleniia, 2(4):4–5, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-2-pp-4-5.pdf.

Filippov:1992:EES

[83] A. F. Filippov. Ellipsoidal estimates for a solution of a system of differential equations. Interval Computations = Interval’nye vychisleniia, 2 (4):6–17, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-2-pp-6-17.pdf.

Sharyi:1992:NCA

[84] Sergey P. Sharyĭ. A new class of algorithms for optimal solution of interval linear systems. Interval Computations = Interval’nye vychisleniia, 2(4):18–29, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-2-pp-18-29.pdf.

Pankov:1992:PFD

[85] P. S. Pankov. Proof of finite-dimensional inequalities by means of interval analysis. Interval Computations = Interval’nye vychisleniia, 2(4): 30–38, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-2-pp-30-38.pdf.

Zyuzin:1992:MIS

[86] Vladimir S. Zyuzin. On metrization of interval sets I(R), I(R^{n}). Interval Computations = Interval’nye vychisleniia, 2(4):39–50, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-2-pp-39-50.
pdf.

Kreinovich:1992:IEC

[87] Vladik Kreinovich, Andrew Bernat, Olga Kosheleva, and Andrei Finkelstein. Interval estimates for closure-phase and closure-amplitude imaging in radio astronomy. Interval Computations = Interval’nye vychisleniia, 2 (4):51–71, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-2-pp-51-71.pdf.

Davidenkoff:1992:AXY

[88] Alexander Davidenkoff. ACRITH–XSC: Yazik programmirovaniya dla nauchno-texnicheckix vychisleniĭ. (Russian) [ACRITH–XSC: a programming language for scientific/engineering computation]. Interval Computations = Interval’nye vychisleniia, 2(4):72–81, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-2-pp-72-81.
pdf.

Menshikov:1992:TEI

[89] G. G. Menshikov. The teaching experience of interval computations at the Department of Applied Mathematics and Control Processes of St. Petersburg University. Interval Computations = Interval’nye vychisleniia, 2(4):83–85, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-2-pp-82-85.pdf.

Anonymous:1992:ICIb

[90] A. P. Voshchinin. International Conference on Interval and Stochastic Methods in Science and Engineering “INTERVAL-92”. Interval Computations = Interval’nye vychisleniia, 2(4):86–91, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-2-pp-86-91.
pdf.

Anonymous:1992:LA

[91] Anonymous. List of abstracts from “Referativnyi Zhurnal ‘Matematika’ for 1991”. Interval Computations = Interval’nye vychisleniia, 2(4):105– 106, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-2-pp-105-106.pdf.

Yakovlev:1992:BSWc

[92] Alexander G. Yakovlev and R. Baker Kearfott. Bibliography of Soviet works on interval computations. III. Interval Computations = Interval’nye vychisleniia, 2(4):107–115, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-2-pp-107-115.pdf.

Anonymous:1992:RMP

[93] Anonymous. Requirements for manuscript preparation. Interval Computations = Interval’nye vychisleniia, 2(4):116–117, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-2-pp-116-117.
pdf.

Anonymous:1992:TKO

[94] Anonymous. Trebovaniya k oformleniyu rykopici. (Russian) [requirements for manuscript preparation]. Interval Computations = Interval’nye vychisleniia, 2(4):118–119, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-2-pp-118-119.pdf.

Anonymous:1992:AEB

[95] Anonymous. Addresses of the Editorial Board Members. Interval Computations = Interval’nye vychisleniia, 2(4):120–122, 1992. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-2-pp-120-122.
pdf.

Anonymous:1992:Cc

[96] Anonymous. Contents. Interval Computations = Interval’nye vychisleniia, 2(4):123–125, 1992. CODEN ???? ISSN 0135-4868. URL http://

interval.louisiana.edu/reliable-computing-journal/1992/interval-
computations-1992-2-pp-123-125.pdf.

Anonymous:1992:FAC

[97] Anonymous. First announcement and call for papers: International Conference on Mathematical Modelling and Scientific Computation (MMSC-93), September 14–17, 1993, Sozopol. Interval Computations = Interval’nye vychisleniia, 2(4), 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-2-pp-99-104.pdf.

Krishchuk:1992:IOF

[98] Vladimir N. Krishchuk, Nikolay M. Vasilega, and Galina L. Kozina. Interval operations and functions library for FORTRAN 77 programming system and its practice using. Interval Computations = Interval’nye vychisleniia, 2(4(6)):2–8, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-2-8.pdf. Interval

’92 (Moscow, 1992).

Krymsky:1992:AAR

[99] Victor G. Krymsky. Algorithmic aims of reliability provision for largescale dynamic systems with interval parameters. Interval Computations = Interval’nye vychisleniia, 2(4(6)):9–13, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-9-13.pdf. Interval

’92 (Moscow, 1992).

Litvinov:1992:EAC

[100] Grigory L. Litvinov. Error auto-correction in rational approximation. Interval Computations = Interval’nye vychisleniia, 2(4(6)):14–18, 1992. CODEN ???? ISSN 0135-4868. Interval ’92 (Moscow, 1992).

Markov:1992:PRM

[101] Svetoslav M. Markov. On the presentation of ranges of monotone functions using interval arithmetic. Interval Computations = Interval’nye vychisleniia, 2(4(6)):19–31, 1992. CODEN ???? ISSN 0135-4868. See correction [175]. Interval ’92 (Moscow, 1992).

Menshikov:1992:ICI

[102] Grigory G. Menshikov. Interval co-integration of differential equations connected by a substitution of the variable. Interval Computations = Interval’nye vychisleniia, 2(4(6)):32–36, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-32-36.pdf. Inter-

val ’92 (Moscow, 1992).

Musaev:1992:ARC

[103] Eldar A. Musaev. An approach to reliable computations with the minimal representation. Interval Computations = Interval’nye vychisleniia, 2(4(6)):37–41, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-37-41.pdf. Inter-

val ’92 (Moscow, 1992).

Narinyani:1992:FNP

[104] Alexander S. Narin′yani. NE-factors and natural pragmatics: what do the intervals represent? Interval Computations = Interval’nye vychisleniia, 2(4(6)):42–46, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-42-46.pdf. Inter-

val ’92 (Moscow, 1992).

Nesterov:1992:ERV

[105] Vyacheslav M. Nesterov. Estimating a range of values of functions using extended interval arithmetics. Interval Computations = Interval’nye vychisleniia, 2(4(6)):47–53, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-47-53.pdf. Inter-

val ’92 (Moscow, 1992).

Pankov:1992:UIM

[106] Pavel S. Pankov and Batyigul D. Bayachorova. Using interval methods in cluster analysis and verified representation of connected sets. Interval Computations = Interval’nye vychisleniia, 2(4(6)):54–58, 1992. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-4-

pp-54-58.pdf. Interval ’92 (Moscow, 1992).

Senio:1992:SSS

[107] Pyotr S. Senio and Pyotr S. Vengersky. Solving systems of special form nonlinear equations by means of some modifications of Runge type interval iterative method. Interval Computations = Interval’nye vychisleniia, 2(4(6)):59–65, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-59-65.pdf. Inter-

val ’92 (Moscow, 1992).

Sharyi:1992:CSS

[108] Sergey P. Sharyĭ. On controlled solution set of interval algebraic systems. Interval Computations = Interval’nye vychisleniia, 2(4(6)):66–75, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-4-

pp-66-75.pdf. Interval ’92 (Moscow, 1992).

Shiriaev:1992:PXP

[109] Dmitri Shiriaev. PASCAL–XSC. A portable programming system for scientific computations. Interval Computations = Interval’nye vychisleniia, 2(4(6)):76–82, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-76-82.pdf. Inter-

val ’92 (Moscow, 1992).

Simoff:1992:IAR

[110] Simeon J. Simoff. Interval approximate reasoning for expert systems. Interval Computations = Interval’nye vychisleniia, 2(4(6)):83–87, 1992.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1992/interval-computations-1992-4-

pp-83-87.pdf. Interval ’92 (Moscow, 1992).

Skybytsky:1992:CLD

[111] Nikita V. Skybytsky and Tian Yuping. Control of the linear dynamic plant with intervally given parameters from the guarantee condition of the required accuracy of the solution. Interval Computations = Interval’nye vychisleniia, 2(4(6)):88–93, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-88-93.pdf. Inter-

val ’92 (Moscow, 1992).

Smagina:1992:GPA

[112] Elena M. Smagina. General problem of the asymptotic steady-output tracking for plant with interval parameters. Interval Computations = Interval’nye vychisleniia, 2(4(6)):94–99, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-94-99.pdf. Inter-

val ’92 (Moscow, 1992).

Ten:1992:SOC

[113] Iosif G. Ten. Synthesis of optimal control under interval uncertainty in models. Interval Computations = Interval’nye vychisleniia, 2(4(6)):100–106, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-100-106.pdf. In-

terval ’92 (Moscow, 1992).

Voshchinin:1992:SQA

[114] A. P. Voshchinin. Some questions of application of interval mathematics in parameter estimation and decision making. Interval Computations = Interval’nye vychisleniia, 2(4(6)):107–115, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-107-115.pdf. In-

terval ’92 (Moscow, 1992).

vonGudenberg:1992:PLS

[115] Jürgen Wolff von Gudenberg. Programming language support for scientific computation. Interval Computations = Interval’nye vychisleniia, 2(4(6)):116–126, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-116-126.pdf. In-

terval ’92 (Moscow, 1992).

Zyuzin:1992:EFD

[116] Vladimir S. Zyuzin. The extension of the Fréchet derivative concept in the interval-segment analysis. Interval Computations = Interval’nye vychisleniia, 2(4(6)):127–132, 1992. CODEN ???? ISSN 0135-4868. Interval ’92 (Moscow, 1992).

Anonymous:1992:Cd

[117] Anonymous. Contents. Interval Computations = Interval’nye vychisleniia, 2(4(6)):133–136, 1992. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1992/interval-computations-1992-4-pp-133-136.pdf.

Anonymous:1992:ICIe

[118] Anonymous. Interval computations — interval’nye vychisleniia: Special issue: Proceedings of the conference “Interval ’92”, Moscow, September 22–25, 1992. Interval Computations = Interval’nye vychisleniia, 2(4(6)): cover, 1992. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1992/interval-computations-
1992-2-p-1.pdf.

Hansen:1993:CZF

[119] Eldon R. Hansen. Computing zeros of functions using generalized interval arithmetic. Interval Computations = Interval’nye vychisleniia, 2 (3(5)):3–28, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Hu:1993:BRS

[120] Chen-Yi Hu, R. Baker Kearfott, and Abdulhamid Awad. On bounding the range of some elementary functions in FORTRAN-77. Interval Computations = Interval’nye vychisleniia, 2(3(5)):29–39, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

VanIwaarden:1993:ADA

[121] Ronald Van Iwaarden. Automatic differentiation applied to unconstrained nonlinear optimization with result verification. Interval Computations = Interval’nye vychisleniia, 2(3(5)):41–60, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Jaulin:1993:GNP

[122] Luc Jaulin and Eric Walter. Guaranteed nonlinear parameter estimation via interval computations. Interval Computations = Interval’nye vychisleniia, 2(3(5)):61–75, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Keiper:1993:IAM

[123] Jerry B. Keiper. Interval arithmetic in Mathematica. Interval Computations = Interval’nye vychisleniia, 2(3(5)):76–87, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Kohout:1993:IVI

[124] Ladislav J. Kohout and Isabel Stabile. Interval-valued inference in medical knowledge-based system CLINAID. Interval Computations = Interval’nye vychisleniia, 2(3(5)):88–115, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Korn:1993:VSL

[125] Carlos Falcó Korn and Christian P. Ullrich. Verified solution of linear systems based on common software libraries. Interval Computations = Interval’nye vychisleniia, 2(3(5)):116–132, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Kristinsdottir:1993:MTI

[126] Birna P. Kristinsdottir, Zelda B. Zabinsky, Tibor Csendes, and Mark E. Tuttle. Methodologies for tolerance intervals. Interval Computations = Interval’nye vychisleniia, 2(3(5)):133–147, ???? 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Leclerc:1993:PIG

[127] Anthony Leclerc. Parallel interval global optimization and its implementation in C++. Interval Computations = Interval’nye vychisleniia, 2(3(5)): 148–163, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Markov:1993:SIP

[128] Svetoslav M. Markov. Some interpolation problems involving interval data. Interval Computations = Interval’nye vychisleniia, 2(3(5)):164–182, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Nakamura:1993:GIK

[129] Mitsumi Nakamura, Ray Mines, and Vladik Kreinovich. Guaranteed intervals for Kolmogorov’s Theorem (and their possible relation to neural networks). Interval Computations = Interval’nye vychisleniia, 2(3(5)): 183–199, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Alvarado:1993:DSI

[130] Fernando L. Alvarado and Zian Wang. Direct sparse interval hull computations for thin non-M-matrices. Interval Computations = Interval’nye vychisleniia, 2(4):5–28, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Babichev:1993:UNA

[131] A. B. Babichev, O. B. Kadyrova, T. P. Kashevarova, A. S. Leshchenko, and Alexander L. Semenov. UniCalc, a novel approach to solving systems of algebraic equations. Interval Computations = Interval’nye vychisleniia, 2(4):29–47, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Berleant:1993:AVR

[132] Daniel Berleant. Automatically verified reasoning with both intervals and probability density functions. Interval Computations = Interval’nye vychisleniia, 2(4):48–70, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Caprani:1993:URV

[133] Ole Caprani, Brian Godthaab, and Kaj Madsen. Use of a real-valued local minimum in parallel interval global optimization. Interval Computations = Interval’nye vychisleniia, 2(4):71–82, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Cervesato:1993:NMB

[134] Iliano Cervesato, Angelo Montanari, and Alessandro Provetti. On the nonmonotonic behaviour of event calculus for deriving maximal time intervals. Interval Computations = Interval’nye vychisleniia, 2(4):83–119, 1993. CO-DEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Connell:1993:EIA

[135] Amanda E. Connell and Robert. M. Corless. An experimental interval arithmetic package in Maple. Interval Computations = Interval’nye vychisleniia, 2(4):120–134, 1993. CODEN ???? ISSN 0135-4868.

Ely:1993:VSP

[136] J. S. Ely. The VPI software package for variable precision interval arithmetic. Interval Computations = Interval’nye vychisleniia, 2(4):135–154, 1993. CODEN ???? ISSN 0135-4868.

Garloff:1993:BA

[137] Jürgen Garloff. The Bernstein algorithm. Interval Computations = Interval’nye vychisleniia, 2(4):154–168, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Hager:1993:SLS

[138] Gregory D. Hager. Solving large systems of nonlinear constraints with application to data modeling. Interval Computations = Interval’nye vychisleniia, 2(4):169–200, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Nesterov:1993:HUM

[139] Vyacheslav M. Nesterov. How to use monotonicity-type information to get better estimates of the range of real-valued functions. Interval Computations = Interval’nye vychisleniia, 2(4(6)):3–12, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Rohn:1993:CTB

[140] Jiří Rohn. Cheap and tight bounds: The recent result by E. Hansen can be made more efficient. Interval Computations = Interval’nye vychisleniia, 2(4(6)):13–21, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Schaefer:1993:PZA

[141] Mark J. Schaefer. Precise zeros of analytic functions using interval arithmetic. Interval Computations = Interval’nye vychisleniia, 2(4(6)):22–39, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Schnepper:1993:API

[142] Carol A. Schnepper and Mark A. Stadtherr. Application of a parallel interval Newton/generalized bisection algorithm to equation-based chemical process flowsheeting. Interval Computations = Interval’nye vychisleniia, 2(4(6)):40–64, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Schulte:1993:PHD

[143] Michael J. Schulte and Earl E. Swartzlander, Jr. Parallel hardware designs for correctly rounded elementary functions. Interval Computations = Interval’nye vychisleniia, 2(4(6)):65–88, ???? 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Simcik:1993:BBI

[144] L. Simcik and P. Linz. Boundary-based interval Newton’s method. Interval Computations = Interval’nye vychisleniia, 2(4(6)):89–99, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA,

Sirisaengtaksin:1993:NNS

[145] Ongard Sirisaengtaksin and Vladik Kreinovich. Neural networks that are not sensitive to the imprecision of hardware neurons. Interval Computations = Interval’nye vychisleniia, 2(4(6)):100–113, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Storck:1993:VCN

[146] Ulrike Storck. Verified calculation of the nodes and weights for Gaussian quadrature formulas. Interval Computations = Interval’nye vychisleniia, 2(4(6)):114–124, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Turksen:1993:IVF

[147] I. Burhan Türksen. Interval valued fuzzy sets and fuzzy connectives. Interval Computations = Interval’nye vychisleniia, 2(4(6)):125–142, 1993. CO-DEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Veliov:1993:CIU

[148] V. M. Veliov. Computation of integrals of uncertain vector functions. Interval Computations = Interval’nye vychisleniia, 2(4(6)):143–153, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Villa:1993:EEI

[149] Elsa Villa, Andrew Bernat, and Vladik Kreinovich. Estimating errors of indirect measurement on realistic parallel machines: routings on 2-D and 3-D meshes that are nearly optimal. Interval Computations = Interval’nye vychisleniia, 2(4(6)):154–175, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Villaverde:1993:LTA

[150] Karen Villaverde and Vladik Kreinovich. A linear-time algorithm that locates local extrema of a function of one variable from interval measurement results. Interval Computations = Interval’nye vychisleniia, 2 (4(6)):176–194, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Yakovlev:1993:ML

[151] Alexander G. Yakovlev. Multiaspectness and localization. Interval Computations = Interval’nye vychisleniia, 2(4(6)):195–209, 1993. CODEN ???? ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Nesterov:1993:DC

[152] V. M. Nesterov. Dear colleagues! Interval Computations = Interval’nye vychisleniia, 3(1):2–3, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-2-3.pdf.

Nesterov:1993:UKR

[153] V. M. Nesterov. Uvaszaemije kollegi! (Russian) [Dear colleagues]. Interval Computations = Interval’nye vychisleniia, 3(1):4–5, 1993. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1993/interval-computations-1993-1-pp-4-5.pdf.

Kreinovich:1993:OSI

[154] Vladik Kreinovich, Anatoly V. Lakeyev, and Sergey I. Noskov. Optimal solution of interval linear systems is intractable (NP-hard). Interval Computations = Interval’nye vychisleniia, 3(1):6–14, 1993. CODEN ???? ISSN 0135-4868.

Kearfott:1993:PSH

[155] R. Baker Kearfott and Xiaofa Shi. A preconditioner selection heuristic for efficient iteration with decomposition of arithmetic expressions for nonlinear algebraic systems. Interval Computations = Interval’nye vychisleniia, 3(1):15–33, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-15-33.pdf.

Petkovic:1993:SOI

[156] Ljiljana D. Petković and Miroslav Trajković. On some optimal inclusion approximations by disks. Interval Computations = Interval’nye vychisleniia, 3(1):34–50, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-34-50.pdf.

Perepelitsa:1993:IDM

[157] Vitaly A. Perepelitsa and Galina L. Kozina. Interval discrete models and multiobjectivity. Complexity estimates. Interval Computations = Interval’nye vychisleniia, 3(1):51–59, 1993. CODEN ???? ISSN 0135-4868.

Akyildiz:1993:NPI

[158] Yılmaz Akyıldiz and Mohammad I. Al-Suwaiyel. No pathologies for interval Newton’s method. Interval Computations = Interval’nye vychisleniia, 3(1):60–72, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-60-72.pdf.

Anonymous:1993:SAC

[159] Anonymous. Second announcement and call for papers: International Conference on Interval and Computer-Algebraic Methods in Science and Engineering (INTERVAL’94), March 7-10, 1994, St. Petersburg, Russia. Interval Computations = Interval’nye vychisleniia, 3(1):73–77, 1993.

CODEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1993/interval-computations-1993-1-
pp-73-77.pdf.

Khlebalin:1993:RBL

[160] N. A. Khlebalin. Review of the book: Ludyk, G., CAE von Dynamischen Systemen. Analyse, Simulation, Entwurf von Regelungssystemen, Springer-Verlag, Berlin–Heidelberg, 1990, 335 p. Interval Computations = Interval’nye vychisleniia, 3(1):78–82, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-78-82.pdf.

Khlebalin:1993:RRR

[161] N. A. Khlebalin. Retsinsija. (Russian) [review of the book Ludyk, G., CAE von Dynamischen Systemen. Analyse, Simulation, Entwurf von Regelungssystem, Springer-Verlag, Berlin–Heidelberg, 1990, 335 p.]. Interval Computations = Interval’nye vychisleniia, 3(1):83–87, 1993. CO-

DEN ???? ISSN 0135-4868. URL http://interval.louisiana.edu/
reliable-computing-journal/1993/interval-computations-1993-1-
pp-83-87.pdf.

Kolev:1993:UIS

[162] L. V. Kolev. Use of interval slopes for the irrational part of factorable functions. Interval Computations = Interval’nye vychisleniia, 3(1):83–93, 1993. CODEN ???? ISSN 0135-4868.

Anonymous:1993:ICC

[163] Anonymous. International Congress on Computer Systems and Applied Mathematics (CSAM-93). Interval Computations = Interval’nye vychisleniia, 3(1):88–90, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-88-90.pdf.

Anonymous:1993:MKP

[164] Anonymous. Mezhdunarodnyĭ Kongress po kompyuternim sistemam i prikladnoĭ matematike (CSAM-93). (Russian) [international congress on computer systems and applied mathematics (CSAM-93)]. Interval Computations = Interval’nye vychisleniia, 3(1):91–93, 1993. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1993/interval-computations-1993-1-pp-91-93.
pdf.

Herzberger:1993:RIG

[165] J. Herzberger. Report on the IMACS–GAMM International Workshop on Validated Computation, August 30–September 3, 1993, Universität Oldenburg, Germany. Interval Computations = Interval’nye vychisleniia, 3(1):94, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-p-94.pdf.

Shary:1993:ICM

[166] Sergey Shary. International Conference on Mathematical Modelling and Scientific Computations (MMSC0-93), September 14–17, 1993, Sozopol, Bulgaria. Interval Computations = Interval’nye vychisleniia, 3(1):95– 96, 1993. CODEN ???? ISSN 0135-4868. URL http://interval.

louisiana.edu/reliable-computing-journal/1993/interval-computations-
1993-1-pp-95-96.pdf.

Sharyi:1993:MKP

[167] Sergey Sharyĭ. Mezhdunarodnaya Konferentsiya po Matematicheskomy Modelirovanioyu i Naychnym Vychisleniiam (MMSC-93), Sozopol, Bulgariya, 14–17 sentyabrya 1993 goda. Interval Computations = Interval’nye vychisleniia, 3(1):97–98, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-97-98.pdf.

Anonymous:1993:ISS

[168] Anonymous. International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics “SCAN-93”, September 26–29, 1993, Vienna, Austria. Interval Computations = Interval’nye vychisleniia, 3(1):99–101, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-99-101.pdf.

Kreinovich:1993:CPA

[169] Vladik Kreinovich. Call for papers: Announcing special student issue of the international journal Interval Computations. Interval Computations = Interval’nye vychisleniia, 3(1):102, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-p-102.pdf.

Yakovlev:1993:BSW

[170] Alexander G. Yakovlev and R. Baker Kearfott. Bibliography of Soviet works on interval computations. IV. Interval Computations = Interval’nye vychisleniia, 3(1):103–115, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-pp-103-115.pdf. Text

in English and Russian.

Anonymous:1993:RMP

[171] Anonymous. Requirements for manuscript preparation. Interval Computations = Interval’nye vychisleniia, 3(1):116, 1993. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1993/interval-computations-1993-1-p-116.pdf.

Anonymous:1993:TKO

[172] Anonymous. Trebovaniya k oformleniyu rykopici. (Russian) [requirements for manuscript preparation]. Interval Computations = Interval’nye vychisleniia, 3(1):117, 1993. CODEN ???? ISSN 0135-

4868. URL http://interval.louisiana.edu/reliable-computing-
journal/1993/interval-computations-1993-1-p-117.pdf.

Anonymous:1993:AEB

[173] Anonymous. Addresses of the Editorial Board Members. Interval Computations = Interval’nye vychisleniia, 3(1):118–119, 1993. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1993/interval-computations-1993-1-pp-118-119.
pdf.

Anonymous:1993:C

[174] Anonymous. Contents. Interval Computations = Interval’nye vychisleniia, 3(1):120–122, 1993. CODEN ???? ISSN 0135-4868. URL http://

interval.louisiana.edu/reliable-computing-journal/1993/interval-
computations-1993-1-pp-120-122.pdf.

Markov:1993:CPR

[175] Svetoslav M. Markov. Correction to: “On the presentation of ranges of monotone functions using interval arithmetic” [Interval Comput./Interval. Vychisl. 1992, no. 4, 19–31; see MR1253135 (94g:65004)]. Interval Computations = Interval’nye vychisleniia, 3:122–123, 1993. CODEN ????

ISSN 0135-4868. URL http://interval.louisiana.edu/reliable-
computing-journal/1992/interval-computations-1992-3-pp-122-123.

pdf. See [101].

Alefeld:1994:CAE

[176] Götz Alefeld and Günter Mayer. A computer aided existence and uniqueness proof for an inverse matrix eigenvalue problem. Interval Computations = Interval’nye vychisleniia, 1(1):4–27, 1994. CODEN ???? ISSN 0135-4868.

Luther:1994:VIE

[177] Wolfram J. Luther and Werner Otten. Verified inclusion for eigenvalues of Hill’s Equation. Interval Computations = Interval’nye vychisleniia, 1 (1):28–41, 1994. CODEN ???? ISSN 0135-4868.

Kozina:1994:IST

[178] Galina L. Kozina and Vitaly A. Perepelitsa. Interval spanning trees problem: Solvability and computational complexity. Interval Computations = Interval’nye vychisleniia, 1(1):42–50, 1994. CODEN ???? ISSN 0135-4868.

Schwandt:1994:SEI

[179] Hartmut Schwandt. Suboptimal enclosures for the interval Buneman algorithm for arbitrary block dimension. Interval Computations = Interval’nye vychisleniia, 1(1):51–89, 1994. CODEN ???? ISSN 0135-4868.

Ermakov:1994:SSO

[180] Oleg B. Ermakov. Solving systems of ordinary differential equations using Adams’ interpolation method with guaranteed accuracy. Interval Computations = Interval’nye vychisleniia, 1(1):90–95, 1994. CODEN ???? ISSN 0135-4868.

Menshikov:1994:DDI

[181] G. G. Menshikov. On different definitions of interval extension: Problems of teaching. Interval Computations = Interval’nye vychisleniia, 1(1):96– 98, 1994. CODEN ???? ISSN 0135-4868.

Nesterov:1994:BSWa

[182] V. M. Nesterov. Bibliography of Soviet works on interval computations. V. Interval Computations = Interval’nye vychisleniia, 1:100–109, 1994. CODEN ???? ISSN 0135-4868.

Nesterov:1994:BSWb

[183] V. M. Nesterov. Bibliography of Soviet works on interval computations. VI. Interval Computations = Interval’nye vychisleniia, 2(1):116–126, 1994. CODEN ???? ISSN 0135-4868.

Hager:1994:SLS

[184] Gregory D. Hager. Solving large systems of non-linear constraints with application to data modeling. Interval Computations = Interval’nye vychisleniia, 2(1):169–200, ???? 1994. CODEN ???? ISSN 0135-4868.

Kreinovich:1994:PAI

[185] Vladik Kreinovich and Andrew Bernat. Parallel algorithms for interval computations: An introduction. Interval Computations = Interval’nye vychisleniia, 2(3(5)):6–62, 1994. CODEN ???? ISSN 0135-4868.

Lyager:1994:FLE

[186] Erlang Lyager. Finding local extremal points by using parallel interval methods. Interval Computations = Interval’nye vychisleniia, 2(3(5)):63– 80, 1994. CODEN ???? ISSN 0135-4868.

Madsen:1994:PML

[187] Kaj Madsen and Ole Toft. A parallel method for linear interval equations. Interval Computations = Interval’nye vychisleniia, 2(3(5)):81–105, 1994. CODEN ???? ISSN 0135-4868.

Plum:1994:ESP

[188] Michael Plum. Enclosures for solutions of parameter-dependent nonlinear elliptic boundary value problems: Theory and implementation on a parallel computer. Interval Computations = Interval’nye vychisleniia, 2(3(5)): 106–121, 1994. CODEN ???? ISSN 0135-4868.

Sharyi:1994:STP

[189] Sergey P. Sharyĭ. Solving the tolerance problem for interval linear systems. Interval Computations = Interval’nye vychisleniia, 2(4):6–26, 1994. CODEN ???? ISSN 0135-4868.

Dimitrova:1994:VNT

[190] Neli S. Dimitrova and Svetoslav M. Markov. On validated Newton type method for nonlinear equations. Interval Computations = Interval’nye vychisleniia, 2(4):27–51, 1994. CODEN ???? ISSN 0135-4868.

Podchukayev:1994:AMC

[191] Vladimir A. Podchukayev and Igor M. Svetlov. An analytical method of constructing Hurwitz interval polynomials. Interval Computations = Interval’nye vychisleniia, 2(4):52–67, 1994. CODEN ???? ISSN 0135-4868.

Berz:1994:EBL

[192] Martin Berz and Georg Hoffstätter. Exact bounds on the Long term stability of weakly nonlinear systems applied to the design of large storage rings. Interval Computations = Interval’nye vychisleniia, 2(4):68–89, 1994. CODEN ???? ISSN 0135-4868.

Kreinovich:1994:IAT

[193] V. Kreinovich, T. Swenson, and A. Elentukh. Interval approach to testing software. Interval Computations = Interval’nye vychisleniia, 2(4):90–109, 1994. CODEN ???? ISSN 0135-4868.

Christiansen:1994:IMC

[194] Søren Christiansen. Interval methods and condition numbers of linear algebraic systems. Interval Computations = Interval’nye vychisleniia, 2 (4(6)):5–27, 1994. CODEN ???? ISSN 0135-4868. SCAN-93 (Vienna, 1993).

Daumas:1994:RFP

[195] Marc Daumas and David W. Matula. Rounding of floating point intervals. Interval Computations = Interval’nye vychisleniia, 2(4(6)):28–45, 1994. CODEN ???? ISSN 0135-4868. SCAN-93 (Vienna, 1993).

Daumas:1994:TUT

[196] Marc Daumas, Christophe Mazenc, and Jean-Michel Muller. Towards a user transparent interval arithmetic. Interval Computations = Interval’nye vychisleniia, 2(4(6)):46–63, 1994. CODEN ???? ISSN 0135-4868. SCAN-93 (Vienna, 1993).

Nakao:1994:NVS

[197] Mitsuhiro T. Nakao. Numerical verifications of solutions for nonlinear hyperbolic equations. Interval Computations = Interval’nye vychisleniia, 2(4(6)):64–77, 1994. CODEN ???? ISSN 0135-4868. SCAN-93 (Vienna, 1993).

Luther:1994:CSI

[198] Wolfram Luther and Werner Otten. Computation of standard interval functions in multiple-precision interval arithmetic. Interval Computations = Interval’nye vychisleniia, 2(4(6)):78–99, 1994. CODEN ???? ISSN 0135-4868. SCAN-93 (Vienna, 1993).

Popova:1994:EIA

[199] Evgenija D. Popova. Extended interval arithmetic in IEEE floating-point environment. Interval Computations = Interval’nye vychisleniia, 2(4(6)): 100–129, 1994. CODEN ???? ISSN 0135-4868. SCAN-93 (Vienna, 1993).

Kearfott:1995:P

[200] R. B. Kearfott, E. A. Musaev, V. M. Nesterov, and A. G. Yakovlev. Preface. Reliable Computing = Nadezhnye vychisleniia, 1(1):3–4, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1995/reliable-computing-1995-1-pp-3-4.pdf;   http://
link.springer.com/accesspage/article/10.1007/BF02390516;   http:
//link.springer.com/article/10.1007/BF02390516;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
1&issue=1&spage=3-4.

Kirfott:1995:PRF

[201] R. B. Kirfott, Je. A. Musaev, V. M. Nesmerov, and A. G. Jakovlev. Predislovie. (Russian) [Foreword]. Reliable Computing = Nadezhnye vychisleniia, 1(1):5–7, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
1-pp-5-7.pdf;   http://link.springer.com/article/10.1007/BF02390517;
 http://link.springer.com/article/10.1007/BF02390517/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=1&spage=5-7.

Ratschek:1995:FWI

[202] Helmut Ratschek and Jon G. Rokne. Formulas for the width of interval products. Reliable Computing = Nadezhnye vychisleniia, 1(1):9–14, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1995/reliable-computing-1995-1-pp-9-14.pdf;   http:/
/link.springer.com/article/10.1007/BF02390518;   http://link.

springer.com/article/10.1007/BF02390518/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=1&issue=
1&spage=9-14.

Kupriyanova:1995:IEU

[203] Ludmila Kupriyanova. Inner estimation of the united solution set of interval linear algebraic system. Reliable Computing = Nadezhnye vychisleniia, 1(1):15–31, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
1-pp-15-31.pdf;   http://link.springer.com/article/10.1007/BF02390519;
 http://link.springer.com/article/10.1007/BF02390519/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=1&spage=15-31.

Kreinovich:1995:WIS

[204] Vladik Kreinovich. Why intervals? A simple limit theorem that is similar to limit theorems from statistics. Reliable Computing = Nadezhnye vychisleniia, 1(1):33–40, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
1-pp-33-40.pdf;   http://link.springer.com/article/10.1007/BF02390520;
 http://link.springer.com/article/10.1007/BF02390520/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=1&spage=33-40.

Jager:1995:CME

[205] Christine Jäger and Dietmar Ratz. A combined method for enclosing all solutions of nonlinear systems of polynomial equations. Reliable Computing = Nadezhnye vychisleniia, 1(1):41–64, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1995/reliable-computing-
1995-1-pp-41-64.pdf;   http://link.springer.com/article/10.1007/
BF02390521;   http://link.springer.com/article/10.1007/BF02390521/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=1&spage=41-64.

Cooke:1995:IIH

[206] Daniel E. Cooke. An informal introduction to a high level language with applications to interval mathematics. Reliable Computing = Nadezhnye vychisleniia, 1(1):65–75, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-

1-pp-65-75.pdf;   http://link.springer.com/article/10.1007/BF02390522;
 http://link.springer.com/article/10.1007/BF02390522/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=1&spage=65-75.

Eriksson:1995:PIM

[207] Jerry Eriksson and Per Lindström. A parallel interval method implementation for global optimization using dynamic load balancing. Reliable Computing = Nadezhnye vychisleniia, 1(1):77–91, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1995/reliable-computing-1995-1-pp-77-91.pdf;   http:
//link.springer.com/article/10.1007/BF02390523;   http://link.
springer.com/article/10.1007/BF02390523/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=1&issue=
1&spage=77-91.

Kreinovich:1995:ISN

[208] Vladik Kreinovich and Hung T. Nguyen. Interval sessions at NAFIPS/ IFIS/NASA’94. Reliable Computing = Nadezhnye vychisleniia, 1 (1):93–98, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
1-pp-93-98.pdf;   http://link.springer.com/article/10.1007/BF02390524;
 http://link.springer.com/article/10.1007/BF02390524/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=1&spage=93-98.

Nesterov:1995:DCa

[209] V. M. Nesterov. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 1(2):107, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.edu/

reliable-computing-journal/1995/reliable-computing-1995-2-p-
107.pdf;   http://link.springer.com/accesspage/article/10.1007/
BF02384049;   http://link.springer.com/article/10.1007/BF02384049;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=2&spage=107.

Nesterov:1995:UKRa

[210] V. M. Nesterov. Uvazhaemye kollegi. (Russian) [Dear colleagues]. Reliable Computing = Nadezhnye vychisleniia, 1(2):108, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1995/reliable-

computing-1995-2-p-108.pdf;   http://link.springer.com/accesspage/
article/10.1007/BF02384050;   http://link.springer.com/article/
10.1007/BF02384050;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=1&issue=2&spage=108.

Kohout:1995:PIB

[211] Ladislav J. Kohout, Isabel Stabile, Hasan Kalantar, Maria F. San-Andres, and John Anderson. Parallel interval-based reasoning in medical knowledge-based system CLINAID. Reliable Computing = Nadezhnye vychisleniia, 1(2):109–140, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
2-pp-109-140.pdf;   http://link.springer.com/article/10.1007/
BF02384051;   http://link.springer.com/article/10.1007/BF02384051/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=2&spage=109-140.

Kreinovich:1995:AIC

[212] Vladik Kreinovich, David Nemir, and Efren Gutierrez. Applications of interval computations to earthquake-resistant engineering: How to compute derivatives of interval functions fast. Reliable Computing = Nadezhnye vychisleniia, 1(2):141–172, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
2-pp-141-172.pdf;   http://link.springer.com/article/10.1007/
BF02384052;   http://link.springer.com/article/10.1007/BF02384052/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=2&spage=141-172.

Ullrich:1995:RLA

[213] Christian P. Ullrich and Roman Reith. A reliable linear algebra library for transputer networks. Reliable Computing = Nadezhnye vychisleniia, 1(2): 173–187, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-2-pp-173-187.
pdf;   http://link.springer.com/article/10.1007/BF02384053;
http://link.springer.com/article/10.1007/BF02384053/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=2&spage=173-187.

vonGudenberg:1995:PAL

[214] Jürgen Wolff von Gudenberg. Parallel accurate linear algebra subroutines. Reliable Computing = Nadezhnye vychisleniia, 1(2):189– 199, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-2-pp-189-199.
pdf;   http://link.springer.com/article/10.1007/BF02384054;
http://link.springer.com/article/10.1007/BF02384054/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=2&spage=189-199.

Anonymous:1995:AEBa

[215] Anonymous. Addresses of the Editorial Board Members. Reliable Computing = Nadezhnye vychisleniia, 1(2):201–202, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http:

//link.springer.com/accesspage/article/10.1007/BF02384055;
http://link.springer.com/article/10.1007/BF02384055;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=2&spage=201-202.

Kreinovich:1995:TFI

[216] V. Kreinovich and G. Mayer. Towards the future of interval computations. Reliable Computing = Nadezhnye vychisleniia, 1(3):209– 214, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-209-214.
pdf;   http://link.springer.com/article/10.1007/BF02385252;
http://link.springer.com/article/10.1007/BF02385252/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=209-214.

Balaji:1995:AIN

[217] Gopalan V. Balaji and J. D. Seader. Application of interval Newton’s method to chemical engineering problems. Reliable Computing = Nadezhnye vychisleniia, 1(3):215–223, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
3-pp-215-223.pdf;   http://link.springer.com/article/10.1007/
BF02385253;   http://link.springer.com/article/10.1007/BF02385253/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=3&spage=215-223.

Friesen:1995:AGD

[218] Bo. H. Friesen and Vladik Kreinovich. Ockham’s Razor in interval identification. Reliable Computing = Nadezhnye vychisleniia, 1(3):225– 237, ???? 1995. CODEN rcomf8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-225-237.
pdf;   http://link.springer.com/article/10.1007/bf02385254;
http://link.springer.com/article/10.1007/bf02385254/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=225-237.

Herlocker:1995:AGD

[219] Jonathan Herlocker and Jeffrey Ely. An automatic and guaranteed determination of the number of roots of an analytic function interior to a simple closed curve in the complex plane. Reliable Computing = Nadezhnye vychisleniia, 1(3):239–249, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
3-pp-239-249.pdf;   http://link.springer.com/article/10.1007/
BF02385255;   http://link.springer.com/article/10.1007/BF02385255/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=3&spage=239-249.

Hu:1995:GIS

[220] Chenyi Hu, Anna Frolov, R. Baker Kearfott, and Qing Yang. A general iterative sparse linear solver and its parallelization for interval Newton methods. Reliable Computing = Nadezhnye vychisleniia, 1(3):251– 263, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-251-263.
pdf;   http://link.springer.com/article/10.1007/BF02385256;
http://link.springer.com/article/10.1007/BF02385256/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=251-263.

Hu:1995:OIC

[221] Chenyi Hu, Joe Sheldon, R. Baker Kearfott, and Qing Yang. Optimizing INTBIS on the CRAY Y-MP. Reliable Computing = Nadezhnye vychisleniia, 1(3):265–274, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
3-pp-265-274.pdf;   http://link.springer.com/article/10.1007/
BF02385257;   http://link.springer.com/article/10.1007/BF02385257/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=3&spage=265-274.

McLean:1995:SST

[222] Thomas J. McLean and David H. Xu. Study on sampling techniques with CMMs. Reliable Computing = Nadezhnye vychisleniia, 1(3):275– 284, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-275-284.
pdf;   http://link.springer.com/article/10.1007/BF02385258;
http://link.springer.com/article/10.1007/BF02385258/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=275-284.

Misane:1995:NCS

[223] Driss Misane and Vladik Kreinovich. A new characterization of the set of all intervals, based on the necessity to check consistency easily. Reliable Computing = Nadezhnye vychisleniia, 1(3):285–297, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-285-297.
pdf;   http://link.springer.com/article/10.1007/BF02385259;
http://link.springer.com/article/10.1007/BF02385259/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=285-297.

Nguyen:1995:ILN

[224] Hung T. Nguyen, Vladik Kreinovich, Bob Lea, and Dana Tolbert. Interpolation that leads to the narrowest intervals and its application to expert systems and intelligent control. Reliable Computing = Nadezhnye vychisleniia, 1(3):299–315, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
3-pp-299-315.pdf;   http://link.springer.com/article/10.1007/
BF02385260;   http://link.springer.com/article/10.1007/BF02385260/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=3&spage=299-315.

Schaefer:1995:PCZ

[225] Mark J. Schaefer and Tilmann Bubeck. A parallel complex zero finder. Reliable Computing = Nadezhnye vychisleniia, 1(3):317–323, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-317-323.
pdf;   http://link.springer.com/article/10.1007/BF02385261;

http://link.springer.com/article/10.1007/BF02385261/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=317-323.

Schulte:1995:SIH

[226] Michael J. Schulte and Earl E. Swartzlander, Jr. A software interface and hardware design for variable-precision interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 1(3):325–342, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-325-342.
pdf;   http://link.springer.com/article/10.1007/BF02385262;
http://link.springer.com/article/10.1007/BF02385262/;   http:
//www.springerlink.com/content/hj23t481181084t6/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
1&issue=3&spage=325-342.

Traylor:1995:BSN

[227] Bonnie Traylor and Vladik Kreinovich. A bright side of NPhardness of interval computations: interval heuristics applied to NPproblems. Reliable Computing = Nadezhnye vychisleniia, 1(3):343– 359, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-3-pp-343-359.
pdf;   http://link.springer.com/article/10.1007/BF02385263;
http://link.springer.com/article/10.1007/BF02385263/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=343-359.

Anonymous:1995:CP

[228] Anonymous. Call for papers. Reliable Computing = Nadezhnye vychisleniia, 1(3):361, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.edu/

reliable-computing-journal/1995/reliable-computing-1995-3-p-
361.pdf;   http://link.springer.com/accesspage/article/10.1007/
BF02385264;   http://link.springer.com/article/10.1007/BF02385264;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=3&spage=361.

Wang:1995:CPI

[229] Paul P. Wang. Canadian professor’s interval paper chosen best in fuzzy theory and technology. Reliable Computing = Nadezhnye vychisleniia, 1(3):362, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.edu/

reliable-computing-journal/1995/reliable-computing-1995-3-p-
362.pdf;   http://link.springer.com/accesspage/article/10.1007/
BF02385265;   http://link.springer.com/article/10.1007/BF02385265;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=3&spage=362.

Anonymous:1995:AEBb

[230] Anonymous. Addresses of the Editorial Board Members. Reliable Computing = Nadezhnye vychisleniia, 1(3):363–364, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http:

//link.springer.com/accesspage/article/10.1007/BF02385266;
http://link.springer.com/article/10.1007/BF02385266;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=3&spage=363-364.

Nesterov:1995:DCb

[231] V. Nesterov. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 1(4):371, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.edu/

reliable-computing-journal/1995/reliable-computing-1995-4-p-
371.pdf;   http://link.springer.com/accesspage/article/10.1007/
BF02391680;   http://link.springer.com/article/10.1007/BF02391680;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=4&spage=371.

Nesterov:1995:UKRb

[232] V. Nesterov. Uvazhaemye kollegi. (Russian) [Dear colleagues]. Reliable Computing = Nadezhnye vychisleniia, 1(4):372–373, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1995/reliable-
computing-1995-4-pp-372-373.pdf;   http://link.springer.com/
accesspage/article/10.1007/BF02391681;   http://link.springer.
com/article/10.1007/BF02391681;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=1&issue=4&spage=
372-373.

Akritas:1995:MCS

[233] A. G. Akritas, E. K. Akritas, and G. I. Malaschonok. Matrix computation of subresultant polynomial remainder sequences in integral domains. Reliable Computing = Nadezhnye vychisleniia, 1(4):375– 381, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-

computing-journal/1995/reliable-computing-1995-4-pp-375-381.
pdf;   http://link.springer.com/article/10.1007/BF02391682;
http://link.springer.com/article/10.1007/BF02391682/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=4&spage=375-381.

Dobronets:1995:NMU

[234] Boris S. Dobronets. Numerical methods using defects. Reliable Computing = Nadezhnye vychisleniia, 1(4):383–391, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1995/reliable-computing-
1995-4-pp-383-391.pdf;   http://link.springer.com/article/10.
1007/BF02391683;   http://link.springer.com/article/10.1007/
BF02391683/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=1&issue=4&spage=383-391.

Ostylovsky:1995:EAV

[235] A. N. Ostylovsky. An estimate of the absolute value and width of the solution of a linear system of equations with tridiagonal interval matrix by the interval sweep method. Reliable Computing = Nadezhnye vychisleniia, 1 (4):393–401, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-4-pp-393-401.
pdf;   http://link.springer.com/article/10.1007/BF02391684;
http://link.springer.com/article/10.1007/BF02391684/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=4&spage=393-401.

Petkovic:1995:VMI

[236] Ljiljana D. Petković and Miroslav Trajković. Verification methods for inclusion disks. Reliable Computing = Nadezhnye vychisleniia, 1(4):403– 410, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-4-pp-403-410.
pdf;   http://link.springer.com/article/10.1007/BF02391685;
http://link.springer.com/article/10.1007/BF02391685/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=4&spage=403-410.

vonGudenberg:1995:DPL

[237] Jürgen Wolff von Gudenberg. Design of a parallel linear algebra library for verified computation. Reliable Computing = Nadezhnye vychisleniia, 1(4):

411–419, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1995/reliable-computing-1995-4-pp-411-419.
pdf;   http://link.springer.com/article/10.1007/BF02391686;
http://link.springer.com/article/10.1007/BF02391686/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=1&issue=4&spage=411-419.

Wu:1995:CSR

[238] Q.-H. Wu and M. Mansour. Computation of the stability radius of a Schur polynomial: an orthogonal projection approach. Reliable Computing = Nadezhnye vychisleniia, 1(4):421–430, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1995/reliable-computing-1995-4-pp-421-430.pdf;   http:
//link.springer.com/article/10.1007/BF02391687;   http://link.
springer.com/article/10.1007/BF02391687/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=1&issue=
4&spage=421-430.

Anonymous:1995:FAC

[239] Anonymous. First announcement and call for papers. Reliable Computing = Nadezhnye vychisleniia, 1(4):431–432, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1995/reliable-computing-
1995-4-pp-431-432.pdf;   http://link.springer.com/accesspage/
article/10.1007/BF02391688;   http://link.springer.com/article/
10.1007/BF02391688;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=1&issue=4&spage=431-432.

Anonymous:1995:CC

[240] Anonymous. Call for contributions. Reliable Computing = Nadezhnye vychisleniia, 1(4):433, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.edu/

reliable-computing-journal/1995/reliable-computing-1995-4-p-
433.pdf;   http://link.springer.com/accesspage/article/10.1007/
BF02391689;   http://link.springer.com/article/10.1007/BF02391689;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=1&issue=4&spage=433.

Shary:1995:IGI

[241] Sergey Shary. IMACS–GAMM International Symposium on Numerical Methods and Error Bounds. Reliable Computing = Nadezhnye vychisleniia, 1(4):434–435, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1995/reliable-computing-1995-
4-pp-434-435.pdf;   http://link.springer.com/accesspage/article/
10.1007/BF02391690;   http://link.springer.com/article/10.1007/
BF02391690;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=1&issue=4&spage=434-435.

Anonymous:1995:AEBc

[242] Anonymous. Addresses of the Editorial Board Members. Reliable Computing = Nadezhnye vychisleniia, 1(4):436–437, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1995/reliable-
computing-1995-4-pp-436-437.pdf;   http://link.springer.com/
accesspage/article/10.1007/BF02391691;   http://link.springer.
com/article/10.1007/BF02391691;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=1&issue=4&spage=
436-437.

Anonymous:1995:Aa

[243] Anonymous. Amendments. Reliable Computing = Nadezhnye vychisleniia, 1(4):439, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02391692;   http://link.springer.com/article/
10.1007/BF02391692;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=1&issue=4&spage=439.

Anonymous:1995:Ab

[244] Anonymous. Amendments. Reliable Computing = Nadezhnye vychisleniia, 1(4):439, ???? 1995. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02391693;   http://link.springer.com/article/
10.1007/BF02391693;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=1&issue=4&spage=439.

Shary:1996:AAI

[245] Sergey P. Shary. Algebraic approach to the interval linear static identification, tolerance, and control problems, or one more application of Kaucher arithmetic. Reliable Computing = Nadezhnye vychisleniia, 2(1):3–33, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1996/reliable-computing-1996-1-pp-3-33.pdf;   http:/
/link.springer.com/article/10.1007/BF02388185;   http://link.

springer.com/article/10.1007/BF02388185/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=2&issue=
1&spage=3-33.

Dobner:1996:BHQ

[246] Hans-Jürgen Dobner. Bounds of high quality for first kind Volterra integral equations. Reliable Computing = Nadezhnye vychisleniia, 2(1):35–45, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1996/reliable-computing-1996-1-pp-35-45.pdf;   http:
//link.springer.com/article/10.1007/BF02388186;   http://link.
springer.com/article/10.1007/BF02388186/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=2&issue=
1&spage=35-45.

Schulte:1996:VPI

[247] Michael J. Schulte and Earl E. Swartzlander, Jr. Variable-precision, interval arithmetic coprocessors. Reliable Computing = Nadezhnye vychisleniia, 2(1):47–62, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
1-pp-47-62.pdf;   http://link.springer.com/article/10.1007/BF02388187;
 http://link.springer.com/article/10.1007/BF02388187/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=1&spage=47-62.

Kreinovich:1996:MEI

[248] Vladik Kreinovich. Maximum entropy and interval computations (September notes on summer impressions). Reliable Computing = Nadezhnye vychisleniia, 2(1):63–79, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
1-pp-63-79.pdf;   http://link.springer.com/article/10.1007/BF02388188;
 http://link.springer.com/article/10.1007/BF02388188/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=1&spage=63-79.

Anonymous:1996:CPRa

[249] Anonymous. Call for papers: Reliable Computing: An International Journal. Reliable Computing = Nadezhnye vychisleniia, 2(1): 81, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-1-p-81.pdf;

http://link.springer.com/accesspage/article/10.1007/BF02388190;
 http://link.springer.com/article/10.1007/BF02388190;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=1&spage=81.

Kreinovich:1996:BSP

[250] Vladik Kreinovich and Günter Mayer. Best student paper award. Reliable Computing = Nadezhnye vychisleniia, 2(1):81, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1996/reliable-
computing-1996-1-p-81.pdf;   http://link.springer.com/accesspage/
article/10.1007/BF02388189;   http://link.springer.com/article/
10.1007/BF02388189;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=2&issue=1&spage=81.

Anonymous:1996:CPRb

[251] Anonymous. Call for papers: Reliable Computing: An International Journal: Special Issue Applications to Geosciences. Reliable Computing = Nadezhnye vychisleniia, 2(1):82–83, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-1-pp-82-83.pdf;   http://link.springer.com/accesspage/article/
10.1007/BF02388191;   http://link.springer.com/article/10.1007/
BF02388191;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=1&spage=82-83.

Frommer:1996:SIS

[252] Andreas Frommer. SCAN’95: International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics, September 26–29, 1995, Wuppertal, Germany. Reliable Computing = Nadezhnye vychisleniia, 2(1):84–85, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
1-pp-84-85.pdf;   http://link.springer.com/accesspage/article/
10.1007/BF02388192;   http://link.springer.com/article/10.1007/
BF02388192;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=1&spage=84-85.

Kreinovich:1996:SSM

[253] Vladik Kreinovich. SONIC’95: Student Mini-Symposium on Interval Computations. Reliable Computing = Nadezhnye vychisleniia, 2(1):86–87, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-

journal/1996/reliable-computing-1996-1-pp-86-87.pdf;   http:
//link.springer.com/accesspage/article/10.1007/BF02388193;
http://link.springer.com/article/10.1007/BF02388193;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=1&spage=86-87.

Kreinovich:1996:ITF

[254] Vladik Kreinovich. Interval talks at the First El Paso Shell Oil Symposium. Reliable Computing = Nadezhnye vychisleniia, 2(1):88, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-1-p-88.pdf;
http://link.springer.com/accesspage/article/10.1007/BF02388194;
 http://link.springer.com/article/10.1007/BF02388194;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=1&spage=88.

Anonymous:1996:AEB

[255] Anonymous. Addresses of the Editorial Board Members. Reliable Computing = Nadezhnye vychisleniia, 2(1):89–90, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-1-pp-89-90.pdf;   http://link.springer.com/accesspage/article/
10.1007/BF02388195;   http://link.springer.com/article/10.1007/
BF02388195;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=1&spage=89-90.

Campos:1996:MTI

[256] Marcilia A. Campos, Augusto C. A. Sampaio, and Alexandre H. F. Brainer. Mechanising the theory of intervals using OBJ3. Reliable Computing = Nadezhnye vychisleniia, 2(2):97–102, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1996/reliable-
computing-1996-2-pp-97-102.pdf;   http://link.springer.com/article/
10.1007/BF02425910;   http://link.springer.com/article/10.1007/
BF02425910/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=2&spage=97-102.

Diverio:1996:EVP

[257] Tiarajú A. Diverio, Ursula A. Fernandes, and Dalcidio M. Claudio. Errors in vector processing and the library libavi.a. Reliable Computing = Nadezhnye vychisleniia, 2(2):103–109, ???? 1996. CODEN RCOMF8.

ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-2-pp-103-109.pdf;   http://link.springer.com/article/10.
1007/BF02425911;   http://link.springer.com/article/10.1007/
BF02425911/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=2&spage=103-109.

Heuveline:1996:CAT

[258] Vincent Heuveline and Miloud Sadkane. Chebyshev acceleration techniques for large complex non Hermitian eigenvalue problems. Reliable Computing = Nadezhnye vychisleniia, 2(2):111–117, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1996/reliable-computing-1996-2-pp-111-117.pdf;   http:
//link.springer.com/article/10.1007/BF02425912;   http://link.
springer.com/article/10.1007/BF02425912/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=2&issue=
2&spage=111-117.

Kreinovich:1996:IMG

[259] Vladik Kreinovich, Vyacheslav M. Nesterov, and Nina A. Zheludeva. Interval methods that are guaranteed to underestimate (and the resulting new justification of Kaucher arithmetic). Reliable Computing = Nadezhnye vychisleniia, 2(2):119–124, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
2-pp-119-124.pdf;   http://link.springer.com/article/10.1007/
BF02425913;   http://link.springer.com/article/10.1007/BF02425913/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=2&issue=2&spage=119-124.

Lakeyev:1996:CCS

[260] Anatoly V. Lakeyev. On the computational complexity of the solution of linear systems with moduli. Reliable Computing = Nadezhnye vychisleniia, 2(2):125–131, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
2-pp-125-131.pdf;   http://link.springer.com/article/10.1007/
BF02425914;   http://link.springer.com/article/10.1007/BF02425914/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=2&issue=2&spage=125-131.

Lynch:1996:SHR

[261] Thomas Lynch and Michael J. Schulte. Software for high radix on-line arithmetic. Reliable Computing = Nadezhnye vychisleniia, 2(2):133– 138, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-2-pp-133-138.
pdf;   http://link.springer.com/article/10.1007/BF02425915;
http://link.springer.com/article/10.1007/BF02425915/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=133-138.

Noubir:1996:SCP

[262] Guevara Noubir and Henri J. Nussbaumer. Self-correcting polynomial programs. Reliable Computing = Nadezhnye vychisleniia, 2(2):139– 145, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-2-pp-139-145.
pdf;   http://link.springer.com/article/10.1007/BF02425916;
http://link.springer.com/article/10.1007/BF02425916/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=139-145.

Oberman:1996:RDL

[263] Stuart F. Oberman and Michael J. Flynn. Reducing division latency with reciprocal caches. Reliable Computing = Nadezhnye vychisleniia, 2(2): 147–153, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-2-pp-147-153.
pdf;   http://link.springer.com/article/10.1007/BF02425917;
http://link.springer.com/article/10.1007/BF02425917/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=147-153.

Pesonen:1996:IAC

[264] Janne Pesonen and Eero Hyvönen. Interval approach challenges Monte Carlo simulation. Reliable Computing = Nadezhnye vychisleniia, 2(2): 155–160, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-2-pp-155-160.
pdf;   http://link.springer.com/article/10.1007/BF02425918;
http://link.springer.com/article/10.1007/BF02425918/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=155-160.

Popova:1996:IOI

[265] Evgenija D. Popova. Interval operations involving NaNs. Reliable Computing = Nadezhnye vychisleniia, 2(2):161–165, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1996/reliable-
computing-1996-2-pp-161-165.pdf;   http://link.springer.com/
article/10.1007/BF02425919;   http://link.springer.com/article/
10.1007/BF02425919/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=2&issue=2&spage=161-165.

Rohn:1996:ESO

[266] Jiří Rohn. Enclosing solutions of overdetermined systems of linear interval equations. Reliable Computing = Nadezhnye vychisleniia, 2(2): 167–171, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-2-pp-167-171.
pdf;   http://link.springer.com/article/10.1007/BF02425920;
http://link.springer.com/article/10.1007/BF02425920/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=167-171.

Sugihara:1996:NSB

[267] Masaaki Sugihara and Seiji Fujino. Numerical solutions of Burgers’ equation with a large Reynolds number. Reliable Computing = Nadezhnye vychisleniia, 2(2):173–179, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
2-pp-173-179.pdf;   http://link.springer.com/article/10.1007/
BF02425921;   http://link.springer.com/article/10.1007/BF02425921/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=2&issue=2&spage=173-179.

Szulc:1996:RCC

[268] Tomasz Szulc. Rank of convex combinations of matrices. Reliable Computing = Nadezhnye vychisleniia, 2(2):181–185, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

interval.louisiana.edu/reliable-computing-journal/1996/reliable-
computing-1996-2-pp-181-185.pdf;   http://link.springer.com/
article/10.1007/BF02425922;   http://link.springer.com/article/
10.1007/BF02425922/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=2&issue=2&spage=181-185.

Vrahatis:1996:LCC

[269] Michael N. Vrahatis and Evangelia C. Triantafyllou. Locating, characterizing and computing the stationary points of a function. Reliable Computing = Nadezhnye vychisleniia, 2(2):187–193, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1996/reliable-computing-1996-2-pp-187-193.pdf;   http:
//link.springer.com/article/10.1007/BF02425923;   http://link.
springer.com/article/10.1007/BF02425923/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=2&issue=
2&spage=187-193.

Anonymous:1996:ARS

[270] Anonymous. Reviews: Applications of reliable scientific computing. Reliable Computing = Nadezhnye vychisleniia, 2(2):195–203, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-2-pp-195-203.
pdf;   http://link.springer.com/article/10.1007/BF02425924;
http://link.springer.com/article/10.1007/BF02425924/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=195-203.

Alefeld:1996:Pa

[271] Götz E. Alefeld, Andreas Frommer, and Bruno Lang. Preface. Reliable Computing = Nadezhnye vychisleniia, 2(2):i, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http:

//link.springer.com/accesspage/article/10.1007/BF02425909;
http://link.springer.com/article/10.1007/BF02425909;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=2&spage=i.

Alefeld:1996:Pb

[272] Götz E. Alefeld, Andreas Frommer, and Bruno Lang. Preface. Reliable Computing = Nadezhnye vychisleniia, 2(2):iv, 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-2-p-iv.pdf.

Eick:1996:RAL

[273] Christoph Eick and Karen Villaverde. Robust algorithms that locate local extrema of a function of one variable from interval measurement results: a remark. Reliable Computing = Nadezhnye vychisleniia, 2(3): 213–218, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-3-pp-213-218.
pdf;   http://link.springer.com/article/10.1007/BF02391695;
http://link.springer.com/article/10.1007/BF02391695/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=3&spage=213-218.

Ferregut:1996:FEE

[274] Carlos Ferregut, Soheil Nazarian, Krishnamohan Vennalaganti, Ching-Chuan Chang, and Vladik Kreinovich. Fast error estimates for indirect measurements: applications to pavement engineering. Reliable Computing = Nadezhnye vychisleniia, 2(3):219–228, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-3-pp-219-228.pdf;   http://link.springer.com/article/10.
1007/BF02391696;   http://link.springer.com/article/10.1007/
BF02391696/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=3&spage=219-228.

Holzmann:1996:NCG

[275] Oliver Holzmann, Bruno Lang, and Holger Schütt. Newton’s constant of gravitation and verified numerical quadrature. Reliable Computing = Nadezhnye vychisleniia, 2(3):229–239, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-3-pp-229-239.pdf;   http://link.springer.com/article/10.
1007/BF02391697;   http://link.springer.com/article/10.1007/
BF02391697/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=3&spage=229-239.

Kramer:1996:TAG

[276] Walter Krämer and Stefan Wedner. Two adaptive Gauss–Legendre type algorithms for the verified computation of definite integrals. Reliable Computing = Nadezhnye vychisleniia, 2(3):241–253, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://interval.louisiana.edu/reliable-computing-
journal/1996/reliable-computing-1996-3-pp-241-253.pdf;   http:
//link.springer.com/article/10.1007/BF02391698;   http://link.
springer.com/article/10.1007/BF02391698/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=2&issue=
3&spage=241-253.

Kreinovich:1996:QTA

[277] Vladik Kreinovich and Karen Villaverde. A quadratic-time algorithm for smoothing interval functions. Reliable Computing = Nadezhnye vychisleniia, 2(3):255–264, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
3-pp-255-264.pdf;   http://link.springer.com/article/10.1007/
BF02391699;   http://link.springer.com/article/10.1007/BF02391699/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=2&issue=3&spage=255-264.

Lea:1996:OIE

[278] Robert N. Lea, Vladik Kreinovich, and Raul Trejo. Optimal interval enclosures for fractionally-linear functions, and their application to intelligent control. Reliable Computing = Nadezhnye vychisleniia, 2(3): 265–285, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-3-pp-265-285.
pdf;   http://link.springer.com/article/10.1007/BF02391700;
http://link.springer.com/article/10.1007/BF02391700/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=3&spage=265-285.

Lorkowski:1996:IWM

[279] Joe Lorkowski and Vladik Kreinovich. If we measure a number, we get an interval. What if we measure a function or an operator? Reliable Computing = Nadezhnye vychisleniia, 2(3):287–297, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-3-pp-287-297.
pdf;   http://link.springer.com/article/10.1007/BF02391701;
http://link.springer.com/article/10.1007/BF02391701/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=3&spage=287-297.

Oliveira:1996:NSM

[280] João B. Oliveira. New slope methods for sharper interval functions and a note on Fischer’s acceleration method. Reliable Computing = Nadezhnye vychisleniia, 2(3):299–320, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
3-pp-299-320.pdf;   http://link.springer.com/article/10.1007/

BF02391702;   http://link.springer.com/article/10.1007/BF02391702/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=2&issue=3&spage=299-320.

Provetti:1996:OEI

[281] Alessandro Provetti. Ordering events: Intervals are sufficient, more general sets are usually not necessary. Reliable Computing = Nadezhnye vychisleniia, 2(3):321–327, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
3-pp-321-327.pdf;   http://link.springer.com/article/10.1007/
BF02391703;   http://link.springer.com/article/10.1007/BF02391703/
;   http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=2&issue=3&spage=321-327.

Beltran:1996:RAR

[282] M. Beltran and D. E. Cooke. Reviews: Applications of reliable scientific computing. Reliable Computing = Nadezhnye vychisleniia, 2(3): 329–331, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-3-pp-329-331.
pdf;   http://link.springer.com/article/10.1007/BF02391704;
http://link.springer.com/article/10.1007/BF02391704/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=3&spage=329-331.

Mayer:1996:DC

[283] Günther Mayer and Vladik Kreinovich. Dear colleagues! Reliable Computing = Nadezhnye vychisleniia, 2(3):i, 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-3-p-i.pdf.

Mayer:1996:P

[284] Günther Mayer and Vladik Kreinovich. Preface. Reliable Computing = Nadezhnye vychisleniia, 2(3):i, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
3-p-i.pdf;   http://link.springer.com/accesspage/article/10.
1007/BF02391694;   http://link.springer.com/article/10.1007/
BF02391694;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=2&issue=3&spage=i.

Kreinovich:1996:LIE

[285] V. Kreinovich and A. V. Lakeyev. Linear interval equations: Computing enclosures with bounded relative or absolute overestimation is NP-hard. Reliable Computing = Nadezhnye vychisleniia, 2(4):341– 350, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-4-pp-341-350.
pdf;   http://link.springer.com/article/10.1007/BF02389894;
http://link.springer.com/article/10.1007/BF02389894/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=4&spage=341-350.

Lyashko:1996:SCT

[286] Marina A. Lyashko. On the speed of convergence of the total step iterative method for a class of interval linear algebraic systems. Reliable Computing = Nadezhnye vychisleniia, 2(4):351–356, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-4-pp-351-356.
pdf;   http://link.springer.com/article/10.1007/BF02389895;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=2&issue=4&spage=351-356.

Shi:1996:IEI

[287] Yixun Shi. Improving the efficiency index in enclosing a root of an equation. Reliable Computing = Nadezhnye vychisleniia, 2(4):357– 372, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-4-pp-357-372.
pdf;   http://link.springer.com/article/10.1007/BF02389896;
http://link.springer.com/article/10.1007/BF02389896/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=4&spage=357-372.

Starner:1996:SEU

[288] John W. Starner. Some examples using the interval data type in the relational database model. Reliable Computing = Nadezhnye vychisleniia, 2 (4):373–381, ???? 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-4-pp-373-381.
pdf;   http://link.springer.com/article/10.1007/BF02389897;
http://link.springer.com/article/10.1007/BF02389897/;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=2&issue=4&spage=373-381.

Kreinovich:1996:PE

[289] V. Kreinovich. Paul Erdős, 1913–1996. Reliable Computing = Nadezhnye vychisleniia, 2(4):383–386, December 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-4-pp-383-386.pdf;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=2&issue=4&spage=383.

Kreinovich:1996:DBM

[290] Vladik Kreinovich. Double bubble minimizes: Interval computations help in solving a long-standing geometric problems. Reliable Computing = Nadezhnye vychisleniia, 2(4):387–388, 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.

louisiana.edu/reliable-computing-journal/1996/reliable-computing-
1996-4-pp-387-388.pdf.

Anonymous:1996:IJD

[291] Anonymous. An international journal devoted to reliable mathematical computations based on finite representation and guaranteed accuracy: Reliable Computing. Reliable Computing = Nadezhnye vychisleniia, 2 (4):389–390, 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-4-pp-389-390.
pdf.

Campos:1996:WIW

[292] Marcilia A. Campos, Erich Peter Klement, and Vladik Kreinovich. WAI’96: II Workshop on Computer Arithmetic, Interval and Symbolic Computation. Reliable Computing = Nadezhnye vychisleniia, 2 (4):391–401, December 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://interval.louisiana.

edu/reliable-computing-journal/1996/reliable-computing-1996-
4-pp-391-401.pdf;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=2&issue=4&spage=391.

Nesterov:1996:DC

[293] V. Nesterov. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 2(4):i, 1996. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://interval.louisiana.edu/reliable-
computing-journal/1996/reliable-computing-1996-4-p-i.pdf.

Rump:1997:IZN

[294] Siegfried M. Rump. Inclusion of zeros of nowhere differentiable ndimensional functions. Reliable Computing = Nadezhnye vychisleniia, 3(1):5–16, February 1997. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009967901122/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=3&issue=1&spage=5-16.

Hansen:1997:SIC

[295] Eldon R. Hansen. Sharpness in interval computations. Reliable Computing = Nadezhnye vychisleniia, 3(1):17–29, February 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009917818868/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=1&spage=17-29.

Wolfe:1997:IAB

[296] Michael A. Wolfe. An interval algorithm for bounding the ranges of real-valued functions of one real variable. Reliable Computing = Nadezhnye vychisleniia, 3(1):31–50, February 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009973905665/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=1&spage=31-50.

Lakeyev:1997:NHC

[297] A. V. Lakeyev and V. Kreinovich. NP-hard classes of linear algebraic systems with uncertainties. Reliable Computing = Nadezhnye vychisleniia, 3(1):51–81, February 1997. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009938325229/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=3&issue=1&spage=51-81.

Kolev:1997:UIS

[298] Lubomir V. Kolev. Use of interval slopes for the irrational part of factorable functions. Reliable Computing = Nadezhnye vychisleniia, 3(1):83– 93, February 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009902813842/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=1&spage=83-93.

Nguyen:1997:NIS

[299] Hung T. Nguyen and Vladik Kreinovich. From numerical intervals to set intervals (interval-related results presented at the First International Workshop on Applications and Theory of Random Sets). Reliable Computing = Nadezhnye vychisleniia, 3(1): 95–102, February 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000013418.91863.98/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
3&issue=1&spage=95-102.

Shary:1997:AAO

[300] Sergey P. Shary. Algebraic approach in the “outer problem” for interval linear equations. Reliable Computing = Nadezhnye vychisleniia, 3(2):103– 135, May 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009975421252/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=2&spage=103-135.

Minamoto:1997:NVS

[301] Teruya Minamoto and Mitsuhiro T. Nakao. Numerical verifications of solutions for nonlinear parabolic equations in one-space dimensional case. Reliable Computing = Nadezhnye vychisleniia, 3(2):137–147, May 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009933805322/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=2&spage=137-147.

Kreinovich:1997:SSS

[302] Vladik Kreinovich and Andrew Bernat. Is solar system stable? A remark. Reliable Computing = Nadezhnye vychisleniia, 3(2):149–154, May 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009985822160/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=2&spage=149-154.

Chee:1997:CVB

[303] Leticia S. Chee. Computing the value of a Boolean expression with interval inputs is NP-Hard. Reliable Computing = Nadezhnye vychisleniia, 3 (2):155–172, May 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009937906231/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=2&spage=155-172.

Nedoma:1997:SSS

[304] Josef Nedoma. Sign-stable solutions of column-vague linear equation systems. Reliable Computing = Nadezhnye vychisleniia, 3(2):173–180, May 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009989923069/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=2&spage=173-180.

vonGudenberg:1997:DC

[305] Jürgen Wolff von Gudenberg. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 3(3):195–198, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009942015805/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=195-198.

Markov:1997:IEA

[306] Svetoslav M. Markov. Isomorphic embeddings of abstract interval systems. Reliable Computing = Nadezhnye vychisleniia, 3(3):199–207, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009967232643/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=199-207.

Kulpa:1997:DRI

[307] Zenon Kulpa. Diagrammatic representation of interval space in proving theorems about interval relations. Reliable Computing = Nadezhnye vychisleniia, 3(3):209–217, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009919304728/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=209-217.

Berthelot:1997:CSE

[308] David Berthelot and Marc Daumas. Computing on sequences of embedded intervals. Reliable Computing = Nadezhnye vychisleniia, 3(3):219–227, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009962521566/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=3&spage=219-227.

Weinhofer:1997:CUP

[309] Jürgen K. Weinhofer and Werner C. Haas. H_{∞}-control using polynomial matrices and interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 3(3):229–237, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009914605637/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=229-237.

Hofschuster:1997:COA

[310] Werner Hofschuster and Walter Krämer. A computer oriented approach to get sharp reliable error bounds. Reliable Computing = Nadezhnye vychisleniia, 3(3):239–248, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009966622475/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=239;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=239-248.

Luther:1997:AEE

[311] Wolfram Luther and Werner Otten. Approximation error and error accumulation for the Landen transform. Reliable Computing = Nadezhnye vychisleniia, 3(3):249–258, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009918706545/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=249;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=249-258.

Bliek:1997:FEP

[312] Christian Bliek. Fast evaluation of partial derivatives and interval slopes. Reliable Computing = Nadezhnye vychisleniia, 3(3):259–268, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009970723383/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=259;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=259-268.

Caprani:1997:ETA

[313] Ole Caprani, Kaj Madsen, and Ole Stauning. Existence test for asynchronous interval iteration. Reliable Computing = Nadezhnye vychisleniia, 3(3):269–275, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009922807454/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=3&issue=3&spage=269;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=269-275.

Mrozek:1997:HCK

[314] Marian Mrozek and Marcin Żelawski. Heteroclinic connections in the Kuramoto–Sivashinsky equation: a computer assisted proof. Reliable Computing = Nadezhnye vychisleniia, 3(3):277–285, August 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009974824292/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=277;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=277-285.

Dobner:1997:ACM

[315] Hans-Jürgen Dobner and Stefan Ritter. Attacking a conjecture in mathematical physics by combining methods of computational analysis and scientific computing. Reliable Computing = Nadezhnye vychisleniia, 3(3): 287–295, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009927009271/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=3&spage=287;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=3&issue=3&spage=287-295.

Dobronets:1997:TSM

[316] Boris S. Dobronets. Two-sided multigrid method for elliptic boundary value problems. Reliable Computing = Nadezhnye vychisleniia, 3(3):297– 303, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009983126109/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=3&spage=297;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=3&issue=3&spage=297-303.

Acioly:1997:QMT

[317] Benedito Melo Acióly and Benjamín R. Callejas Bedregal. A quasi-metric topology compatible with inclusion monotonicity on interval space. Reliable Computing = Nadezhnye vychisleniia, 3(3):305–313, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009935210180/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=305;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=305-313.

Rohn:1997:CSL

[318] Jiří Rohn. Complexity of some linear problems with interval data. Reliable Computing = Nadezhnye vychisleniia, 3(3):315–323, August 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009987227018/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=315;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=315-323.

Rueher:1997:CCS

[319] Michel Rueher and Christine Solnon. Concurrent cooperating solvers over reals. Reliable Computing = Nadezhnye vychisleniia, 3(3):325–333, August 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009939327927/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=3&spage=325;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=3&issue=3&spage=325-333.

Benhamou:1997:AGN

[320] Frédéric Benhamou and Laurent Granvilliers. Automatic generation of numerical redundancies for non-linear constraint solving. Reliable Computing = Nadezhnye vychisleniia, 3(3):335–344, August 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009943413814/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=335;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=335-344.

Anonymous:1997:FAC

[321] Anonymous. First announcement and call for papers. Reliable Computing = Nadezhnye vychisleniia, 3(3):345–347, August 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017130516540/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=3&spage=345;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
3&spage=345-347.

Petkovic:1997:SMH

[322] Ljiljana D. Petković, Slobodan Tricković, and Miodrag S. Petković. Slope methods of higher order for the inclusion of complex roots of polynomials. Reliable Computing = Nadezhnye vychisleniia, 3(4):349–362, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009903702722/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=4&spage=349;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&

volume=3&issue=4&spage=349-362. Special issue on Reliable Geometric Computations.

Rohn:1997:OPI

[323] Jiří Rohn. On overestimations produced by the interval Gaussian algorithm (dedicated to Prof. Dr. Gerhard Heindl on the occasion of his 60th birthday). Reliable Computing = Nadezhnye vychisleniia, 3(4):363–368, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009993319560/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=4&spage=363;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=3&issue=4&spage=363-368. Dedicated to Prof. Dr. Gerhard

Heindl on the occasion of his 60th birthday. Special issue on Reliable Geometric Computations.

Nesterov:1997:ITA

[324] Vyacheslav M. Nesterov. Interval and twin arithmetics. Reliable Computing = Nadezhnye vychisleniia, 3(4):369–380, November 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009945403631/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=369;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=369-380. Special issue on Reliable Geometric Computations.

Kalovics:1997:FGM

[325] Ferenc Kálovics and Gabriella Mészáros. Finding global minima of maximum functions by using exclusion functions without derivatives. Reliable Computing = Nadezhnye vychisleniia, 3(4):381–399, November 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009997420469/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=381;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=381-399. Special issue on Reliable Geometric Computations.

Smagina:1997:NAM

[326] Yelena M. Smagina. A new approach to the modal regulator synthesis for interval plant with scalar input. Reliable Computing = Nadezhnye vychisleniia, 3(4):401–410, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009949504539/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=401;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=401-410. Special issue on Reliable Geometric Computations.

Stahl:1997:ERT

[327] Volker Stahl. Error reduction of the Taylor centered form by half and an inner estimation of the range. Reliable Computing = Nadezhnye vychisleniia, 3(4):411–420, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009901621378/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=411;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=411-420. Special issue on Reliable Geometric Computations.

Heindl:1997:HCI

[328] Gerhard Heindl. How to compute interval inclusions of geodetic coordinates from interval inclusions of Cartesian coordinates. Reliable Computing = Nadezhnye vychisleniia, 3(4):421–435, November 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009953605448/;

 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=421;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=421-435. Special issue on Reliable Geometric Computations.

Kreinovich:1997:TJC

[329] Vladik Kreinovich, Scott Starks, and Günter Mayer. On a theoretical justification of the choice of epsilon-inflation in PASCAL–XSC. Reliable Computing = Nadezhnye vychisleniia, 3(4):437–445, November 1997. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009905822286/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=437;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=437-445. Special issue on Reliable Geometric Computations.

Gabaldon:1997:ARS

[330] A. Gabaldon. Applications of reliable scientific computing. Reliable Computing = Nadezhnye vychisleniia, 3(4):447–448, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1009957806357;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=3&issue=4&spage=447;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&

volume=3&issue=4&spage=447-448. Special issue on Reliable Geometric Computations.

Anonymous:1997:AC

[331] Anonymous. Applications to control. Reliable Computing = Nadezhnye vychisleniia, 3(4):448–452, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017116623195/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=
4&spage=448-452.

Longpre:1997:ICW

[332] Luc Longpré and Martin Berz. Interval and complexity workshops backto-back with 1997 ACM Symposium on Theory of Computing (STOC’97). Reliable Computing = Nadezhnye vychisleniia, 3(4):453–457, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009909907265/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=3&issue=4&spage=453;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&

volume=3&issue=4&spage=453-457. Special issue on Reliable Geometric Computations.

Nogueira:1997:IRT

[333] Monica Nogueira. Interval-related talks at NASA URC conference. Reliable Computing = Nadezhnye vychisleniia, 3(4):459–460, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1009914024104;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=3&issue=4&spage=459;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&

volume=3&issue=4&spage=459-460. Special issue on Reliable Geometric Computations.

Ratschek:1997:RCS

[334] H. Ratschek and J. Rokne. Reliable Computing special issue on reliable geometric computations. Reliable Computing = Nadezhnye vychisleniia, 3(4):461–462, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1009966008174;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=3&issue=4&spage=461;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=3&issue=4&spage=

461-462. Special Issue on Reliable Geometric Computations.

Anonymous:1997:BPA

[335] Anonymous. Best paper award to Zdzislaw Pawlak. Reliable Computing = Nadezhnye vychisleniia, 3(4):463, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017161925012;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=3&issue=4&spage=463. Special Issue on Reliable Geomet-

ric Computations.

Anonymous:1997:PS

[336] Anonymous. Patrick Suppes is 75. Reliable Computing = Nadezhnye vychisleniia, 3(4):465, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017170109082;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=3&issue=4&spage=465. Special Issue on Reliable Geomet-

ric Computations.

Anonymous:1997:AIV

[337] Anonymous. Author index/volume contents. Reliable Computing = Nadezhnye vychisleniia, 3(4):467–473, November 1997. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017173108660/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=3&issue=4&spage=467;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=3&issue=

4&spage=467-473. Special Issue on Reliable Geometric Computations.

Mayer:1998:F

[338] Günter Mayer and Vladik Kreinovich. Foreword. Reliable Computing = Nadezhnye vychisleniia, 4(1):1–2, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1009990214039;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=4&issue=1&spage=1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=4&issue=1&spage=
1-2.

Hungerbuhler:1998:BRB

[339] Ralf Hungerbühler and Jürgen Garloff. Bounds for the range of a bivariate polynomial over a triangle. Reliable Computing = Nadezhnye vychisleniia, 4(1):3–13, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009942430877/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=4&issue=1&spage=3;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=4&issue=1&spage=3-13.

Huang:1998:IEP

[340] Zhenyu Huang. An interval entropy penalty method for nonlinear global optimization. Reliable Computing = Nadezhnye vychisleniia, 4(1):15–25, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009994414947/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=1&spage=15;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=15-25. See errata and comments [610].

Hu:1998:IPI

[341] Chenyi Hu, Angelina Cardenas, Stephanie Hoogendoorn, and Pedro Sepulveda, Jr. An interval polynomial interpolation problem and its Lagrange solution. Reliable Computing = Nadezhnye vychisleniia, 4(1):27– 38, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009946531786/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=1&spage=27;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=27-38.

Beltran:1998:ASP

[342] Maria Beltran, Gilbert Castillo, and Vladik Kreinovich. Algorithms that still produce a solution (maybe not optimal) even when interrupted: Shary’s idea justified. Reliable Computing = Nadezhnye vychisleniia, 4(1): 39–53, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009998515856/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=1&spage=39;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=39-53.

Morales:1998:IMR

[343] David Morales and Tran Cao Son. Interval methods in robot navigation. Reliable Computing = Nadezhnye vychisleniia, 4(1):55–61, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009950632694/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=1&spage=55;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
1&spage=55-61.

Hu:1998:ROP

[344] Zhihui Huey Hu. Reliable optimal production control with Cobb–Douglas model. Reliable Computing = Nadezhnye vychisleniia, 4(1):63–69, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009902716765/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=1&spage=63;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=63-69.

Berleant:1998:STA

[345] Daniel Berleant and Hang Cheng. A software tool for automatically verified operations on intervals and probability distributions. Reliable Computing = Nadezhnye vychisleniia, 4(1):71–82, February 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009954817673/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=1&spage=71;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
1&spage=71-82.

Berz:1998:CAT

[346] Martin Berz and Georg Hoffstätter. Computation and application of Taylor polynomials with interval remainder bounds. Reliable Computing = Nadezhnye vychisleniia, 4(1):83–97, February 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009958918582/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=1&spage=83;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
1&spage=83-97.

Davis:1998:NAL

[347] Clifton Davis. A new application of local minima of interval functions: Interval-valued fuzzy control. Reliable Computing = Nadezhnye vychisleniia, 4(1):99–101, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009963019490/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
1&spage=99;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=4&issue=1&spage=99-101.

Gonzalez:1998:IHD

[348] Graciela Gonzalez. Intervals help to design an imaging system. Reliable Computing = Nadezhnye vychisleniia, 4(1):103–104, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1009967121308;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=4&issue=1&spage=103;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=103-104.

Koshelev:1998:BDG

[349] Misha Koshelev and Luc Longpré. A brief description of Gell-Mann’s lecture and how intervals may help to describe complexity in the real world. Reliable Computing = Nadezhnye vychisleniia, 4(1):105–107, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009975305378/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=1&spage=105;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=105-107.

Levichev:1998:IST

[350] Alexander Levichev and Olga Kosheleva. Intervals in space–time: A. D. Alexandrov is 85. Reliable Computing = Nadezhnye vychisleniia, 4(1):109– 112, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009927422216/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=1&spage=109;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=1&spage=109-112.

Garloff:1998:CPR

[351] J. Garloff and É. Walter. Call for papers: Reliable computing special issue on applications to control, signals, and systems. Reliable Computing = Nadezhnye vychisleniia, 4(1):113–114, February 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017154506287;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=4&issue=1&spage=113;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=4&issue=1&spage=
113-114.

Nakao:1998:CEE

[352] Mitsuhiro T. Nakao, Nobito Yamamoto, and Yoshitaka Watanabe. Constructive L^{2} error estimates for finite element solutions of the Stokes equations. Reliable Computing = Nadezhnye vychisleniia, 4(2):115–124, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009979408417/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=2&spage=115;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=2&spage=115-124.

Kolev:1998:NMG

[353] Lubomir V. Kolev. A new method for global solution of systems of nonlinear equations. Reliable Computing = Nadezhnye vychisleniia, 4(2):125– 146, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009981025255/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=2&spage=125;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=2&spage=125-146.

Berleant:1998:BRA

[354] Daniel Berleant and Chaim Goodman-Strauss. Bounding the results of arithmetic operations on random variables of unknown dependency using intervals. Reliable Computing = Nadezhnye vychisleniia, 4(2):147– 165, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://ee.iastate.edu/~djb/Research/Pdfs/
unknownDependency.ps;   http://link.springer.com/article/10.
1023/A%3A1009933109326/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=4&issue=2&spage=147;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=2&spage=147-165.

Popov:1998:RBM

[355] Antony T. Popov. A relation between morphological and interval operations. Reliable Computing = Nadezhnye vychisleniia, 4(2):167–178, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009985126164/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=2&spage=167;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=2&spage=167-178.

Kosheleva:1998:WPI

[356] Olga Kosheleva and Piet G. Vroegindeweij. When is the product of intervals also an interval? Reliable Computing = Nadezhnye vychisleniia, 4(2): 179–190, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009937210234/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=2&spage=179;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=2&spage=179-190.

Dennis:1998:IOC

[357] David Dennis, Vladik Kreinovich, and Siegfried M. Rump. Intervals and the origins of calculus. Reliable Computing = Nadezhnye vychisleniia, 4 (2):191–197, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/

A%3A1009989211143/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=2&spage=191;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=2&spage=191-197.

Plotnikov:1998:OCE

[358] Anatoly D. Plotnikov. One criterion of existence of a Hamiltonian cycle. Reliable Computing = Nadezhnye vychisleniia, 4(2):199–202, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009993312051/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=2&spage=199;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
2&spage=199-202.

Anonymous:1998:FA

[359] Anonymous. First announcement. Reliable Computing = Nadezhnye vychisleniia, 4(2):203–204, May 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1023/A%3A1017143730707;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
2&spage=203;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=4&issue=2&spage=203-204.

Mayer:1998:AIG

[360] Günter Mayer and Jiří Rohn. On the applicability of the interval Gaussian algorithm. Reliable Computing = Nadezhnye vychisleniia, 4(3):205–222, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009997411503/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=3&spage=205;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=3&spage=205-222.

Wolfe:1998:SDT

[361] Michael A. Wolfe. On a second derivative test due to Qi. Reliable Computing = Nadezhnye vychisleniia, 4(3):223–234, August 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009999328341/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=3&spage=223;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
3&spage=223-234.

Baker:1998:UAT

[362] Mark R. Baker and Rajendra B. Patil. Universal approximation theorem for interval neural networks. Reliable Computing = Nadezhnye vychisleniia, 4(3):235–239, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009951412412/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=3&spage=235;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
3&spage=235-239.

Doser:1998:EUG

[363] Diane I. Doser, Kevin D. Crain, Mark R. Baker, Vladik Kreinovich, and Matthew C. Gerstenberger. Estimating uncertainties for geophysical tomography. Reliable Computing = Nadezhnye vychisleniia, 4(3):241–268, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009903529250/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=3&spage=241;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=3&spage=241-268.

Birdie:1998:UIA

[364] Tiraz R. Birdie and Karan S. Surana. The use of interval analysis in hydrologic systems. Reliable Computing = Nadezhnye vychisleniia, 4(3):269– 281, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009955613320/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=3&spage=269;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=3&spage=269-281.

Popov:1998:AFM

[365] Antony T. Popov, Hung T. Nguyen, and Leonid K. Reznik. An application of fuzzy mathematical morphology to interval-valued knowledge representation: a remark. Reliable Computing = Nadezhnye vychisleniia, 4(3):283– 290, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009907730159/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=3&spage=283;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=3&spage=283-290.

Brito:1998:IIW

[366] Alejandro E. Brito and Olga Kosheleva. Interval + image = wavelet: For image processing under interval uncertainty, wavelets are optimal. Reliable Computing = Nadezhnye vychisleniia, 4(3):291–301, August 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009959714229/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=3&spage=291;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
3&spage=291-301.

Gabaldon:1998:EOL

[367] Alfredo Gabaldon and Hung T. Nguyen. Exclusive OR operation that leads to the narrowest intervals. Reliable Computing = Nadezhnye vychisleniia, 4(3):303–306, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009911831067/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=3&spage=303;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
3&spage=303-306.

Moore:1998:BRA

[368] Ramon Moore. Book review: Aberth, O.: Precise Numerical Methods Using C++. Reliable Computing = Nadezhnye vychisleniia, 4(3):307–308, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/accesspage/article/
10.1023/A%3A1009963815138;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=4&issue=3&spage=307;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=3&spage=307-308. See [776].

vonGudenberg:1998:ICI

[369] Jürgen Wolff von Gudenberg. International Conference Interval’98. Reliable Computing = Nadezhnye vychisleniia, 4(3):309, August 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1009915931976;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=4&issue=3&spage=309.

Anguelov:1998:WEW

[370] Roumen Anguelov and Svetoslav Markov. Wrapping effect and wrapping function. Reliable Computing = Nadezhnye vychisleniia, 4(4):311–330, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1024414730820/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=4&spage=311;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=4&spage=311-330.

Leclerc:1998:RCD

[371] Anthony Leclerc and Jeff Ely. Reliable collision detection for timedependent parametric surfaces. Reliable Computing = Nadezhnye vychisleniia, 4(4):331–344, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1024411614890/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=331;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=4&issue=4&spage=331-344.

Jansson:1998:NHR

[372] Christian Jansson. An NP-hardness result for nonlinear systems. Reliable Computing = Nadezhnye vychisleniia, 4(4):345–350, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024463631728/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=345;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=345-350.

Koshelev:1998:OEQ

[373] Misha Koshelev, Luc Longpré, and Patrick Taillibert. Optimal enclosure of quadratic interval functions. Reliable Computing = Nadezhnye vychisleniia, 4(4):351–360, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024415715798/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=351;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=351-360.

Berz:1998:VIO

[374] Martin Berz and Kyoko Makino. Verified integration of ODEs and flows using differential algebraic methods on high-order Taylor models. Reliable Computing = Nadezhnye vychisleniia, 4(4):361–369, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024467732637/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=361;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=361-369.

Rump:1998:NEI

[375] Siegfried M. Rump. A note on epsilon-inflation. Reliable Computing = Nadezhnye vychisleniia, 4(4):371–375, November 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024419816707/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=371;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=371-375.

Menshikov:1998:ISI

[376] Gregory G. Menshikov. Intersection subdistributivity and interval hull superdistributivity with respect to the interval maps. Reliable Computing = Nadezhnye vychisleniia, 4(4):377–381, November 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024471800777/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=377;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=377-381.

Heindl:1998:SLI

[377] Gerhard Heindl, Vladik Kreinovich, and Anatoly V. Lakeyev. Solving linear interval systems is NP-hard even if we exclude overflow and underflow. Reliable Computing = Nadezhnye vychisleniia, 4(4):383–388, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1024423917616/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=4&issue=4&spage=383;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=4&issue=4&spage=383-388.

Nogueira:1998:WIB

[378] Monica Nogueira and Amarendra Nandigam. Why intervals? because if we allow other sets, tractable problems become intractable. Reliable Computing = Nadezhnye vychisleniia, 4(4):389–394, November 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024475901686/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=389;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=389-394.

Merenyi:1998:HSS

[379] Erzsébet Merényi, Scott Starks, and Karen Villaverde. Hyper-spectral satellite images: Interval methods may be helpful. Reliable Computing = Nadezhnye vychisleniia, 4(4):395–397, November 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024428118524/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=395;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=395-397.

Kosheleva:1998:HPA

[380] Olga M. Kosheleva. Hilbert problems (almost) 100 years later (from the viewpoint of interval computations). Reliable Computing = Nadezhnye vychisleniia, 4(4):399–403, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024480119433/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=399;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=399-403.

Kearfott:1998:BRK

[381] R. Baker Kearfott. Book review: Kreinovich, V., Lakeyev, A., Rohn, J., and Kahl, P.: Computational Complexity and Feasibility of Data Processing and Interval Computations. Reliable Computing = Nadezhnye vychisleniia, 4(4):405–409, November 1998. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024484203503/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=4&issue=4&spage=405;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=4&issue=
4&spage=405-409. See [777].

Corliss:1998:SAM

[382] George Corliss and Ramon Moore. 1998 SIAM Annual Meeting: Interval Highlights — Toronto, July 13–17, 1998. Reliable Computing = Nadezhnye vychisleniia, 4(4):411–412, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1024436320342;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=4&issue=4&spage=411;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=4&issue=4&spage=
411-412.

Anonymous:1998:FAC

[383] Anonymous. First Announcement and Call for Papers ICRA99 — International Conference on Rational Approximation June 6–11, 1999 University of Antwerp (UIA). Reliable Computing = Nadezhnye vychisleniia, 4(4):413–414, November 1998. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1024482404412;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=4&issue=4&spage=413;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=4&issue=4&spage=
413-414.

Zuhe:1999:P

[384] Shen Zuhe and Vladik Kreinovich. Preface. Reliable Computing = Nadezhnye vychisleniia, 5(1):1, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1026471414671;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=5&issue=1&spage=1.

Makino:1999:ECD

[385] Kyoko Makino and Martin Berz. Efficient control of the dependency problem based on Taylor model methods. Reliable Computing = Nadezhnye vychisleniia, 5(1):3–12, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026485406803/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=1&spage=3;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
1&spage=3-12.

Berz:1999:NMH

[386] Martin Berz and Kyoko Makino. New methods for high-dimensional verified quadrature. Reliable Computing = Nadezhnye vychisleniia, 5(1):13– 22, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026437523641/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=1&spage=13;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=1&spage=13-22.

Zhang:1999:SUS

[387] Danqing Zhang, Weiguo Li, and Zuhe Shen. Solving underdetermined systems with interval methods. Reliable Computing = Nadezhnye vychisleniia, 5(1):23–33, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026489507711/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
1&spage=23;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=5&issue=1&spage=23-33.

Funge:1999:RKW

[388] John Funge. Representing knowledge within the situation calculus using interval-valued epistemic fluents. Reliable Computing = Nadezhnye vychisleniia, 5(1):35–61, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026441624550/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=1&spage=35;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
1&spage=35-61.

Heindl:1999:HGF

[389] Gerhard Heindl. How to guarantee finite termination of verifying global optimization codes. Reliable Computing = Nadezhnye vychisleniia, 5(1): 63–68, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1026493608620/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=1&spage=63;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=1&spage=63-68.

Lerch:1999:ETD

[390] Michael Lerch and Jürgen Wolff von Gudenberg. Expression templates for dot product expressions. Reliable Computing = Nadezhnye vychisleniia, 5(1):69–80, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026445725458/;   http://www.springerlink.com/openurl.

asp?genre=article&issn=1385-3139&volume=5&issue=1&spage=69;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=5&issue=1&spage=69-80.

Kosheleva:1999:EEI

[391] Olga Kosheleva and Vladik Kreinovich. Error estimation for indirect measurements: Interval computation problem is (slightly) harder than a similar probabilistic computational problem. Reliable Computing = Nadezhnye vychisleniia, 5(1):81–95, February 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026497709529/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=1&spage=81;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
1&spage=81-95.

Dimitrova:1999:SLS

[392] N. Dimitrova, S. Markov, and Ch. Ullrich. Solving linear systems with error control: Minisymposium within the 4th International Conference NMA’98. Reliable Computing = Nadezhnye vychisleniia, 5(1):97–100, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026449926367/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=1&spage=97;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=1&spage=97-100.

Csendes:1999:SIG

[393] Tibor Csendes. SCAN-98: an IMACS/GAMM International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics. Reliable Computing = Nadezhnye vychisleniia, 5(1):101–102, February 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1026402010437;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=5&issue=1&spage=101;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=1&spage=101-102.

Kolev:1999:IMG

[394] Lubomir V. Kolev. An improved method for global solution of nonlinear systems. Reliable Computing = Nadezhnye vychisleniia, 5(2):103– 111, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009935303343/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=2&spage=103;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=2&spage=103-111.

Zelawski:1999:RNA

[395] Marcin Zelawski. Rigorous numerical approach to isolation in dynamical systems on the example of the Kuramoto–Sivashinsky equation. Reliable Computing = Nadezhnye vychisleniia, 5(2):113–129, May 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009945220181/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=2&spage=113;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
2&spage=113-129.

Neumaier:1999:SDH

[396] Arnold Neumaier. A simple derivation of the Hansen–Bliek–Rohn– Ning–Kearfott enclosure for linear interval equations. Reliable Computing = Nadezhnye vychisleniia, 5(2):131–136, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009997221089/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=2&spage=131;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=

2&spage=131-136. See erratum [441].

Coxson:1999:CEB

[397] Gregory E. Coxson. Computing exact bounds on elements of an inverse interval matrix is NP-hard. Reliable Computing = Nadezhnye vychisleniia, 5 (2):137–142, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009901405160/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=2&spage=137;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=2&spage=137-142.

Anguelov:1999:WFI

[398] Roumen Anguelov. Wrapping function of the initial value problem for ODE: Applications. Reliable Computing = Nadezhnye vychisleniia, 5(2): 143–164, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/

A%3A1009953406068/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=2&spage=143;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=2&spage=143-164.

Jaulin:1999:EDE

[399] Luc Jaulin, Jean-Louis Boimond, and Laurent Hardouin. Estimation of discrete-event systems using interval computation. Reliable Computing = Nadezhnye vychisleniia, 5(2):165–173, May 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009957622907/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=2&spage=165;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
2&spage=165-173.

Kosheleva:1999:IES

[400] Olga Kosheleva, Sergio D. Cabrera, Glenn A. Gibson, and Sreedhar Cherukuri. Interval estimates for signal processing: Special purpose hardware. Reliable Computing = Nadezhnye vychisleniia, 5(2):175–196, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009909706977/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=2&spage=175;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=2&spage=175-196.

Anonymous:1999:ICH

[401] Anonymous. Interval computations help in proving the Kepler conjecture. Reliable Computing = Nadezhnye vychisleniia, 5(2):197–199, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017205323815/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=2&spage=197-199.

Longpre:1999:BRG

[402] Luc Longpré and Vladik Kreinovich. Book review: Gasarch, W. I. and Martin, G. A.: Bounded Queries in Recursion Theory. Reliable Computing = Nadezhnye vychisleniia, 5(2):201–203, May 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009961707886/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=2&spage=201;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
2&spage=201-203. See [779].

Garloff:1999:KNT

[403] Jürgen Garloff. Karl Nickel is turned 75. Reliable Computing = Nadezhnye vychisleniia, 5(2):205–206, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1009965824724;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=5&issue=2&spage=205;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=5&issue=2&spage=
205-206.

Vehi:1999:IWM

[404] Josep Vehí. Interval Workshop MISC’99. Reliable Computing = Nadezhnye vychisleniia, 5(2):207–208, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1009917908794;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=5&issue=2&spage=207;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=5&issue=2&spage=
207-208.

Corliss:1999:RBG

[405] George F. Corliss. Reminiscences “back in the good old days … ”. Reliable Computing = Nadezhnye vychisleniia, 5(2):209, May 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1009969925633;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=5&issue=2&spage=209.

Anonymous:1999:DC

[406] Anonymous. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 5(3):211–213, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017218110602/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=211-213.

Collavizza:1999:CPC

[407] Hélène Collavizza, François Delobel, and Michel Rueher. Comparing partial consistencies. Reliable Computing = Nadezhnye vychisleniia, 5(3):213– 228, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009922003700/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=3&spage=213;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=3&spage=213-228.

Dimitrova:1999:VCF

[408] Neli S. Dimitrova and Svetoslav M. Markov. Verified computation of fast decreasing polynomials. Reliable Computing = Nadezhnye vychisleniia, 5 (3):229–240, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009972120539/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=5&issue=3&spage=229;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=229-240.

Dyllong:1999:ADC

[409] Eva Dyllong, Wolfram Luther, and Werner Otten. An accurate distance-calculation algorithm for convex polyhedra. Reliable Computing = Nadezhnye vychisleniia, 5(3):241–253, August 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009924204609/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=241;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=241-253.

Frommer:1999:VEB

[410] Andreas Frommer and Andre Weinberg. Verified error bounds for linear systems through the Lanczos process. Reliable Computing = Nadezhnye vychisleniia, 5(3):255–267, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009976221447/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=255;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=255-267.

Heindl:1999:RIH

[411] Gerhard Heindl. A representation of the interval hull of a tolerance polyhedron describing inclusions of function values and slopes. Reliable Computing = Nadezhnye vychisleniia, 5(3):269–278, August 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009928406426/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=269;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=269-278.

Muller:1999:FRT

[412] Jean-Michel Muller. A few results on table-based methods. Reliable Computing = Nadezhnye vychisleniia, 5(3):279–288, August 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009984523264/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=279;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=279-288.

Nedialkov:1999:IHO

[413] Nedialko S. Nedialkov and Kenneth R. Jackson. An interval Hermite– Obreschkoff method for computing rigorous bounds on the solution of an initial value problem for an ordinary differential equation. Reliable Computing = Nadezhnye vychisleniia, 5(3):289–310, August 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009936607335/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=289;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=289-310.

Schulte:1999:IEG

[414] Michael J. Schulte, Vitaly Zelov, Ahmet Akkas, and James Craig C. Burley. The interval-enhanced GNU Fortran compiler. Reliable Computing = Nadezhnye vychisleniia, 5(3):311–322, August 1999. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009988620481/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=311;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=311-322.

Shary:1999:OEG

[415] Sergey P. Shary. Outer estimation of generalized solution sets to interval linear systems. Reliable Computing = Nadezhnye vychisleniia, 5(3):323– 335, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009972404551/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=3&spage=323;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=3&spage=323-335.

Strzebonski:1999:RPD

[416] Adam Strzebonski. A real polynomial decision algorithm using arbitrary-precision floating point arithmetic. Reliable Computing = Nadezhnye vychisleniia, 5(3):337–346, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009924521389/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=337;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=337-346.

Watanabe:1999:NVM

[417] Yoshitaka Watanabe, Nobito Yamamoto, and Mitsuhiro T. Nakao. A numerical verification method of solutions for the Navier–Stokes equations. Reliable Computing = Nadezhnye vychisleniia, 5(3):347–357, August 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009976505460/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=3&spage=347;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
3&spage=347-357.

Min:1999:CKM

[418] Kong Min, Liu Qi, and Shen Zuhe. On the componentwise Krawczyk– Moore iteration. Reliable Computing = Nadezhnye vychisleniia, 5(4):359– 370, November 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1009928629577/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=4&spage=359;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=4&spage=359-370.

Wolfe:1999:DMP

[419] Michael A. Wolfe. On discrete minimax problems in R using interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 5(4):371–383, November 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009930013647/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=4&spage=371;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=4&spage=371-383.

Kosheleva:1999:OIP

[420] Olga Kosheleva and Vladik Kreinovich. Only intervals preserve the invertibility of arithmetic operations. Reliable Computing = Nadezhnye vychisleniia, 5(4):385–394, November 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009982014556/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=4&spage=385;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
4&spage=385-394.

Kalovics:1999:SNC

[421] Ferenc Kálovics. Solving nonlinear constrained minimization problems with a new interval valued function. Reliable Computing = Nadezhnye vychisleniia, 5(4):395–406, November 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009986131394/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=5&issue=4&spage=395;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=5&issue=
4&spage=395-406.

Nataraj:1999:SRS

[422] Paluri S. V. Nataraj and Suresh Srivastava. Synthesis of robustly stabilizing general order compensators for interval plants using interval analysis. Reliable Computing = Nadezhnye vychisleniia, 5(4):407–422, November 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009938215464/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=5&issue=4&spage=407;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=5&issue=4&spage=407-422.

Moore:1999:D

[423] R. E. Moore. The dawning. Reliable Computing = Nadezhnye vychisleniia, 5(4):423–424, November 1999. CODEN RCOMF8.

ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017250200443;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=5&issue=4&spage=423;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=5&issue=4&spage=
423-424.

Ratschek:2000:ISI

[424] Helmut Ratschek and Jon G. Rokne. Introduction to the special issue: What can one learn from box-plane intersections? Reliable Computing = Nadezhnye vychisleniia, 6(1):1–8, February 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009984921053/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=1&spage=1;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
1&spage=1-8.

Caprani:2000:RER

[425] Ole Caprani, Lars Hvidegaard, Mikkel Mortensen, and Thomas Schneider. Robust and efficient ray intersection of implicit surfaces. Reliable Computing = Nadezhnye vychisleniia, 6(1):9–21, February 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009921806032/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=1&spage=9;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
1&spage=9-21.

Ely:2000:CDT

[426] Jeffrey S. Ely and Anthony P. Leclerc. Correct Delaunay triangulation in the presence of inexact inputs and arithmetic. Reliable Computing = Nadezhnye vychisleniia, 6(1):23–38, February 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009977923779/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=1&spage=23;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
1&spage=23-38.

Gavrilova:2000:ECD

[427] Marina Gavrilova, Helmut Ratschek, and Jon G. Rokne. Exact computation of Delaunay and Power triangulations. Reliable Computing = Nadezhnye vychisleniia, 6(1):39–60, February 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009934225596/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=1&spage=39;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
1&spage=39-60.

Alliez:2000:RDP

[428] Pierre Alliez, Olivier Devillers, and Jack Snoeyink. Removing degeneracies by perturbing the problem or perturbing the world. Reliable Computing = Nadezhnye vychisleniia, 6(1):61–79, February 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009942427413/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=1&spage=61;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
1&spage=61-79.

Hickey:2000:ICP

[429] Timothy J. Hickey, Zhe Qju, and Maarten H. Van Emden. Interval constraint plotting for interactive visual exploration of implicitly defined relations. Reliable Computing = Nadezhnye vychisleniia, 6(1):81–92, February 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009950630139/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=6&issue=1&spage=81;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=6&issue=1&spage=81-92.

Rall:2000:II

[430] L. B. Rall. Interval integration. Reliable Computing = Nadezhnye vychisleniia, 6(1):93–94, February 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017235301006;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=6&issue=1&spage=93;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=6&issue=1&spage=
93-94.

Hansen:2000:HPI

[431] Eldon R. Hansen. The hull of preconditioned interval linear equations. Reliable Computing = Nadezhnye vychisleniia, 6(2):95–103, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009962903365/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=2&spage=95;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=95-103.

Yamamura:2000:FAS

[432] Kiyotaka Yamamura. Finding all solutions of nonlinear equations using linear combinations of functions. Reliable Computing = Nadezhnye vychisleniia, 6(2):105–113, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009956920204/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=105;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=2&spage=105-113.

Hungerbuhler:2000:CBC

[433] Ralf Hungerbühler and Jürgen Garloff. Computation of the Bernstein coefficients on subdivided triangles. Reliable Computing = Nadezhnye vychisleniia, 6(2):115–121, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009909004274/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=115;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=2&spage=115-121.

Stipanovic:2000:JGS

[434] Dusan M. Stipanović and Dragoslav D. Siljak. Jacobi and Gauss–Seidel iterations for polytopic systems: Convergence via convex M-matrices. Reliable Computing = Nadezhnye vychisleniia, 6(2):123–137, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009961021112/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=2&spage=123;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=123-137.

Dimuro:2000:CSR

[435] Graçaliz P. Dimuro, Antônio Carlos Da R. Costa, and Dalcidio M. Claudio. A coherence space of rational intervals for a construction of IR. Reliable Computing = Nadezhnye vychisleniia, 6(2):139–178, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009913122021/;

 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=2&spage=139;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=139-178.

Casado:2000:IBB

[436] Leocadio G. Casado, Inmaculada F. García, and Yaroslav D. Sergeyev. Interval branch and bound algorithm for finding the first-zero-crossing-point in one-dimensional functions. Reliable Computing = Nadezhnye vychisleniia, 6(2):179–191, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009917222929/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=179;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=2&spage=179-191.

Revol:2000:ASA

[437] Nathalie Revol and Jean-Claude Yakoubsohn. Accelerated shift-and-add algorithms. Reliable Computing = Nadezhnye vychisleniia, 6(2):193– 205, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009921407000/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=6&issue=2&spage=193;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=6&issue=2&spage=193-205.

Hofschuster:2000:MFS

[438] Werner Hofschuster and Walter Krämer. Mathematical function software on the Web — are such codes useful for verification algorithms? Reliable Computing = Nadezhnye vychisleniia, 6(2):207–218, May 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009973407908/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=2&spage=207;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=207-218.

Koshelev:2000:ESF

[439] Misha Koshelev. Every superinterval of the function range can be an interval-computations enclosure. Reliable Computing = Nadezhnye vychisleniia, 6(2):219–223, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009977508817/;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=219;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=2&spage=219-223.

Hansen:2000:R

[440] Eldon R. Hansen. Reminiscences. Reliable Computing = Nadezhnye vychisleniia, 6(2):225–226, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1023/A%3A1009981625655;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
2&spage=225;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=2&spage=225-226.

Neumaier:2000:ESD

[441] Arnold Neumaier. Erratum to: “A Simple Derivation of the Hansen– Bliek–Rohn–Ning–Kearfott Enclosure for Linear Interval Equations” (Reliable Computing Vol. 5 (2) (1999)). Reliable Computing = Nadezhnye vychisleniia, 6(2):227, May 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1009933709726;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=6&issue=2&spage=227. See [396].

Garloff:2000:FDC

[442] Jürgen Garloff and Éric Walter. Foreword — dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 6(3):229–230, August 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1009985711668;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=3&spage=229;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=6&issue=3&spage=229-230.

Wang:2000:CIC

[443] Long Wang. Composite interval control systems: Some strong Kharitonovlike properties. Reliable Computing = Nadezhnye vychisleniia, 6(3):231– 246, August 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009974328506/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=6&issue=3&spage=231;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=6&issue=3&spage=231-246.

Ohta:2000:NPI

[444] Yuzo Ohta. Nonconvex polygon interval arithmetic as a tool for the analysis and design of robust control systems. Reliable Computing = Nadezhnye vychisleniia, 6(3):247–279, August 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009926413485/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=3&spage=247;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
3&spage=247-279.

Vehi:2000:ARP

[445] Josep Vehí, José Rodellar, Miguel Sainz, and Joaquim Armengol. Analysis of the robustness of predictive controllers via modal intervals. Reliable Computing = Nadezhnye vychisleniia, 6(3):281–301, August 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009982530323/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=3&spage=281;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
3&spage=281-301.

Garloff:2000:ABE

[446] Jürgen Garloff. Application of Bernstein expansion to the solution of control problems. Reliable Computing = Nadezhnye vychisleniia, 6(3):303– 320, August 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1009934614393/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=6&issue=3&spage=303;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=6&issue=3&spage=303-320.

Edmonson:2000:IMS

[447] William W. Edmonson, Wen H. Lee, and John M. M. Anderson. Interval methods for sinusoidal parameter estimation: a comparative analysis. Reliable Computing = Nadezhnye vychisleniia, 6(3):321–336, August 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009986615302/;
 http://www.springerlink.com/content/p4t219447v056l00/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=6&issue=3&spage=321;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=6&issue=3&spage=321-

Kieffer:2000:RAR

[448] Michel Kieffer, Luc Jaulin, Éric Walter, and Dominique Meizel. Robust autonomous robot localization using interval analysis. Reliable Computing = Nadezhnye vychisleniia, 6(3):337–362, August 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009990700281/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=3&spage=337;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
3&spage=337-362.

Verdonk:2000:DC

[449] Brigitte Verdonk. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 6(4):363, November 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1009946905794;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=6&issue=4&spage=363.

Beckermann:2000:ECR

[450] Bernhard Beckermann and George Labahn. Effective computation of rational approximants and interpolants. Reliable Computing = Nadezhnye vychisleniia, 6(4):365–390, November 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009942122633/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=4&spage=365;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
4&spage=365-390.

Fournier:2000:RIS

[451] Jean-Daniel Fournier and Maciej Pindor. Rational interpolation from stochastic data: a new Froissart’s phenomenon. Reliable Computing = Nadezhnye vychisleniia, 6(4):391–409, November 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009994123541/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=4&spage=391;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
4&spage=391-409.

Graves-Morris:2000:RLT

[452] Peter R. Graves-Morris. Reliability of Lanczos-type product methods from perturbation theory. Reliable Computing = Nadezhnye vychisleniia, 6(4): 411–428, November 2000. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009998225359/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=6&issue=4&spage=411;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=4&spage=411-428.

Kai:2000:HRF

[453] Hiroshi Kai and Matu-Tarow Noda. Hybrid rational function approximation and its accuracy analysis. Reliable Computing = Nadezhnye vychisleniia, 6(4):429–438, November 2000. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1009906513972/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
4&spage=429;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=6&issue=4&spage=429-438.

Karcanias:2000:NCL

[454] Nicos Karcanias and Marilena Mitrouli. Numerical computation of the least common multiple of a set of polynomials. Reliable Computing = Nadezhnye vychisleniia, 6(4):439–457, November 2000. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1009979130810/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=4&spage=439;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=6&issue=
4&spage=439-457.

Sakurai:2000:FAF

[455] Tetsuya Sakurai and Hiroshi Sugiura. On factorization of analytic functions and its verification. Reliable Computing = Nadezhnye vychisleniia, 6 (4):459–470, November 2000. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1009931231719/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=6&issue=4&spage=459;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=6&issue=4&spage=459-470.

Jaulin:2001:PPU

[456] Luc Jaulin. Path planning using intervals and graphs. Reliable Computing = Nadezhnye vychisleniia, 7(1):1–15, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011400431065/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=1&spage=1;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
1&spage=1-15.

Kolev:2001:ACL

[457] Lubomir V. Kolev. Automatic computation of a linear interval enclosure. Reliable Computing = Nadezhnye vychisleniia, 7(1):17–28, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011470916953/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=1&spage=17;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
1&spage=17-28.

Dobner:2001:KI

[458] Hans-Jürgen Dobner. On kernel inclusions. Reliable Computing = Nadezhnye vychisleniia, 7(1):29–39, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011431217861/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=1&spage=29;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
1&spage=29-39.

Kreinovich:2001:RFN

[459] Vladik Kreinovich. Roundoff-free number fields for interval computations. Reliable Computing = Nadezhnye vychisleniia, 7(1):41–47, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011483201932/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=1&spage=41;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
1&spage=41-47.

Wolfe:2001:NUT

[460] Michael A. Wolfe. A note on a uniqueness theorem for the secondderivative test of Qi. Reliable Computing = Nadezhnye vychisleniia, 7(1):

49–52, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1011435318770/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=1&spage=49;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=1&spage=49-52.

Menshikov:2001:SSR

[461] Gregory G. Menshikov and Alexey V. Tomashevsky. On ∧-subdistributivity and ∨-superdistributivity with respect to interval map in Kaucher arithmetic. Reliable Computing = Nadezhnye vychisleniia, 7(1):53–57, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1011487302840/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=1&spage=53;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=1&spage=53-57.

Schafer:2001:FIG

[462] Uwe Schäfer. The feasibility of the interval Gaussian algorithm for arrowhead matrices. Reliable Computing = Nadezhnye vychisleniia, 7(1):59–62, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1011439403749/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=1&spage=59;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=1&spage=59-62.

Markov:2001:MI

[463] Svetoslav Markov. The mystery of intervals. Reliable Computing = Nadezhnye vychisleniia, 7(1):63–65, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011451822404/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=1&spage=63;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
1&spage=63-65.

Starks:2001:IMT

[464] Scott A. Starks and Vladik Kreinovich. Interval methods and their applications: Invited sessions at the World Automation Congress (WAC’2000) Maui, Hawaii, June 10–16, 2000. Reliable Computing = Nadezhnye vychisleniia, 7(1):67–68, February 2001. CODEN RCOMF8.

ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1011403908292;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=7&issue=1&spage=67;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=7&issue=1&spage=
67-68.

Nesterov:2001:SIC

[465] Vyacheslav Nesterov and Vladik Kreinovich. Session “Interval and Computer–Algebraic Methods in Science and Engineering” at the 6th International Association for Mathematics and Computers in Simulation (IMACS) Conference on Applications of Computer Algebra (ACA’2000) Saint Petersburg, Russia, June 25–28, 2000. Reliable Computing = Nadezhnye vychisleniia, 7(1):69–70, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1011416209201;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=7&issue=1&spage=69;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=7&issue=1&spage=
69-70.

Lin:2001:SSG

[466] T. Y. Lin and V. Kreinovich. A Special Session on Granular Computing and Interval Computations at the 19th International Conference of the North American Fuzzy Information Processing Society (NAFIPS) Atlanta, Georgia, July 13–15, 2000. Reliable Computing = Nadezhnye vychisleniia, 7(1):71–72, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1011468210109;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=7&issue=1&spage=71;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=7&issue=1&spage=
71-72.

Garloff:2001:MAI

[467] Jürgen Garloff. Minisymposium on Applications of Interval Computations at the Third World Congress of Nonlinear Analysts Catania, Sicily, Italy, July 19–26, 2000. Reliable Computing = Nadezhnye vychisleniia, 7(1):73–74, February 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1011472311018;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=7&issue=1&spage=73;   http://www.springerlink.com/

openurl.asp?genre=article&issn=1385-3139&volume=7&issue=1&spage=
73-74.

Vehi:2001:E

[468] Josep Vehí and Miguel Á. Sainz. Editorial. Reliable Computing = Nadezhnye vychisleniia, 7(2):75, April 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1011428913339;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=7&issue=2&spage=75.

Gardenes:2001:MI

[469] Ernest Gardeñes, Miguel Á. Sainz, Lambert Jorba, Remei Calm, Rosa Estela, Honorino Mielgo, and Albert Trepat. Modal intervals. Reliable Computing = Nadezhnye vychisleniia, 7(2):77–111, April 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011465930178/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=2&spage=77;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
2&spage=77-111.

Markov:2001:API

[470] Svetoslav Markov. On the algebraic properties of intervals and some applications. Reliable Computing = Nadezhnye vychisleniia, 7(2):113– 127, April 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1011418014248/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=2&spage=113;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=2&spage=113-127.

Popova:2001:MDP

[471] Evgenija D. Popova. Multiplication distributivity of proper and improper intervals. Reliable Computing = Nadezhnye vychisleniia, 7(2):129–140, April 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1011470131086/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=2&spage=129;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=2&spage=129-140.

Shary:2001:IGS

[472] Sergey P. Shary. Interval Gauss–Seidel method for generalized solution sets to interval linear systems. Reliable Computing = Nadezhnye vychisleniia, 7(2):141–155, April 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1011422215157/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
2&spage=141;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=2&spage=141-155.

Hertling:2001:LUT

[473] Peter Hertling. A limitation for underestimation via twin arithmetic. Reliable Computing = Nadezhnye vychisleniia, 7(2):157–169, April 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011474231995/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=2&spage=157;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
2&spage=157-169.

Armengol:2001:AMI

[474] Joaquim Armengol, Josep Vehí, Louise Travé-Massuyès, and Miguel Ángel Sainz. Application of modal intervals to the generation of error-bounded envelopes. Reliable Computing = Nadezhnye vychisleniia, 7(2):171–185, April 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1011426300135/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=2&spage=171;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=2&spage=171-185.

Anonymous:2001:SGI

[475] Anonymous. SCAN 2000: GAMM–IMACS International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics Interval 2000: International Conference on Interval Methods in Science and Engineering Karlsruhe, Germany, September 18–22, 2000. Reliable Computing = Nadezhnye vychisleniia, 7(2):187–191, April 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017363016974/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=2&spage=187;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
2&spage=187-191.

Anonymous:2001:DC

[476] Anonymous.  Dear colleagues.  Reliable Computing = Nadezh-

nye vychisleniia, 7(3):193–194, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017382418791;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=7&issue=3&spage=193-194.

Wolfe:2001:BSU

[477] Michael A. Wolfe. On bounding solutions of underdetermined systems. Reliable Computing = Nadezhnye vychisleniia, 7(3):195–207, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011430402861/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=3&spage=195;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
3&spage=195-207.

Sweidan:2001:TEC

[478] Andraos Sweidan and Ahmad A. Hiasat. On the theory of error control based on moduli with common factors. Reliable Computing = Nadezhnye vychisleniia, 7(3):209–218, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011442703770/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=3&spage=209;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
3&spage=209-218.

Shashikhin:2001:RCU

[479] Vladimir N. Shashikhin. Robust control using interval analysis. Reliable Computing = Nadezhnye vychisleniia, 7(3):219–230, June 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011446804678/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=3&spage=219;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
3&spage=219-230.

Jaulin:2001:RMP

[480] Luc Jaulin. Reliable minimax parameter estimation. Reliable Computing = Nadezhnye vychisleniia, 7(3):231–246, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011451021517/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=3&spage=231;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
3&spage=231-246.

Do:2001:SAR

[481] Dang-Khoa Do. Spigot algorithm and root computing. Reliable Computing = Nadezhnye vychisleniia, 7(3):247–273, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011403105587/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=3&spage=247;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
3&spage=247-273.

Alefeld:2001:CSS

[482] Götz Alefeld, Vladik Kreinovich, Günter Mayer, and Michael Huth. A comment on the shape of the solution set for systems of interval linear equations with dependent coefficients. Reliable Computing = Nadezhnye vychisleniia, 7(3):275–277, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011455122425/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=3&spage=275;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
3&spage=275-277.

Anonymous:2001:ITI

[483] Anonymous. Interval talks at the International Conference on Intelligent Technologies InTech’2000 Bangkok, Thailand, December 12–14, 2000. Reliable Computing = Nadezhnye vychisleniia, 7(3):279–280, June 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1017390606496;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=3&spage=279;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=3&spage=279-280.

Schwandt:2001:SAI

[484] Hartmut Schwandt. Synchronous and asynchronous interval Newton– Schwarz methods for a class of large systems of nonlinear equations. Reliable Computing = Nadezhnye vychisleniia, 7(4):281–306, August 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011407223334/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=4&spage=281;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
4&spage=281-306.

Neher:2001:VBT

[485] Markus Neher. Validated bounds for Taylor coefficients of analytic functions. Reliable Computing = Nadezhnye vychisleniia, 7(4):307–319, August 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1011411307404/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=4&spage=307;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=4&spage=307-319.

Kramer:2001:AFE

[486] Walter Krämer and Armin Bantle. Automatic forward error analysis for floating point algorithms. Reliable Computing = Nadezhnye vychisleniia, 7(4):321–340, August 2001. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1011463324243/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=7&issue=4&spage=321;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=4&spage=321-340.

Worden:2001:IMN

[487] Keith Worden, Roberto Osegueda, Carlos Ferregut, Soheil Nazarian, Debra L. George, V. Kreinovich, O. Kosheleva, and S. Cabrera. Interval methods in non-destructive testing of material structures. Reliable Computing = Nadezhnye vychisleniia, 7(4):341–352, August 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011415408313/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=4&spage=341;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
4&spage=341-352.

Litvinov:2001:IIA

[488] Grigori L. Litvinov and Andrei N. Sobolevskii. Idempotent interval analysis and optimization problems. Reliable Computing = Nadezhnye vychisleniia, 7(5):353–377, October 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1011487725803/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
5&spage=353;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=5&spage=353-377.

Berz:2001:VHO

[489] Martin Berz and Jens Hoefkens. Verified high-order inversion of functional dependencies and interval Newton methods. Reliable Computing = Nadezhnye vychisleniia, 7(5):379–398, October 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011423909873/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=5&spage=379;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
5&spage=379-398.

Kolev:2001:CTB

[490] Lubomir V. Kolev and Ivo P. Nenov. Cheap and tight bounds on the solution set of perturbed systems of nonlinear equations. Reliable Computing = Nadezhnye vychisleniia, 7(5):399–408, October 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011475926711/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=5&spage=399;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
5&spage=399-408.

Sharaya:2001:MIE

[491] Irene A. Sharaya. On maximal inner estimation of the solution sets of linear systems with interval parameters. Reliable Computing = Nadezhnye vychisleniia, 7(5):409–424, October 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011428127620/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=5&spage=409;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
5&spage=409-424.

Kearfott:2001:ESN

[492] R. Baker Kearfott. An example of singularity in nonlinear systems. Reliable Computing = Nadezhnye vychisleniia, 7(5):425–429, October 2001.

CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1011484228528/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=5&spage=425;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
5&spage=425-429.

Starks:2001:CS

[493] Scott A. Starks and Vladik Kreinovich. Claude E. Shannon (1916-2001). Reliable Computing = Nadezhnye vychisleniia, 7(5):431–432, October 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1011436312599;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=5&spage=431;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=5&spage=431-432.

Beaumont:2001:LIT

[494] Oliver Beaumont and Bernard Philippe. Linear interval tolerance problem and linear programming techniques. Reliable Computing = Nadezhnye vychisleniia, 7(6):433–447, December 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014758201565/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=6&spage=433;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
6&spage=433-447.

Nedialkov:2001:EHO

[495] Nedialko S. Nedialkov, Kenneth R. Jackson, and John D. Pryce. An effective high-order interval method for validating existence and uniqueness of the solution of an IVP for an ODE. Reliable Computing = Nadezhnye vychisleniia, 7(6):449–465, December 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014798618404/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=6&spage=449;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
6&spage=449-465.

Granvilliers:2001:CIC

[496] Laurent Granvilliers. On the combination of interval constraint solvers. Reliable Computing = Nadezhnye vychisleniia, 7(6):467–483, December 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014750702474/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=7&issue=6&spage=467;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=6&spage=467-483.

Menshikov:2001:PEO

[497] Gregory G. Menshikov. The preliminary enclosing of the ODE solutions on the base of the Cauchy–Duhamel identity. Reliable Computing = Nadezhnye vychisleniia, 7(6):485–495, December 2001. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014702819312/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=6&spage=485;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
6&spage=485-495.

Shary:2001:SAI

[498] Sergey P. Shary. A surprising approach in interval global optimization. Reliable Computing = Nadezhnye vychisleniia, 7(6):497–505, December 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014754803382/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=7&issue=6&spage=497;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=

6&spage=497-505. Dedicated to Prof. Dr. Gregory G. Menshikov on the occasion of his 70th anniversary.

Dobronets:2001:GGM

[499] Boris S. Dobronets, Anatoly V. Lakeyev, Vyacheslav M. Nesterov, Sergey P. Shary, and Yuri I. Shokin. Gregory G. Menshikov has turned 70. Reliable Computing = Nadezhnye vychisleniia, 7(6):507–508, December 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1014706920221;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=6&spage=507;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=7&issue=6&spage=507-508.

Anonymous:2001:GII

[500] Anonymous. 10th GAMM–IMACS International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics. Reliable Computing = Nadezhnye vychisleniia, 7(6):509–510, December 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/A%
3A1017476504291;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=6&spage=509-510.

Anonymous:2001:RC

[501] Anonymous. Reliable computing. Reliable Computing = Nadezhnye vychisleniia, 7(6):511–515, December 2001. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017489810284/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=7&issue=
6&spage=511;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=7&issue=6&spage=511-515.

Schafer:2002:TWE

[502] Uwe Schäfer. Two ways to extend the Cholesky decomposition to block matrices with interval entries. Reliable Computing = Nadezhnye vychisleniia, 8(1):1–20, February 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1014717701732/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
1&spage=1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=8&issue=1&spage=1-20.

Ratschan:2002:AQC

[503] Stefan Ratschan. Approximate quantified constraint solving by cylindrical box decomposition. Reliable Computing = Nadezhnye vychisleniia, 8(1): 21–42, February 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1014785518570/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=1&spage=21;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=1&spage=21-42.

Farouki:2002:EMP

[504] Rida T. Farouki and Helmut Pottmann. Exact Minkowski products of N complex disks. Reliable Computing = Nadezhnye vychisleniia, 8(1):43–66, February 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014737602641/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=1&spage=43;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=1&spage=43-66.

Hoefkens:2002:VIC

[505] Jens Hoefkens and Martin Berz. Verification of invertibility of complicated functions over large domains. Reliable Computing = Nadezhnye vychisleniia, 8(1):67–82, February 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1014789619479/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
1&spage=67;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=8&issue=1&spage=67-82.

Nataraj:2002:NSS

[506] Paluri S. V. Nataraj and Suresh Mandir Sheela. A new subdivision strategy for range computations. Reliable Computing = Nadezhnye vychisleniia, 8(1):83–92, February 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1014741703549/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
1&spage=83;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=8&issue=1&spage=83-92.

Lodwick:2002:RCS

[507] Weldon A. Lodwick. Reliable computing: Special issue on the linkages between interval mathematics and fuzzy set theory. Reliable Computing = Nadezhnye vychisleniia, 8(1):93–95, February 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014793820388/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=1&spage=93;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=

1&spage=93-95. Special Issue on the Linkages between Interval Mathematics and Fuzzy Set Theory.

Kozine:2002:IVF

[508] Igor O. Kozine and Lev V. Utkin. Interval-valued finite Markov chains. Reliable Computing = Nadezhnye vychisleniia, 8(2):97–113, April 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014745904458/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=2&spage=97;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
2&spage=97-113.

Hansen:2002:SBI

[509] Eldon R. Hansen and G. William Walster. Sharp bounds on interval polynomial roots. Reliable Computing = Nadezhnye vychisleniia, 8(2):115– 122, April 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014797921296/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=2&spage=115;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=2&spage=115-122.

Nataraj:2002:PVC

[510] Paluri S. V. Nataraj and Airani Kalathil Prakash. A parallelized version of the covering algorithm for solving parameter-dependent systems of nonlinear equations. Reliable Computing = Nadezhnye vychisleniia, 8(2):123– 130, April 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014750005366/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=2&spage=123;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=2&spage=123-130.

Oliveira:2002:ICV

[511] Jaão Batista Oliveira and Luiz Henrique De Figueiredo. Interval computation of Viswanath’s constant. Reliable Computing = Nadezhnye vychisleniia, 8(2):131–138, April 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1014702122205/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
2&spage=131;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=8&issue=2&spage=131-138.

Chiu:2002:EIL

[512] Chong-Kan Chiu and Jimmy Ho-Man Lee. Efficient interval linear equality solving in constraint logic programming. Reliable Computing = Nadezhnye vychisleniia, 8(2):139–174, April 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1014754106275/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=2&spage=139;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
2&spage=139-174.

Berleant:2002:RCS

[513] Daniel Berleant. Reliable computing: Special issue on dependable reasoning about uncertainty. Reliable Computing = Nadezhnye vychisleniia, 8(2):175–176, April 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1014706223113;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=8&issue=2&spage=175;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=8&issue=2&spage=

175-176. Special Issue on Dependable Reasoning about Uncertainty.

Wolfe:2002:BPZ

[514] Michael A. Wolfe. Bounding perturbations in zeros of nonlinear systems. Reliable Computing = Nadezhnye vychisleniia, 8(3):177–188, June 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1015559628657/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=3&spage=177;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
3&spage=177-188.

Sainz:2002:FSS

[515] Miguel Á. Sainz, Ernest Gardeñes, and Lambert Jorba. Formal solution to systems of interval linear or non-linear equations. Reliable Computing = Nadezhnye vychisleniia, 8(3):189–211, June 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1015561212728/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=3&spage=189;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
3&spage=189-211.

Heindl:2002:CIM

[516] Gerhard Heindl, Vladik Kreinovich, and Maria Rifqi. In case of interval (or more general) uncertainty, no algorithm can choose the simplest representative. Reliable Computing = Nadezhnye vychisleniia, 8(3):213– 227, June 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1015513329566/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=3&spage=213;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=3&spage=213-227.

Mayer:2002:AOD

[517] Jan Mayer. An approach to overcome division by zero in the interval Gauss algorithm. Reliable Computing = Nadezhnye vychisleniia, 8(3):229– 237, June 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1015565313636/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=3&spage=229;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=3&spage=229-237.

Johnson:2002:IIM

[518] Charles R. Johnson and Ronald L. Smith. Intervals of inverse M-matrices. Reliable Computing = Nadezhnye vychisleniia, 8(3):239–243, June 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1015517430475/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=3&spage=239;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
3&spage=239-243.

Loh:2002:RER

[519] Eugene Loh and G. William Walster. Rump’s example revisited. Reliable Computing = Nadezhnye vychisleniia, 8(3):245–248, June 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1015569431383/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=3&spage=245;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
3&spage=245-248.

Moore:2002:SSF

[520] Ramon E. Moore. Sparse systems in fixed point form. Reliable Computing = Nadezhnye vychisleniia, 8(4):249–265, August 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1016390830247/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=4&spage=249;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
4&spage=249-265.

Kearfott:2002:EUV

[521] R. Baker Kearfott. On existence and uniqueness verification for nonsmooth functions. Reliable Computing = Nadezhnye vychisleniia, 8(4): 267–282, August 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1016381031155/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=4&spage=267;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=4&spage=267-282.

Sainz:2002:IES

[522] Miguel Á. Sainz, Ernest Gardeñes, and Lambert Jorba. Interval estimations of solution sets to real-valued systems of linear or non-linear equations. Reliable Computing = Nadezhnye vychisleniia, 8(4):283–305, August 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1016385132064/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=4&spage=283;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=4&spage=283-305.

vonGudenberg:2002:IAM

[523] Jürgen Wolff von Gudenberg. Interval arithmetic on multimedia architectures. Reliable Computing = Nadezhnye vychisleniia, 8(4):307–312, August 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1016389216134/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=4&spage=307;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=4&spage=307-312.

Neumaier:2002:GCS

[524] Arnold Neumaier. Grand challenges and scientific standards in interval analysis. Reliable Computing = Nadezhnye vychisleniia, 8(4):313–320, August 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1016341317043/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=4&spage=313;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=4&spage=313-320.

Shary:2002:NTS

[525] Sergey P. Shary. A new technique in systems analysis under interval uncertainty and ambiguity. Reliable Computing = Nadezhnye vychisleniia, 8(5): 321–418, October 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1020505620702/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=5&spage=321;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=5&spage=321-418.

Kearfott:2002:SCO

[526] R. Baker Kearfott and G. William Walster. SIAM conference on optimization, validated computing 2002, and the fields institute informal working group on validated optimization. Reliable Computing = Nadezhnye vychisleniia, 8(5):419–424, October 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1020583804772/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=5&spage=419;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
5&spage=419-424.

Lodwick:2002:IRT

[527] Weldon A. Lodwick and Vladik Kreinovich. Interval-related talks at the Annual Conference of the North American Fuzzy Information Processing Society NAFIPS’02. Reliable Computing = Nadezhnye vychisleniia, 8(5):425–426, October 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1020535905681;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=8&issue=5&spage=425;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=8&issue=5&spage=
425-426.

Nataraj:2002:QPG

[528] Paluri S. V. Nataraj and Suresh Mandir Sheela. A QFT procedure for generating design frequencies and bounds of guaranteed accuracy. Reliable Computing = Nadezhnye vychisleniia, 8(6):427–451, December 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021331309574/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=6&spage=427;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
6&spage=427-451.

Kearfott:2002:SPT

[529] R. Baker Kearfott and G. William Walster. Symbolic preconditioning with Taylor models: Some examples. Reliable Computing = Nadezhnye vychisleniia, 8(6):453–468, December 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021364526413/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=6&spage=453;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
6&spage=453-468.

Dobner:2002:KST

[530] Hans-Jürgen Dobner. Kernel-splitting technique for enclosing the solution of Fredholm equations of the first kind. Reliable Computing = Nadezhnye vychisleniia, 8(6):469–479, December 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021316610483/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=6&spage=469;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
6&spage=469-479.

Kreinovich:2002:REN

[531] Vladik Kreinovich. Range estimation is NP-hard for ϵ^{2} accuracy and feasible for ϵ^{2−δ}. Reliable Computing = Nadezhnye vychisleniia, 8(6):481–491, December 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1021368627321/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=8&issue=6&spage=481;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=8&issue=6&spage=481-491.

Kolev:2002:OSL

[532] Lubomir V. Kolev. Outer solution of linear systems whose elements are affine functions of interval parameters. Reliable Computing = Nadezhnye vychisleniia, 8(6):493–501, December 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021320711392/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=6&spage=493;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
6&spage=493-501.

Oberkampf:2002:EUW

[533] William Oberkampf, Jon Helton, Steve Wojtkiewicz, Cliff Joslyn, and Scott Ferson. Epistemic uncertainty workshop. Reliable Computing = Nadezhnye vychisleniia, 8(6):503–505, December 2002. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021372829139/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=8&issue=6&spage=503;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
6&spage=503-505.

Anonymous:2002:CV

[534] Anonymous. Contents volume 8. Reliable Computing = Nadezhnye vychisleniia, 8(6):507–510, December 2002. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1021350118293/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=8&issue=
6&spage=507;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=8&issue=6&spage=507-510.

Kulpa:2003:DAIa

[535] Zenon Kulpa. Diagrammatic analysis of interval linear equations: Part I: Basic notions and the one-dimensional case. Reliable Computing = Nadezhnye vychisleniia, 9(1):1–20, February 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023022826879/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=1&spage=1;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
1&spage=1-20.

Hoefkens:2003:CWE

[536] Jens Hoefkens, Martin Berz, and Kyoko Makino. Controlling the wrapping effect in the solution of ODEs for asteroids. Reliable Computing = Nadezhnye vychisleniia, 9(1):21–41, February 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023009910949/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=1&spage=21;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
1&spage=21-41.

Neumaier:2003:TFU

[537] Arnold Neumaier. Taylor forms — use and limits. Reliable Computing = Nadezhnye vychisleniia, 9(1):43–79, February 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023061927787/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=1&spage=43;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
1&spage=43-79.

Kearfott:2003:CWG

[538] R. Baker Kearfott. COCOS’02 — A Workshop on Global Constrained Optimization and Constraint Satisfaction October 2–4, 2002, Sophia-Antipolis, France. Reliable Computing = Nadezhnye vychisleniia, 9(1):81– 87, February 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023014011858/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=1&spage=81;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=1&spage=81-87.

Kearfott:2003:DCa

[539] R. Baker Kearfott. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 9(2):89–90, April 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1023/A%3A1023022731988;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=89;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=9&issue=2&spage=89-90.

Berleant:2003:STD

[540] Daniel Berleant, Lizhi Xie, and Jianzhong Zhang. Statool: a tool for distribution envelope determination (DEnv), an interval-based algorithm for arithmetic on random variables. Reliable Computing = Nadezhnye vychisleniia, 9(2):91–108, April 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023082100128/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=2&spage=91;   http://www.springerlink.

com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=91-108.

Csendes:2003:NEN

[541] Tibor Csendes. Numerical experiences with a new generalized subinterval selection criterion for interval global optimization. Reliable Computing = Nadezhnye vychisleniia, 9(2):109–125, April 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023086201037/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=2&spage=109;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=109-125.

Lodwick:2003:EVC

[542] Weldon A. Lodwick and K. David Jamison. Estimating and validating the cumulative distribution of a function of random variables: Toward the development of distribution arithmetic. Reliable Computing = Nadezhnye vychisleniia, 9(2):127–141, April 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023090317875/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=2&spage=127;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=127-141.

Nenov:2003:IMA

[543] Ivo P. Nenov and Daniel H. Fylstra. Interval methods for accelerated global search in the Microsoft Excel Solver. Reliable Computing = Nadezhnye vychisleniia, 9(2):143–159, April 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023042418784/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=2&spage=143;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=143-159.

Oliveira:2003:RAO

[544] João Batista Oliveira and Luiz Henrique De Figueiredo. Robust approximation of offsets, bisectors, and medial axes of plane curves. Reliable Computing = Nadezhnye vychisleniia, 9(2):161–175, April 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023046502854/;

 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=2&spage=161;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=161-175.

Rall:2003:EFG

[545] Louis B. Rall. Evaluation of functions, gradients, and Jacobians. Reliable Computing = Nadezhnye vychisleniia, 9(2):177–182, April 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023098519693/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=2&spage=177;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
2&spage=177-182.

Alt:2003:GII

[546] René Alt and Jean Vignes. 10th GAMM–IMACS International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics SCAN’2002. Reliable Computing = Nadezhnye vychisleniia, 9(2):183–184, April 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1023050620601;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=9&issue=2&spage=183;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=9&issue=2&spage=
183-184.

Nataraj:2003:HOC

[547] Paluri S. V. Nataraj and Ketan Kotecha. Higher order convergence for multidimensional functions with a new Taylor–Bernstein form as inclusion function. Reliable Computing = Nadezhnye vychisleniia, 9(3):185– 203, June 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1024618415645/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=3&spage=185;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=3&spage=185-203.

Kulpa:2003:DAIb

[548] Zenon Kulpa. Diagrammatic analysis of interval linear equations. Part II: The two-dimensional case and generalization to n dimensions. Reliable Computing = Nadezhnye vychisleniia, 9(3):205–228, June 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024603432484/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=3&spage=205;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
3&spage=205-228.

Ogita:2003:CSR

[549] Takeshi Ogita, Shin’ichi Oishi, and Yasunori Ushiro. Computation of sharp rigorous componentwise error bounds for the approximate solutions of systems of linear equations. Reliable Computing = Nadezhnye vychisleniia, 9(3):229–239, June 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024655416554/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=3&spage=229;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
3&spage=229-239.

Yamamura:2003:FAS

[550] Kiyotaka Yamamura. Finding all solution sets of piecewise-trapezoidal equations described by set-valued functions. Reliable Computing = Nadezhnye vychisleniia, 9(3):241–250, June 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1024607500624/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
9&issue=3&spage=241;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=3&spage=241-250.

Puig:2003:WCS

[551] Vicenç Puig, Jordi Saludes, and Joseba Quevedo. Worst-case simulation of discrete linear time-invariant interval dynamic systems. Reliable Computing = Nadezhnye vychisleniia, 9(4):251–290, August 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024666428387/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=4&spage=251;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
4&spage=251-290.

Fang:2003:OBF

[552] Youkang Fang. Optimal bicentered form. Reliable Computing = Nadezhnye vychisleniia, 9(4):291–302, August 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1024664312457/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=4&spage=291;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
4&spage=291-302.

Sokolova:2003:ASI

[553] Svetlana P. Sokolova and Ruslan S. Ivlev. Asymptotic stability of interval Time–Delay systems. Reliable Computing = Nadezhnye vychisleniia, 9(4): 303–313, August 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1024616413366/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=4&spage=303;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=4&spage=303-313.

Kearfott:2003:DCb

[554] R. Baker Kearfott. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 9(5):315, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1023/A%3A1025127211629;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
5&spage=315.

Kearfott:2003:SIP

[555] R. Baker Kearfott. Special issue: Proceedings of the validated computing 2002 conference, Toronto, Canada, May 23–25, 2002. Reliable Computing = Nadezhnye vychisleniia, 9(5):315–316, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http:/

/link.springer.com/article/10.1023/A:1025127211629.

Achenie:2003:IGO

[556] Luke E. K. Achenie and Manish Sinha. Interval global optimization in solvent design. Reliable Computing = Nadezhnye vychisleniia, 9(5):317–338, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025158512652/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=5&spage=317;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=5&spage=317-338.

Kramer:2003:EIP

[557] Walter Krämer and Jürgen Wolff von Gudenberg. Extended interval power function. Reliable Computing = Nadezhnye vychisleniia, 9(5):339–347, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025175029490/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=5&spage=339;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=5&spage=339-347.

Kreinovich:2003:TEC

[558] Vladik Kreinovich, Luc Longpré, and James J. Buckley. Are there easy-to-check necessary and sufficient conditions for straightforward interval computations to be exact? Reliable Computing = Nadezhnye vychisleniia, 9(5):349–358, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025127113561/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=5&spage=349;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
5&spage=349-358.

Nakao:2003:SCA

[559] Mitsuhiro T. Nakao, Yoshitaka Watanabe, Nobito Yamamoto, and Takaaki Nishida. Some computer assisted proofs for solutions of the heat convection problems. Reliable Computing = Nadezhnye vychisleniia, 9(5): 359–372, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1023/
A%3A1025179130399/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=5&spage=359;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=5&spage=359-372.

Nataraj:2003:RCF

[560] Paluri S. V. Nataraj and Jayesh J. Barve. Reliable computation of frequency response plots for nonrational transfer functions to prescribed accuracy. Reliable Computing = Nadezhnye vychisleniia, 9(5):373–389, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025131214469/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=5&spage=373;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=5&spage=373-389.

Schafer:2003:AEM

[561] Uwe Schäfer. Accelerated enclosure methods for ordinary free boundary problems. Reliable Computing = Nadezhnye vychisleniia, 9(5):391–403, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025183231308/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=5&spage=391;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=5&spage=391-403.

Starks:2003:IRT

[562] Scott A. Starks and Vladik Kreinovich. Interval-related talks at the 2003 IEEE International Conference on Fuzzy Systems — St. Louis, Missouri, May 25–28, 2003. Reliable Computing = Nadezhnye vychisleniia, 9(5):405–406, October 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1025135315378;
http://www.springerlink.com/openurl.asp?genre=article&issn=1385-
3139&volume=9&issue=5&spage=405;   http://www.springerlink.com/
openurl.asp?genre=article&issn=1385-3139&volume=9&issue=5&spage=
405-406.

Berleant:2003:DHU

[563] Daniel Berleant, Mei-Peng Cheong, Chris Chu, Yong Guan, Ahmed Kamal, Gerald Shedblé, Scott Ferson, and James F. Peters. Dependable handling of uncertainty. Reliable Computing = Nadezhnye vychisleniia, 9 (6):407–418, December 2003. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025888503247/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=9&issue=6&spage=407;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=6&spage=407-418.

Horowitz:2003:CBP

[564] Joel L. Horowitz, Charles F. Manski, Maria Ponomareva, and Jörg Stoye. Computation of bounds on population parameters when the data are incomplete. Reliable Computing = Nadezhnye vychisleniia, 9(6):419–440, December 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025865520086/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=6&spage=419;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=6&spage=419-440.

Kreinovich:2003:EUB

[565] Vladik Kreinovich, Scott Ferson, and Lev Ginzburg. Exact upper bound on the mean of the product of many random variables with known expectations. Reliable Computing = Nadezhnye vychisleniia, 9(6):441–463, December 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025841220835/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=6&spage=441;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=6&spage=441-463.

Pelessoni:2003:CIP

[566] Renato Pelessoni and Paolo Vicig. Convex imprecise previsions. Reliable Computing = Nadezhnye vychisleniia, 9(6):465–485, December 2003. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025870204905/;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=9&issue=6&spage=465;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=9&issue=
6&spage=465-485.

Zaffalon:2003:TBC

[567] Marco Zaffalon and Enrico Fagiuoli. Tree–Based credal networks for classification. Reliable Computing = Nadezhnye vychisleniia, 9(6):487–509, December 2003. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025822321743/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=9&issue=6&spage=487;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=9&issue=6&spage=487-509.

Jaulin:2004:SCN

[568] Luc Jaulin, Stefan Ratschan, and Laurent Hardouin. Set computation for nonlinear control. Reliable Computing = Nadezhnye vychisleniia, 10(1):1–26, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000003994.92390.a4/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=1-26.

Nataraj:2004:GOH

[569] Paluri S. V. Nataraj and Ketan Kotecha. Global optimization with higher order inclusion function forms. Part 1: a combined Taylor– Bernstein form. Reliable Computing = Nadezhnye vychisleniia, 10 (1):27–44, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000003995.08805.2a/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=27;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=27-44.

VanEmden:2004:SDI

[570] Maarten H. Van Emden. On the significance of digits in interval notation. Reliable Computing = Nadezhnye vychisleniia, 10 (1):45–58, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000003996.60728.4e/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=45;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=45-58.

Markov:2004:SSS

[571] Svetoslav Markov and Vladik Kreinovich. Special session “Set-Valued Numerics and Reliable Computing” — Sozopol, Bulgaria, June 4–8, 2003. Reliable Computing = Nadezhnye vychisleniia, 10 (1):59–61, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000004032.19358.61/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=59;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=59-61.

Garloff:2004:IRT

[572] Jürgen Garloff and Vladik Kreinovich. Interval-related talks at the 4th International Conference on Frontiers in Global Optimization — Santorini, Greece, June 8–12, 2003. Reliable Computing = Nadezhnye vychisleniia, 10(1):63–70, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000004033.14825.f9/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=63;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=63-70.

Lodwick:2004:IRT

[573] Weldon Lodwick and Vladik Kreinovich. Interval-related talks at the 22th International Conference of the North American Fuzzy Information Processing Society NAFIPS’03 — Chicago, Illinois, USA, July 24–26, 2003. Reliable Computing = Nadezhnye vychisleniia, 10(1):71–73, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000004034.85635.3a/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=71;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=71-73.

Brattka:2004:CCA

[574] Vasco Brattka and Vladik Kreinovich. Computability and complexity in analysis (CCA). A view from Interval Computations — Cincinnati, Ohio, USA, August 28–30, 2003. Reliable Computing = Nadezhnye vychisleniia, 10(1):75–80, February 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000004035.08919.ed/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=1&spage=75;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=1&spage=75-80.

Anonymous:2004:DC

[575] Anonymous. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 10(2):81–82, April 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AREOM.0000015944.75276.df;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=10&issue=2&spage=81-82.

Nguyen:2004:DPL

[576] Hung T. Nguyen, Vladik Kreinovich, and Luc Longpré. Dirty pages of logarithm tables, lifetime of the universe, and (subjective) probabilities on finite and infinite intervals. Reliable Computing = Nadezhnye vychisleniia, 10(2):83–106, April 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000015848.19449.12/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=2&spage=83;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=10&issue=2&spage=83-106.

Tonon:2004:URS

[577] Fulvio Tonon. On the use of random set theory to bracket the results of Monte Carlo simulations. Reliable Computing = Nadezhnye vychisleniia, 10(2):107–137, April 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000015849.35108.9c/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=2&spage=107;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=2&spage=107-
137.

Berleant:2004:UPC

[578] Daniel Berleant and Jianzhong Zhang. Using Pearson correlation to improve envelopes around the distributions of functions. Reliable Computing = Nadezhnye vychisleniia, 10(2):139–161, April 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1023/B%3AREOM.0000015850.27690.
3b/;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=2&spage=139;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
2&spage=139-161.

Munoz:2004:SIG

[579] Humberto Muñoz and R. Baker Kearfott. Slope intervals, generalized gradients, semigradients, slant derivatives, and Csets. Reliable Computing = Nadezhnye vychisleniia, 10(3):163–193, June 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1023/B%3AREOM.0000032107.85627.
45/;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=3&spage=163;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
3&spage=163-193.

Huyer:2004:IAR

[580] Waltraud Huyer and Arnold Neumaier. Integral approximation of rays and verification of feasibility. Reliable Computing = Nadezhnye vychisleniia, 10(3):195–207, June 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032108.23609.bc/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=3&spage=195;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=3&spage=195-

Meer:2004:RAS

[581] Klaus Meer. On a refined analysis of some problems in interval arithmetic using real number complexity theory. Reliable Computing = Nadezhnye vychisleniia, 10(3):209–225, June 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032109.55408.1a/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=3&spage=209;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=3&spage=209-
225.

Kolev:2004:MOI

[582] Lubomir V. Kolev. A method for outer interval solution of linear parametric systems. Reliable Computing = Nadezhnye vychisleniia, 10(3):227–239, June 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032110.34735.ca/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=3&spage=227;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=3&spage=227-
239.

Ceberio:2004:FMI

[583] Martine Ceberio and Vladik Kreinovich. Fast multiplication of interval matrices (interval version of Strassen’s algorithm). Reliable Computing = Nadezhnye vychisleniia, 10(3):241–243, June 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1023/B%3AREOM.0000032111.16328.
b2/;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=3&spage=241;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
3&spage=241-243.

Nguyen:2004:IRT

[584] Hung T. Nguyen and Vladik Kreinovich. Interval-related talks at InTech’03 — Chiang Mai, Thailand, December 17–19, 2003. Reliable Computing = Nadezhnye vychisleniia, 10(3):245–246, June 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

URL http://link.springer.com/accesspage/article/10.1023/B%
3AREOM.0000032131.41547.bd;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=3&spage=245;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=10&issue=3&spage=245-246.

Lodwick:2004:DC

[585] Weldon A. Lodwick. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 10(4):247–248, August 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AREOM.0000032140.30155.7a;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=10&issue=4&spage=247;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
4&spage=247-248.

Neumaier:2004:CFS

[586] Arnold Neumaier. Clouds, fuzzy sets, and probability intervals. Reliable Computing = Nadezhnye vychisleniia, 10(4):249–272, August 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340

(electronic). URL http://link.springer.com/article/10.1023/B%
3AREOM.0000032114.08705.cd/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=4&spage=249;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=10&issue=4&spage=249-272.

Dubois:2004:PPT

[587] Didier Dubois, Laurent Foulloy, Gilles Mauris, and Henri Prade. Probability-possibility transformations, triangular fuzzy sets, and probabilistic inequalities. Reliable Computing = Nadezhnye vychisleniia, 10(4):273–297, August 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032115.22510.b5/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=4&spage=273;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=4&spage=273-
297.

Kohout:2004:CIF

[588] Ladislav J. Kohout and Eunjin Kim. Characterization of interval fuzzy logic systems of connectives by group transformations. Reliable Computing = Nadezhnye vychisleniia, 10(4):299–334, August 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1023/B%3AREOM.0000032116.98264.
16/;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=4&spage=299;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
4&spage=299-334.

Anile:2004:MUS

[589] Angelo Marcello Anile and Salvatore Spinella. Modeling uncertain sparse data with fuzzy B-splines. Reliable Computing = Nadezhnye vychisleniia, 10(5):335–355, October 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032117.04378.9a/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=5&spage=335;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=5&spage=335-
355.

Inuiguchi:2004:FLP

[590] Masahiro Inuiguchi and Tetsuzo Tanino. Fuzzy linear programming with interactive uncertain parameters. Reliable Computing = Nadezhnye vychisleniia, 10(5):357–367, October 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032118.34323.f2/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=5&spage=357;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=5&spage=357-
367.

Bondia:2004:AFI

[591] Jorge Bondia and Jesús Picó. Application of functional intervals to the response evaluation of linear time-invariant systems with fuzzy input. Reliable Computing = Nadezhnye vychisleniia, 10 (5):369–387, October 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032119.66122.e5/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=5&spage=369;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=5&spage=369-
387.

Ida:2004:SPS

[592] Masaaki Ida. Solutions for the portfolio selection problem with interval and fuzzy coefficients. Reliable Computing = Nadezhnye vychisleniia, 10(5):389–400, October 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032120.83979.d4/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=5&spage=389;   http://www.springerlink.com/openurl.

asp?genre=article&issn=1385-3139&volume=10&issue=5&spage=389-
400.

Torres:2004:EDU

[593] Roberto Torres, G. Randy Keller, Vladik Kreinovich, Luc Longpré, and Scott A. Starks. Eliminating duplicates under interval and fuzzy uncertainty: An asymptotically optimal algorithm and its geospatial applications. Reliable Computing = Nadezhnye vychisleniia, 10 (5):401–422, October 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000032121.21617.38/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=5&spage=401;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=5&spage=401-
422.

Alefeld:2004:ESL

[594] Götz Alefeld, Zhengyu Wang, and Zuhe Shen. Enclosing solutions of linear complementarity problems for H-matrices. Reliable Computing = Nadezhnye vychisleniia, 10(6):423–435, December 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.

com/article/10.1023/B%3AREOM.0000047093.79994.8f/;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=10&issue=6&spage=423;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=6&spage=423-
435.

Petkovic:2004:OLM

[595] Miodrag S. Petković and Dusan M. Milosević. Ostrowski-like method with corrections for the inclusion of polynomial zeros. Reliable Computing = Nadezhnye vychisleniia, 10(6):437–467, December 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1023/B%3AREOM.0000047094.39609.
f6/;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=6&spage=437;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
6&spage=437-467.

Kozina:2004:DOP

[596] Galina L. Kozina. Discrete optimization problems with interval data: Pareto set of solutions or set of weak solutions? Reliable Computing = Nadezhnye vychisleniia, 10(6):469–487, December 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1023/B%3AREOM.0000047095.22096.

16/;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=10&issue=6&spage=469;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=10&issue=
6&spage=469-487.

Do:2004:SAR

[597] Dang-Khoa Do. Spigot algorithm and reliable computation of natural logarithm. Reliable Computing = Nadezhnye vychisleniia, 10 (6):489–500, December 2004. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1023/B%3AREOM.0000047096.58634.fe/;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
10&issue=6&spage=489;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=10&issue=6&spage=489-
500.

Jaulin:2005:COI

[598] Luc Jaulin and Didier Henrion. Contracting optimally an interval matrix without loosing any positive semi-definite matrix is a tractable problem. Reliable Computing = Nadezhnye vychisleniia, 11(1):1–17, February 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-5939-3;   http://link.springer.com/article/10.1007/s11155-
005-5939-3/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=1&spage=1;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
11&issue=1&spage=1-17.

Utkin:2005:CSR

[599] Lev V. Utkin and Igor O. Kozine. Computing system reliability given interval-valued characteristics of the components. Reliable Computing = Nadezhnye vychisleniia, 11(1):19–34, February 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-5940-x;   http:
//link.springer.com/article/10.1007/s11155-005-5940-x/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=1&spage=19;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=1&spage=19-
34.

Rohn:2005:NFS

[600] Jiří Rohn. A normal form supplement to the Oettli–Prager theorem. Reliable Computing = Nadezhnye vychisleniia, 11(1):35–39, February 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-5941-9;   http://link.springer.com/article/10.1007/s11155-
005-5941-9/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=1&spage=35;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=1&spage=35-39. Dedicated to Prof. Dr. Gerhard

Heindl on the occasion of his retirement.

Salama:2005:ISS

[601] Abdelhay A. Salama and Emad Hamdy. Interval schemes for singularly perturbed initial value problems. Reliable Computing = Nadezhnye vychisleniia, 11(1):41–58, February 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-5942-8;   http:
//link.springer.com/article/10.1007/s11155-005-5942-8/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=1&spage=41;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=1&spage=41-
58.

Kreinovich:2005:ODU

[602] Vladik Kreinovich, Luc Longpré, Praveen Patangay, Scott Ferson, and Lev Ginzburg. Outlier detection under interval uncertainty: Algorithmic solvability and computational complexity. Reliable Computing = Nadezhnye vychisleniia, 11(1):59–76, February 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-5943-7;   http:
//link.springer.com/article/10.1007/s11155-005-5943-7/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=1&spage=59;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=1&spage=59-
76.

Schnurr:2005:PSS

[603] Marco Schnurr. On the proofs of some statements concerning the theorems of Kantorovich, Moore, and Miranda. Reliable Computing = Nadezhnye vychisleniia, 11(1):77–85, February 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-5944-6;   http:
//link.springer.com/article/10.1007/s11155-005-5944-6/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=1&spage=77;   http://www.springerlink.com/openurl.

asp?genre=article&issn=1385-3139&volume=11&issue=1&spage=77-
85.

Arndt:2005:SIS

[604] Hans-Robert Arndt and Günter Mayer. On the solutions of the interval system [x] = [A][x]+[b]. Reliable Computing = Nadezhnye vychisleniia, 11 (2):87–103, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-005-3031-7;   http://link.springer.com/article/10.1007/
s11155-005-3031-7/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=11&issue=2&spage=87;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=2&spage=87-103. Dedicated to Professor Dr. J.

Herzberger on the occasion of his 65-th birthday.

Lyashko:2005:OSI

[605] Marina A. Lyashko. The optimal solution of an interval system of linear algebraic equations. Reliable Computing = Nadezhnye vychisleniia, 11(2): 105–127, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-005-3032-6;   http://link.springer.com/article/10.1007/
s11155-005-3032-6/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=11&issue=2&spage=105;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=2&spage=105-127.

Rohn:2005:LIE

[606] Jiří Rohn. Linear interval equations: Midpoint preconditioning may produce a 100% overestimation for arbitrarily narrow data even in case n = 4. Reliable Computing = Nadezhnye vychisleniia, 11(2):129–135, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-3033-5;   http://link.springer.com/article/10.1007/s11155-
005-3033-5/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=2&spage=129;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=2&spage=129-135.

Schon:2005:UZO

[607] Steffen Schön and Hansjörg Kutterer. Using zonotopes for overestimationfree interval least-squares — some geodetic applications. Reliable Computing = Nadezhnye vychisleniia, 11(2):137–155, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-3034-4;   http:
//link.springer.com/article/10.1007/s11155-005-3034-4/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=2&spage=137;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=2&spage=137-
155.

Nakao:2005:SGI

[608] Mitsuhiro T. Nakao and Kaori Nagatou. SCAN’2004 11th GAMM– IMACS International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics, Fukuoka, Japan, October 4–8, 2004. Reliable Computing = Nadezhnye vychisleniia, 11(2):157–159, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-3037-1;   http://link.springer.com/article/10.1007/s11155-
005-3037-1/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=2&spage=157;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=2&spage=157-159.

Anonymous:2005:SRM

[609] Anonymous. Second R. E. Moore Prize awarded at SCAN’2004. Reliable Computing = Nadezhnye vychisleniia, 11(2):161–162, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s11155-005-3036-2;   http://link.springer.com/article/10.
1007/s11155-005-3036-2;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=2&spage=161-
162.

Kearfott:2005:EOI

[610] R. Baker Kearfott. Errata and opinion to: “An interval entropy penalty method for nonlinear global optimization” [Reliab. Comput. 4(1) (1998), 15–25; MR1617525] by Zhengyu Huang. Reliable Computing = Nadezhnye vychisleniia, 11(2):163–164, April 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s11155-005-3035-3;   http://link.
springer.com/article/10.1007/s11155-005-3035-3;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
11&issue=2&spage=163;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=2&spage=163-

164. See [340].

Alefeld:2005:ESS

[611] Götz Alefeld and Günter Mayer. Enclosing solutions of singular interval systems iteratively. Reliable Computing = Nadezhnye vychisleniia, 11(3): 165–190, June 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-005-3614-3;   http://link.springer.com/article/10.1007/
s11155-005-3614-3/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=11&issue=3&spage=165;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=3&spage=165-190.

Ogita:2005:FII

[612] Takeshi Ogita and Shin’ichi Oishi. Fast inclusion of interval matrix multiplication. Reliable Computing = Nadezhnye vychisleniia, 11(3):191–205, June 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-3615-2;   http://link.springer.com/article/10.1007/s11155-
005-3615-2/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=3&spage=191;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=3&spage=191-205.

Ferson:2005:EBF

[613] Scott Ferson, Lev Ginzburg, Vladik Kreinovich, Luc Longpré, and Monica Aviles. Exact bounds on finite populations of interval data. Reliable Computing = Nadezhnye vychisleniia, 11(3):207–233, June 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-3616-1;   http:
//link.springer.com/article/10.1007/s11155-005-3616-1/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=3&spage=207;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=3&spage=207-
233.

Wallner:2005:TGC

[614] Johannes Wallner, Hans-Peter Schröcker, and Shi-Min Hu. Tolerances in geometric constraint problems. Reliable Computing = Nadezhnye vychisleniia, 11(3):235–251, June 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1007/s11155-005-3617-0;   http://link.springer.com/
article/10.1007/s11155-005-3617-0/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=11&issue=

3&spage=235;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=3&spage=235-251.

Toth:2005:EIC

[615] Boglárka Tóth and Tibor Csendes. Empirical investigation of the convergence speed of inclusion functions in a global optimization context. Reliable Computing = Nadezhnye vychisleniia, 11(4):253–273, August 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-6890-z;   http://link.springer.com/article/10.1007/s11155-
005-6890-z/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=4&spage=253;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=4&spage=253-273.

Revol:2005:MAP

[616] Nathalie Revol and Fabrice Rouillier. Motivations for an arbitrary precision interval arithmetic and the MPFI library. Reliable Computing = Nadezhnye vychisleniia, 11(4):275–290, August 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-6891-y;   http:
//link.springer.com/article/10.1007/s11155-005-6891-y/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=4&spage=275;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=4&spage=275-
290.

Tupelly:2005:CIT

[617] Kavitha Tupelly, Vladik Kreinovich, and Karen Villaverde. Checking if there exists a monotonic function that is consistent with the measurements: An efficient algorithm. Reliable Computing = Nadezhnye vychisleniia, 11(4):291–312, August 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-6892-x;   http:
//link.springer.com/article/10.1007/s11155-005-6892-x/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=4&spage=291;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=4&spage=291-
312.

Backeljauw:2005:CCC

[618] Franky Backeljauw and Annie Cuyt. A constructive criticism of the C/C++ proposal for complex arithmetic. Reliable Computing = Nadezhnye vychisleniia, 11(4):313–319, August 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-6893-9;   http:
//link.springer.com/article/10.1007/s11155-005-6893-9/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=4&spage=313;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=4&spage=313-
319.

Shary:2005:P

[619] Sergey P. Shary. Preface. Reliable Computing = Nadezhnye vychisleniia, 11(5):321–322, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s11155-005-0043-2;   http://link.
springer.com/article/10.1007/s11155-005-0043-2;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
11&issue=5&spage=321;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=5&spage=321-
322.

Akhmerov:2005:IAG

[620] Ramil R. Akhmerov. Interval-affine Gaussian algorithm for constrained systems. Reliable Computing = Nadezhnye vychisleniia, 11(5):323–341, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-005-0040-5;   http://link.springer.com/article/10.1007/
s11155-005-0040-5/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=11&issue=5&spage=323;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=323-341.

Dolgov:2005:DIG

[621] Yuri G. Dolgov. Developing interval global optimization algorithms on the basis of branch-and-bound and constraint propagation methods. Reliable Computing = Nadezhnye vychisleniia, 11(5):343–358, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-0041-4;   http://link.springer.com/article/10.1007/s11155-
005-0041-4/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=5&spage=343;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=343-358.

Ershov:2005:IML

[622] Alexei G. Ershov and Tamara P. Kashevarova. Interval mathematical library based on Chebyshev and Taylor series expansion. Reliable Computing = Nadezhnye vychisleniia, 11(5):359–367, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-0042-3;   http:
//link.springer.com/article/10.1007/s11155-005-0042-3/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=359;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=5&spage=359-
367.

Herrero:2005:QSI

[623] Pau Herrero, Miguel A. Sainz, Josep Veh, and Luc Jaulin. Quantified set inversion algorithm with applications to control. Reliable Computing = Nadezhnye vychisleniia, 11(5):369–382, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-0044-1;   http:
//link.springer.com/article/10.1007/s11155-005-0044-1/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=369;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=5&spage=369-
382.

Kearfott:2005:VCS

[624] R. Baker Kearfott. Validated constraint solving — practicalities, pitfalls, and new developments. Reliable Computing = Nadezhnye vychisleniia, 11 (5):383–391, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1007/s11155-005-0045-0;   http://link.springer.com/article/10.
1007/s11155-005-0045-0/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=5&spage=383;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=11&issue=5&spage=383-391.

Kubica:2005:IGO

[625] Bartłomiej Jacek Kubica and Krzysztof Malinowski. An interval global optimization algorithm combining symbolic rewriting and componentwise Newton method applied to control a class of queueing systems. Reliable Computing = Nadezhnye vychisleniia, 11(5):393–411, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-0047-y;   http://link.springer.com/article/10.1007/s11155-

005-0047-y/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=5&spage=393;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=393-411.

Pushkov:2005:BRM

[626] Sergey G. Pushkov and Svetlana Yu. Kalinkina. Boundary realizations method for interval linear dynamic systems. Reliable Computing = Nadezhnye vychisleniia, 11(5):413–423, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-0048-x;   http:
//link.springer.com/article/10.1007/s11155-005-0048-x/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=413;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=5&spage=413-
423.

Sharaya:2005:UTS

[627] Irene A. Sharaya. On unbounded tolerable solution sets. Reliable Computing = Nadezhnye vychisleniia, 11(5):425–432, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-0049-9;   http:
//link.springer.com/article/10.1007/s11155-005-0049-9/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=425;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=5&spage=425-
432.

Zhilin:2005:FED

[628] Sergei I. Zhilin. On fitting empirical data under interval error. Reliable Computing = Nadezhnye vychisleniia, 11(5):433–442, October 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-0050-3;   http://link.springer.com/article/10.1007/s11155-
005-0050-3/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=5&spage=433;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=5&spage=433-442.

Goldsztejn:2005:RPP

[629] Alexandre Goldsztejn. A right-preconditioning process for the formalalgebraic approach to inner and outer estimation of AE-solution sets. Reliable Computing = Nadezhnye vychisleniia, 11(6):443–478, December 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-0404-x;   http://link.springer.com/article/10.1007/s11155-
005-0404-x/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=6&spage=443;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=6&spage=443-478.

Kreinovich:2005:OFC

[630] Vladik Kreinovich. Optimal finite characterization of linear problems with inexact data. Reliable Computing = Nadezhnye vychisleniia, 11(6):479– 489, December 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-005-0406-8;   http://link.springer.com/article/10.1007/
s11155-005-0406-8/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=11&issue=6&spage=479;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=6&spage=479-489.

Rohn:2005:HSS

[631] Jiří Rohn. How strong is strong regularity? Reliable Computing = Nadezhnye vychisleniia, 11(6):491–493, December 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-0407-7;   http:
//link.springer.com/article/10.1007/s11155-005-0407-7/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=6&spage=491;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=6&spage=491-
493.

Hansen:2005:TRI

[632] Eldon R. Hansen. A theorem on regularity of interval matrices. Reliable Computing = Nadezhnye vychisleniia, 11(6):495–497, December 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
005-0405-9;   http://link.springer.com/article/10.1007/s11155-
005-0405-9/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=11&issue=6&spage=495;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=6&spage=495-497.

Ceberio:2005:RCT

[633] Martine Ceberio, Vladik Kreinovich, and Michel Rueher. Reliable computations and their applications (RCA) track: a technical track at the 20th ACM Symposium on Applied Computing SAC’2005. Reliable Computing = Nadezhnye vychisleniia, 11(6):499–503, December 2005. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-005-0408-6;   http:
//link.springer.com/article/10.1007/s11155-005-0408-6/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=11&issue=6&spage=499;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=11&issue=6&spage=499-
503.

Candau:2006:CIA

[634] Yves Candau, Tarek Raissi, Nacim Ramdani, and Laurent Ibos. Complex interval arithmetic using polar form. Reliable Computing = Nadezhnye vychisleniia, 12(1):1–20, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-2966-7;   http:
//link.springer.com/article/10.1007/s11155-006-2966-7/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=1;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=1&spage=1-20.

Hansen:2006:SIC

[635] Eldon R. Hansen. Sharpening interval computations. Reliable Computing = Nadezhnye vychisleniia, 12(1):21–34, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-2967-6;   http:
//link.springer.com/article/10.1007/s11155-006-2967-6/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=21;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=1&spage=21-
34.

Djanybekov:2006:IHM

[636] Bakyt S. Djanybekov. Interval Householder method for complex linear systems. Reliable Computing = Nadezhnye vychisleniia, 12(1):35–43, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-2968-5;   http://link.springer.com/article/10.1007/s11155-
006-2968-5/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=1&spage=35;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=35-43.

Patre:2006:RCT

[637] Balasaheb M. Patre and Bijnan Bandyopadhyay. Robust control for two-time-scale discrete interval systems. Reliable Computing = Nadezhnye vychisleniia, 12(1):45–58, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-2971-x;   http:
//link.springer.com/article/10.1007/s11155-006-2971-x/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=45;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=1&spage=45-
58.

Xiang:2006:FAC

[638] Gang Xiang. Fast algorithm for computing the upper endpoint of sample variance for interval data: Case of sufficiently accurate measurements. Reliable Computing = Nadezhnye vychisleniia, 12(1):59–64, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-2965-8;   http://link.springer.com/article/10.1007/s11155-
006-2965-8/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=1&spage=59;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=59-64.

Ross:2006:ANL

[639] Timothy J. Ross and Vladik Kreinovich. Los Alamos National Laboratory Uncertainty Workshop: An interval perspective. Reliable Computing = Nadezhnye vychisleniia, 12(1):65–71, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-2969-4;   http:
//link.springer.com/article/10.1007/s11155-006-2969-4/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=65;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=1&spage=65-
71.

Kreinovich:2006:TRW

[640] Vladik Kreinovich, François Modave, Scott Starks, and Gang Xiang. Towards real world applications: Interval-related talks at NAFIPS’05. Reliable Computing = Nadezhnye vychisleniia, 12(1):73–77, February 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-2970-y;   http://link.springer.com/article/10.1007/s11155-

006-2970-y/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=1&spage=73;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=1&spage=73-77.

Mayer:2006:CFI

[641] Günter Mayer. A contribution to the feasibility of the interval Gaussian algorithm. Reliable Computing = Nadezhnye vychisleniia, 12(2):79–98, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-4876-0;   http://link.springer.com/article/10.1007/s11155-
006-4876-0/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=2&spage=79;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=79-98.

Rohn:2006:RIM

[642] Jiří Rohn. Regularity of interval matrices and theorems of the alternatives. Reliable Computing = Nadezhnye vychisleniia, 12(2):99–105, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-4877-z;   http://link.springer.com/article/10.1007/s11155-
006-4877-z/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=2&spage=99;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=99-105.

Skalna:2006:MOI

[643] Iwona Skalna. A method for outer interval solution of systems of linear equations depending linearly on interval parameters. Reliable Computing = Nadezhnye vychisleniia, 12(2):107–120, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-4878-y;   http:
//link.springer.com/article/10.1007/s11155-006-4878-y/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=107;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=2&spage=107-
120.

Kolev:2006:OIS

[644] Lubomir V. Kolev. Outer interval solution of the eigenvalue problem under general form parametric dependencies. Reliable Computing = Nadezhnye vychisleniia, 12(2):121–140, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-4875-1;   http:
//link.springer.com/article/10.1007/s11155-006-4875-1/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=121;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=2&spage=121-
140.

Beelitz:2006:ETS

[645] Thomas Beelitz, Bruno Lang, and Christian H. Bischof. Efficient task scheduling in the parallel result-verifying solution of nonlinear systems. Reliable Computing = Nadezhnye vychisleniia, 12(2):141–151, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-4872-4;   http://link.springer.com/article/10.1007/s11155-
006-4872-4/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=2&spage=141;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=141-151.

Berleant:2006:DIP

[646] Daniel Berleant, Fabio G. Cozman, Olga Kosheleva, and Vladik Kreinovich. Dealing with imprecise probabilities: Interval-related talks at ISIPTA’05. Reliable Computing = Nadezhnye vychisleniia, 12(2):153–165, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-4874-2;   http://link.springer.com/article/10.1007/s11155-
006-4874-2/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=2&spage=153;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=153-165.

Corliss:2006:SSW

[647] George F. Corliss and Kaj Madsen. Second Scandinavian Workshop on Interval Methods and Their Applications. Reliable Computing = Nadezhnye vychisleniia, 12(2):167–169, April 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-4873-3;   http:
//link.springer.com/article/10.1007/s11155-006-4873-3/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=2&spage=167;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=2&spage=167-

Messine:2006:GRQ

[648] Frédéric Messine and Ahmed Touhami. A general reliable quadratic form: An extension of affine arithmetic. Reliable Computing = Nadezhnye vychisleniia, 12(3):171–192, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1007/s11155-006-7217-4;   http://link.springer.com/
article/10.1007/s11155-006-7217-4/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=12&issue=
3&spage=171;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=3&spage=171-192.

Kolev:2006:IDM

[649] Lubomir V. Kolev. Improvement of a direct method for outer solution of linear parametric systems. Reliable Computing = Nadezhnye vychisleniia, 12(3):193–202, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1007/s11155-006-7218-3;   http://link.springer.com/article/10.
1007/s11155-006-7218-3/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=3&spage=193;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=12&issue=3&spage=193-202.

Schrocker:2006:GCD

[650] Hans-Peter Schröcker and Johannes Wallner. Geometric constructions with discretized random variables. Reliable Computing = Nadezhnye vychisleniia, 12(3):203–223, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1007/s11155-006-7219-2;   http://link.springer.com/
article/10.1007/s11155-006-7219-2/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=12&issue=
3&spage=203;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=3&spage=203-223.

Kirchner:2006:HSI

[651] Reinhard Kirchner and Ulrich W. Kulisch. Hardware support for interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 12(3):225–237, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-7220-9;   http://link.springer.com/article/10.1007/s11155-
006-7220-9/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=3&spage=225;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=3&spage=225-237.

Hansen:2006:SOS

[652] Eldon R. Hansen and G. William Walster. Solving overdetermined systems of interval linear equations. Reliable Computing = Nadezhnye vychisleniia, 12(3):239–243, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print),

1573-1340 (electronic). URL http://link.springer.com/article/10.
1007/s11155-006-7221-8;   http://link.springer.com/article/10.
1007/s11155-006-7221-8/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=3&spage=239;
 http://www.springerlink.com/openurl.asp?genre=article&issn=
1385-3139&volume=12&issue=3&spage=239-243.

Rohn:2006:LE

[653] Jiří Rohn. Letter to the Editor. Reliable Computing = Nadezhnye vychisleniia, 12(3):245–246, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s11155-006-7222-7;   http://link.
springer.com/article/10.1007/s11155-006-7222-7;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
12&issue=3&spage=245;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=3&spage=245-
246.

Castillo:2006:IRT

[654] Oscar Castillo, Patricia Melin, and Vladik Kreinovich. Interval-related talks at the International Conference on Fuzzy Systems, Neural Networks, and Genetic Algorithms, FNG’05, Tijuana, Baja California, Mexico, October 13–14, 2005. Reliable Computing = Nadezhnye vychisleniia, 12(3): 247–251, June 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-7223-6;   http://link.springer.com/article/10.1007/
s11155-006-7223-6/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=12&issue=3&spage=247;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=3&spage=247-251.

Hansen:2006:MIN

[655] Eldon R. Hansen. A multidimensional interval Newton method. Reliable Computing = Nadezhnye vychisleniia, 12(4):253–272, August 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9000-y;   http://link.springer.com/article/10.1007/s11155-
006-9000-y/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=4&spage=253;   http:/

/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=4&spage=253-272.

Dantsin:2006:PVU

[656] Evgeny Dantsin, Vladik Kreinovich, Alexander Wolpert, and Gang Xiang. Population variance under interval uncertainty: a new algorithm. Reliable Computing = Nadezhnye vychisleniia, 12(4):273–280, August 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9001-x;   http://link.springer.com/article/10.1007/s11155-
006-9001-x/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=4&spage=273;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=4&spage=273-280.

Kolev:2006:NFM

[657] Lubomir V. Kolev. New formulae for multiplication of intervals. Reliable Computing = Nadezhnye vychisleniia, 12(4):281–292, August 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9002-9;   http://link.springer.com/article/10.1007/s11155-
006-9002-9/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=4&spage=281;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=4&spage=281-292.

Dyllong:2006:MGO

[658] Eva Dyllong, Wolfram Luther, and Holger Traczinski. Modelling geometric objects and tolerances with intervals: Data exchange with ISO standard STEP. Reliable Computing = Nadezhnye vychisleniia, 12(4): 293–302, August 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9003-8;   http://link.springer.com/article/10.1007/
s11155-006-9003-8/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=12&issue=4&spage=293;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=4&spage=293-302.

Keil:2006:CER

[659] Christian Keil and Christian Jansson. Computational experience with rigorous error bounds for the Netlib Linear Programming Library. Reliable Computing = Nadezhnye vychisleniia, 12(4):303–321, August 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9004-7;   http://link.springer.com/article/10.1007/s11155-
006-9004-7/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=4&spage=303;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=4&spage=303-321.

Shou:2006:RPP

[660] Huahao Shou, Jie Shen, and David Yoon. Robust plotting of polar algebraic curves, space algebraic curves, and offsets of planar algebraic curves. Reliable Computing = Nadezhnye vychisleniia, 12(4):323–335, August 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9005-6;   http://link.springer.com/article/10.1007/s11155-
006-9005-6/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=4&spage=323;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=4&spage=323-335.

Anguelov:2006:SHC

[661] Roumen Anguelov, Svetoslav Markov, and Blagovest Sendov. The set of Hausdorff continuous functions — the largest linear space of interval functions. Reliable Computing = Nadezhnye vychisleniia, 12(5):337–363, October 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9006-5;   http://link.springer.com/article/10.1007/s11155-
006-9006-5/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=5&spage=337;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=5&spage=337-363.

Kreinovich:2006:TOU

[662] Vladik Kreinovich and Siegfried Rump. Towards optimal use of multiprecision arithmetic: a remark. Reliable Computing = Nadezhnye vychisleniia, 12(5):365–369, October 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1007/s11155-006-9007-4;   http://link.springer.com/
article/10.1007/s11155-006-9007-4/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=12&issue=
5&spage=365;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=5&spage=365-369.

Lordelo:2006:ADR

[663] Alfredo D. S. Lordelo, Edvaldo A. Juzzo, and Paulo A. V. Ferreira. Analysis and design of robust controllers using the interval Diophantine equation. Reliable Computing = Nadezhnye vychisleniia, 12(5):371–388, October 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9008-3;   http://link.springer.com/article/10.1007/s11155-
006-9008-3/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=5&spage=371;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=5&spage=371-388.

Tucker:2006:PRB

[664] Warwick Tucker and Vincent Moulton. Parameter reconstruction for biochemical networks using interval analysis. Reliable Computing = Nadezhnye vychisleniia, 12(5):389–402, October 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9009-2;   http:
//link.springer.com/article/10.1007/s11155-006-9009-2/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=5&spage=389;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=5&spage=389-
402.

Anonymous:2006:DC

[665] Anonymous. Dear colleagues. Reliable Computing = Nadezhnye vychisleniia, 12(6):403–404, December 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s11155-006-9010-9;   http://link.
springer.com/article/10.1007/s11155-006-9010-9;   http://www.
springerlink.com/openurl.asp?genre=article&issn=1385-3139&volume=
12&issue=6&spage=403-404.

Moore:2006:IRR

[666] Ramon E. Moore. Introductory remarks on reliable engineering computing. Reliable Computing = Nadezhnye vychisleniia, 12(6):405–408, December 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9011-8;   http://link.springer.com/article/10.1007/s11155-
006-9011-8/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=12&issue=6&spage=405;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=6&spage=405-408.

Johnson:2006:HUD

[667] David B. Johnson and I. David L. Bogle. Handling uncertainty in the development and design of chemical processes. Reliable Computing = Nadezhnye vychisleniia, 12(6):409–426, December 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9012-7;   http:
//link.springer.com/article/10.1007/s11155-006-9012-7/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=6&spage=409;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=6&spage=409-
426.

Lin:2006:RMO

[668] Youdong Lin, C. Ryan Gwaltney, and Mark A. Stadtherr. Reliable modeling and optimization for chemical engineering applications: Interval analysis approach. Reliable Computing = Nadezhnye vychisleniia, 12(6):427– 450, December 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9013-6;   http://link.springer.com/article/10.1007/
s11155-006-9013-6/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=12&issue=6&spage=427;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=6&spage=427-450.

Pereira:2006:UTB

[669] Sebastião C. Pereira, Ulisses T. Mello, Nelson F. F. Ebecken, and Rafi L. Muhanna. Uncertainty in thermal basin modeling: An interval finite element approach. Reliable Computing = Nadezhnye vychisleniia, 12(6):451– 470, December 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9014-5;   http://link.springer.com/article/10.1007/
s11155-006-9014-5/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=12&issue=6&spage=451;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=6&spage=451-470.

Kreinovich:2006:TCP

[670] Vladik Kreinovich, Gang Xiang, Scott A. Starks, Luc Longpré, Martine Ceberio, Roberto Araiza, Jan Beck, Raj Kandathi, Asis Nayak, Roberto Torres, and Janos G. Hajagos. Towards combining probabilistic and interval uncertainty in engineering calculations: Algorithms for computing statistics under interval uncertainty, and their computational complexity. Reliable Computing = Nadezhnye vychisleniia, 12(6):471–501, December 2006. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9015-4;   http:
//link.springer.com/article/10.1007/s11155-006-9015-4/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=12&issue=6&spage=471;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=12&issue=6&spage=471-
501.

Paluri:2007:ERC

[671] Nataraj S. V. Paluri and Shanta Sondur. Experiments with range computations using extrapolation. Reliable Computing = Nadezhnye vychisleniia, 13(1):1–23, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1007/s11155-006-9022-5;   http://link.springer.com/
article/10.1007/s11155-006-9022-5/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=13&issue=
1&spage=1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=1385-3139&volume=13&issue=1&spage=1-23.

Kreinovich:2007:MCT

[672] Vladik Kreinovich, Jan Beck, Carlos Ferregut, Araceli Sanchez, G. Randy Keller, Matthew Averill, and Scott A. Starks. Monte-Carlo-type techniques for processing interval uncertainty, and their potential engineering applications. Reliable Computing = Nadezhnye vychisleniia, 13(1): 25–69, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9021-6;   http://link.springer.com/article/10.1007/
s11155-006-9021-6/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=1&spage=25;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=1&spage=25-69.

Hajagos:2007:IMC

[673] Janos G. Hajagos. Interval Monte Carlo as an alternative to secondorder sampling for estimating ecological risk. Reliable Computing = Nadezhnye vychisleniia, 13(1):71–81, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9019-0;   http:
//link.springer.com/article/10.1007/s11155-006-9019-0/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=1&spage=71;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=1&spage=71-

Smith:2007:RAE

[674] Spencer Smith, Lei Lai, and Ridha Khedri. Requirements analysis for engineering computation: a systematic approach for improving reliability. Reliable Computing = Nadezhnye vychisleniia, 13(1):83–107, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9020-7;   http://link.springer.com/article/10.1007/s11155-
006-9020-7/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=1&spage=83;   http://
www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=1&spage=83-107.

Lodwick:2007:IRT

[675] Weldon Lodwick and Vladik Kreinovich. Interval-related talks at the North American Fuzzy Information Processing Society Annual Conference NAFIPS’06. Reliable Computing = Nadezhnye vychisleniia, 13(1):109– 111, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9017-2;   http://link.springer.com/article/10.1007/
s11155-006-9017-2/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=1&spage=109;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=1&spage=109-111.

Starks:2007:IRT

[676] Scott A. Starks and Vladik Kreinovich. Interval-related talks at the Second International Conference on Fuzzy Sets and Soft Computing in Economics and Finance. Reliable Computing = Nadezhnye vychisleniia, 13(1):113– 117, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9018-1;   http://link.springer.com/article/10.1007/
s11155-006-9018-1/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=1&spage=113;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=1&spage=113-117.

Kreinovich:2007:ITI

[677] Vladik Kreinovich and Scott A. Starks. Interval talks at the International Conference on Information Processing and Management of Uncertainty in Knowledge–Based Systems IPMU’2006. Reliable Computing = Nadezhnye vychisleniia, 13(1):119–124, February 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9016-3;   http:
//link.springer.com/article/10.1007/s11155-006-9016-3/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=1&spage=119;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=1&spage=119-
124.

Corliss:2007:FRA

[678] George Corliss, Christopher Foley, and R. Baker Kearfott. Formulation for reliable analysis of structural frames. Reliable Computing = Nadezhnye vychisleniia, 13(2):125–147, April 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9027-0;   http:
//link.springer.com/article/10.1007/s11155-006-9027-0/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=2&spage=125;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=2&spage=125-
147.

Neumaier:2007:LSL

[679] Arnold Neumaier and Andrzej Pownuk. Linear systems with large uncertainties, with applications to truss structures. Reliable Computing = Nadezhnye vychisleniia, 13(2):149–172, April 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9026-1;   http:
//link.springer.com/article/10.1007/s11155-006-9026-1/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=2&spage=149;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=2&spage=149-
172.

Muhanna:2007:IFE

[680] Rafi L. Muhanna, Hao Zhang, and Robert L. Mullen. Interval finite elements as a basis for generalized models of uncertainty in engineering mechanics. Reliable Computing = Nadezhnye vychisleniia, 13(2):173–194, April 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9024-3;   http://link.springer.com/article/10.1007/s11155-
006-9024-3/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=2&spage=173;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=2&spage=173-194.

Tonon:2007:SAC

[681] Fulvio Tonon. A search algorithm for calculating validated reliability bounds. Reliable Computing = Nadezhnye vychisleniia, 13(2):195–209, April 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9025-2;   http://link.springer.com/article/10.1007/s11155-
006-9025-2/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=2&spage=195;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=2&spage=195-209.

Wang:2007:SIC

[682] Yan Wang and Bartholomew O. Nnaji. Solving interval constraints by linearization in computer-aided design. Reliable Computing = Nadezhnye vychisleniia, 13(2):211–244, April 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://link.springer.com/

article/10.1007/s11155-006-9023-4;   http://link.springer.com/
article/10.1007/s11155-006-9023-4/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1385-3139&volume=13&issue=
2&spage=211;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=2&spage=211-244.

Arndt:2007:ISP

[683] Hans-Robert Arndt. On interval systems [x] = [A][x] + [b] and the powers of interval matrices in complex interval arithmetics. Reliable Computing = Nadezhnye vychisleniia, 13(3):245–259, June 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9032-3;   http:
//link.springer.com/article/10.1007/s11155-006-9032-3/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=3&spage=245;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=3&spage=245-
259.

Berleant:2007:UIL

[684] Daniel J. Berleant, Olga Kosheleva, Vladik Kreinovich, and Hung T. Nguyen. Unimodality, independence lead to NP-hardness of interval probability problems. Reliable Computing = Nadezhnye vychisleniia, 13(3): 261–282, June 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-006-9031-4;   http://link.springer.com/article/10.1007/
s11155-006-9031-4/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=3&spage=261;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=3&spage=261-282.

Ivlev:2007:ESI

[685] Ruslan S. Ivlev and Svetlana P. Sokolova. Exponential stability of interval dynamical systems with quadratic nonlinearity. Reliable Computing = Nadezhnye vychisleniia, 13(3):283–291, June 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-006-9029-y;   http:
//link.springer.com/article/10.1007/s11155-006-9029-y/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=3&spage=283;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=3&spage=283-
291.

Stoye:2007:BGL

[686] Jörg Stoye. Bounds on generalized linear predictors with incomplete outcome data. Reliable Computing = Nadezhnye vychisleniia, 13(3):293–302, June 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9030-5;   http://link.springer.com/article/10.1007/s11155-
006-9030-5/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=3&spage=293;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=3&spage=293-302.

Rall:2007:EAD

[687] Louis B. Rall. Early automatic differentiation: The Ch’in–Horner algorithm. Reliable Computing = Nadezhnye vychisleniia, 13(3):303–308, June 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
006-9028-z;   http://link.springer.com/article/10.1007/s11155-
006-9028-z/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=3&spage=303;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=3&spage=303-308.

Faudot:2007:NRA

[688] Dominique Faudot and Dominique Michelucci. A new robust algorithm to trace curves. Reliable Computing = Nadezhnye vychisleniia, 13(4): 309–324, August 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-007-9036-7;   http://link.springer.com/article/10.1007/

s11155-007-9036-7/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=4&spage=309;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=4&spage=309-324.

Chabert:2007:EHB

[689] Gilles Chabert and Alexandre Goldsztejn. Extension of the Hansen– Bliek method to right-quantified linear systems. Reliable Computing = Nadezhnye vychisleniia, 13(4):325–349, August 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-007-9037-6;   http:
//link.springer.com/article/10.1007/s11155-007-9037-6/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=4&spage=325;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=4&spage=325-
349.

Garloff:2007:GPS

[690] Jürgen Garloff, Ismail Idriss, and Andrew P. Smith. Guaranteed parameter set estimation for exponential sums: The three-terms case. Reliable Computing = Nadezhnye vychisleniia, 13(4):351–359, August 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
007-9034-9;   http://link.springer.com/article/10.1007/s11155-
007-9034-9/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=4&spage=351;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=4&spage=351-359.

Hladik:2007:SSC

[691] Milan Hladík. Solution set characterization of linear interval systems with a specific dependence structure. Reliable Computing = Nadezhnye vychisleniia, 13(4):361–374, August 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-007-9033-x;   http:
//link.springer.com/article/10.1007/s11155-007-9033-x/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=4&spage=361;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=4&spage=361-
374.

Kreinovich:2007:IRT

[692] V. Kreinovich and R. Muhanna. Interval-related talks at the International Conference on Finite Element Methods in Engineering and Science, El Paso, Texas, December 11–15, 2006. Reliable Computing = Nadezhnye vychisleniia, 13(4):375–379, August 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-007-9035-8;   http:
//link.springer.com/article/10.1007/s11155-007-9035-8/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=4&spage=375;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=4&spage=375-
379.

Delanoue:2007:GHT

[693] Nicolas Delanoue, Luc Jaulin, and Bertrand Cottenceau. Guaranteeing the homotopy type of a set defined by non-linear inequalities. Reliable Computing = Nadezhnye vychisleniia, 13(5):381–398, October 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-007-9043-8;   http:
//link.springer.com/article/10.1007/s11155-007-9043-8/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=5&spage=381;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=5&spage=381-
398.

Kolev:2007:OMI

[694] Lubomir V. Kolev. Optimal multiplication of G-intervals. Reliable Computing = Nadezhnye vychisleniia, 13(5):399–408, October 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-007-9041-x;   http:
//link.springer.com/article/10.1007/s11155-007-9041-x/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=5&spage=399;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=5&spage=399-
408.

Lagrange:2007:SCI

[695] Sébastien Lagrange, Nicolas Delanoue, and Luc Jaulin. On sufficient conditions of the injectivity: Development of a numerical test algorithm via interval analysis. Reliable Computing = Nadezhnye vychisleniia, 13(5): 409–421, October 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-007-9042-9;   http://link.springer.com/article/10.1007/
s11155-007-9042-9/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=5&spage=409;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=5&spage=409-421.

Hu:2007:AIM

[696] Chenyi Hu and Ling T. He. An application of interval methods to stock market forecasting. Reliable Computing = Nadezhnye vychisleniia, 13(5): 423–434, October 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-007-9039-4;   http://link.springer.com/article/10.1007/
s11155-007-9039-4/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=5&spage=423;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=5&spage=423-434.

Castillo:2007:IRTa

[697] Oscar Castillo and Vladik Kreinovich. Interval-related talks at the 2007 IEEE Symposium Series on Computational Intelligence, Honolulu, Hawaii, April 1–5, 2007. Reliable Computing = Nadezhnye vychisleniia, 13(5):435– 440, October 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-007-9040-y;   http://link.springer.com/article/10.1007/
s11155-007-9040-y/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=5&spage=435;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=5&spage=435-440.

Castillo:2007:IRTb

[698] Oscar Castillo and Vladik Kreinovich. Interval-related talks at the North American Fuzzy Information Processing Society Annual Conference NAFIPS’07, San Diego, California, June 24–27, 2007. Reliable Computing = Nadezhnye vychisleniia, 13(5):441–443, October 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http://

link.springer.com/article/10.1007/s11155-007-9038-5;   http:
//link.springer.com/article/10.1007/s11155-007-9038-5/;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=5&spage=441;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1385-3139&volume=13&issue=5&spage=441-
443.

Kolev:2007:DPD

[699] Lubomir V. Kolev. Determining the positive definiteness margin of interval matrices. Reliable Computing = Nadezhnye vychisleniia, 13(6):445– 466, December 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-007-9046-5;   http://link.springer.com/article/10.1007/
s11155-007-9046-5/;   http://www.springerlink.com/openurl.asp?

genre=article&issn=1385-3139&volume=13&issue=6&spage=445;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=6&spage=445-466.

Xiang:2007:CPV

[700] Gang Xiang, Martine Ceberio, and Vladik Kreinovich. Computing population variance and entropy under interval uncertainty: Linear-time algorithms. Reliable Computing = Nadezhnye vychisleniia, 13(6):467–488, December 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11155-
007-9045-6;   http://link.springer.com/article/10.1007/s11155-
007-9045-6/;   http://www.springerlink.com/openurl.asp?genre=
article&issn=1385-3139&volume=13&issue=6&spage=467;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=6&spage=467-488.

Chabert:2007:CPI

[701] Gilles Chabert and Luc Jaulin. Computing the pessimism of inclusion functions. Reliable Computing = Nadezhnye vychisleniia, 13(6):489–504, December 2007. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-

1340 (electronic). URL http://link.springer.com/article/10.1007/
s11155-007-9044-7;   http://link.springer.com/article/10.1007/
s11155-007-9044-7/;   http://www.springerlink.com/openurl.asp?
genre=article&issn=1385-3139&volume=13&issue=6&spage=489;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=1385-3139&
volume=13&issue=6&spage=489-504.

Goldsztejn:2010:IAR

[702] A. Goldsztejn and L. Jaulin. Inner approximation of the range of vectorvalued functions. Reliable Computing = Nadezhnye vychisleniia, 14(1): 1–23, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Miyajima:2010:FVA

[703] S. Miyajima, T. Ogita, S. M. Rump, and S. Oishi. Fast verification for all eigenpairs in symmetric positive definite generalized eigenvalue problems. Reliable Computing = Nadezhnye vychisleniia, 14(1):24–45, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Patre:2010:RSF

[704] B. M. Patre and P. J. Deore. Robust state feedback for interval systems: An interval analysis approach. Reliable Computing = Nadezhnye vychisleniia, 14(1):46–60, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Garloff:2010:KN

[705] J. Garloff. Karl L. E. Nickel (1924–2009). Reliable Computing = Nadezhnye vychisleniia, 14(1):61–65, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Beelitz:2010:CCD

[706] T. Beelitz, B. Lang, P. Ueberholz, and P. Willems. Closing the case t = 3 for 3-D spherical t-designs using a result-verifying nonlinear solver. Reliable Computing = Nadezhnye vychisleniia, 14(1):66–77, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Hladik:2010:SSC

[707] M. Hladik. Solution sets of complex linear interval systems of equations. Reliable Computing = Nadezhnye vychisleniia, 14(1):78–87, 2010. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Meslem:2010:GPS

[708] N. Meslem, N. Ramdani, and Y. Candau. Guaranteed parameter set estimation for monotone dynamical systems using hybrid automata. Reliable Computing = Nadezhnye vychisleniia, 14(1):88–104, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Corsaro:2010:AAI

[709] S. Corsaro and M. Marino. Archetypal analysis of interval data. Reliable Computing = Nadezhnye vychisleniia, 14(1):105–116, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Ray:2010:NSS

[710] S. Ray and P. S. V. Nataraj. A new strategy for selecting subdivision point in the Bernstein approach to polynomial optimization. Reliable Computing = Nadezhnye vychisleniia, 14(1):117–137, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Zimmer:2010:SMV

[711] M. Zimmer, W. Kraemer, and W. Hofschuster. Sparse matrices and vectors in C-XSC. Reliable Computing = Nadezhnye vychisleniia, 14(1):138– 160, 2010. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Alt:2011:ASL

[712] R. Alt, J.-L. Lamotte, and S. Markov. On the accuracy of the solution of linear problems on the CELL processor. Reliable Computing = Nadezhnye vychisleniia, 15(1):1–12, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Argaez:2011:SOS

[713] M. Argaez. Solving overdetermined systems in l^{p} quasi-norms. Reliable Computing = Nadezhnye vychisleniia, 15(1):13–25, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Blomquist:2011:SCC

[714] F. Blomquist. Staggered correction computations with enhanced accuracy and extremely wide exponent range. Reliable Computing = Nadezhnye vychisleniia, 15(1):26–35, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Bohlender:2011:DAO

[715] G. Bohlender and U. Kulisch. Definition of the arithmetic operations and comparison relations for an interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 15(1):36–42, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Cheu:2011:TPO

[716] D. Cheu and L. Longpre. Towards the possibility of objective interval uncertainty in physics. Reliable Computing = Nadezhnye vychisleniia, 15 (1):43–46, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Goldsztejn:2011:CCP

[717] A. Goldsztejn, Y. Lebbah, C. Michel, and M. Rueher. Capabilities of constraint programming in safe global optimization. Reliable Computing = Nadezhnye vychisleniia, 15(1):47–59, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Villaverde:2011:EVU

[718] K. Villaverde and G. Xiang. Estimating variance under interval and fuzzy uncertainty: Parallel algorithms. Reliable Computing = Nadezhnye vychisleniia, 15(1):60–68, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Hattangady:2011:BFP

[719] S. Hattangady, W. Edmonson, and W. Alexander. Block floating point interval ALU for digital signal processing. Reliable Computing = Nadezhnye vychisleniia, 15(2):69–80, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Petkovic:2011:CAI

[720] I. Petković. Computational aspects of the implementation of disk inversions. Reliable Computing = Nadezhnye vychisleniia, 15(2):81–90, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Petkovic:2011:HOM

[721] M. S. Petković and D. M. Milosevic. Higher order methods for the inclusion of multiple zeros of polynomials. Reliable Computing = Nadezhnye vychisleniia, 15(2):91–108, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Nataraj:2011:INM

[722] P. S. V. Nataraj and M. Arounassalame. An interval Newton method based on the Bernstein form for bounding the zeros of polynomial systems. Reliable Computing = Nadezhnye vychisleniia, 15(2):109–119, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Diep:2011:SCS

[723] Nguyen Hong Diep and N. Revol. Solving and certifying the solution of a linear system. Reliable Computing = Nadezhnye vychisleniia, 15(2): 120–131, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Pownuk:2011:AOP

[724] A. Pownuk and N. K. G. Ramunigari. Application of order-preserving functions to the modeling of computational mechanics problems with uncertainty. Reliable Computing = Nadezhnye vychisleniia, 15(2):132–143, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

vanEmden:2011:IAI

[725] M. H. van Emden. From interval arithmetic to interval constraints. Reliable Computing = Nadezhnye vychisleniia, 15(2):144–155, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Yamanaka:2011:NVA

[726] N. Yamanaka, M. Kashiwagi, S. Oishi, and T. Ogita. A note on a verified automatic integration algorithm. Reliable Computing = Nadezhnye vychisleniia, 15(2):156–167, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Enszer:2011:VSP

[727] J. A. Enszer and M. A. Stadtherr. Verified solution and propagation of uncertainty in physiological models. Reliable Computing = Nadezhnye vychisleniia, 15(3):168–178, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Kempken:2011:VFM

[728] S. Kempken and W. Luther. Verifed factorization methods for SMP/G/1 queueing systems and their interplay in an integrated problem-solving environment. Reliable Computing = Nadezhnye vychisleniia, 15(3):179–192, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Kolberg:2011:EPS

[729] M. Kolberg, W. Kraemer, and M. Zimmer. Efficient parallel solvers for large dense systems of linear interval equations. Reliable Computing = Nadezhnye vychisleniia, 15(3):193–206, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Kubica:2011:IMS

[730] B. J. Kubica. Interval methods for solving underdetermined nonlinear systems. Reliable Computing = Nadezhnye vychisleniia, 15(3):207–217, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Magoc:2011:UPC

[731] T. Magoc, M. Ceberio, and F. Modave. Using preference constraints to solve multi-criteria decision making problems. Reliable Computing = Nadezhnye vychisleniia, 15(3):218–229, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Argaez:2011:HAG

[732] M. Argaez, L. Velázquez, C. Quintero, H. Klie, and M. Wheeler. A hybrid algorithm for global optimization problems. Reliable Computing = Nadezhnye vychisleniia, 15(3):230–241, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Roy:2011:GOS

[733] J. Roy and R. B. Kearfott. Global optimization and singular nonlinear programs: New techniques. Reliable Computing = Nadezhnye vychisleniia, 15(3):242–250, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Nataraj:2011:ETM

[734] P. S. V. Nataraj and S. Sondur. The extrapolated Taylor model. Reliable Computing = Nadezhnye vychisleniia, 15(3):251–278, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Tischler:2011:SDP

[735] G. Tischler and Jürgen Wolff von Gudenberg. Solving decidability problems with interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 15(3):279–289, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Kapitanski:2011:DRC

[736] L. Kapitanski and S. Zivanovic. Dynamics with a range of choice. Reliable Computing = Nadezhnye vychisleniia, 15(4):290–299, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Magoc:2011:AFM

[737] T. Magoc, Xiaojing Wang, F. Modave, and M. Ceberio. Applications of fuzzy measures and intervals in finance. Reliable Computing = Nadezhnye vychisleniia, 15(4):300–311, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Nehmeier:2011:FET

[738] M. Nehmeier and Jürgen Wolff von Gudenberg. filib++, expression templates and the coming interval standard. Reliable Computing = Nadezhnye vychisleniia, 15(4):312–320, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Rauh:2011:DRO

[739] A. Rauh, E. Auer, M. Freihold, E. P. Hofer, and H. Aschemann. Detection and reduction of overestimation in guaranteed simulations of Hamiltonian systems. Reliable Computing = Nadezhnye vychisleniia, 15(4):321–332, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Rauh:2011:ROC

[740] A. Rauh, J. Minisini, E. P. Hofer, and H. Aschemann. Robust and optimal control of uncertain dynamical systems with state-dependent switchings using interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 15(4):333–344, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Sharaya:2011:TSS

[741] I. A. Sharaya and S. P. Shary. Tolerable solution set for interval linear systems with constraints on coefficients. Reliable Computing = Nadezhnye vychisleniia, 15(4):345–357, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Shary:2011:NIL

[742] S. P. Shary. On nonnegative interval linear systems and their solution. Reliable Computing = Nadezhnye vychisleniia, 15(4):358–369, 2011. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Rauh:2011:VSO

[743] A. Rauh and E. Auer. Verified simulation of ODEs and DAEs in ValEncIA-IVP. Reliable Computing = Nadezhnye vychisleniia, 15(4):370–381, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Kolev:2011:MDR

[744] Lubomir V. Kolev. A method for determining the regularity radius of interval matrices. Reliable Computing = Nadezhnye vychisleniia, 16(1): 1–26, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Hladik:2011:ESM

[745] Milan Hladik and Luc Jaulin. An eigenvalue symmetric matrix contractor. Reliable Computing = Nadezhnye vychisleniia, 16(1):27–37, 2011. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Karmakar:2012:CSD

[746] Samiran Karmakar and A. K. Bhunia. A comparative study of different order relations of intervals. Reliable Computing = Nadezhnye vychisleniia, 16:38–72, January 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Just:2012:SGS

[747] E. Just and B. Lang. A success-guided selection of expanded systems for result-verifying nonlinear solvers. Reliable Computing = Nadezhnye vychisleniia, 16:73–83, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Schodl:2012:CNM

[748] P. Schodl and A. Neumaier. Continuity notions for multi-valued mappings with possibly disconnected images. Reliable Computing = Nadezhnye vychisleniia, 16:84–101, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Miyajima:2012:CEE

[749] S. Miyajima. Componentwise error estimates for solutions obtained by stationary iterative methods. Reliable Computing = Nadezhnye vychisleniia, 16:102–106, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Miyajima:2012:RBT

[750] S. Miyajima. The relation between two types of error bounds for computed matrix eigenvalues. Reliable Computing = Nadezhnye vychisleniia, 16: 107–113, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Kiel:2012:YYA

[751] S. Kiel. YalAA: Yet another library for affine arithmetic. Reliable Computing = Nadezhnye vychisleniia, 16:114–129, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Goldsztejn:2012:MIRa

[752] A. Goldsztejn. Modal intervals revisited, Part 1: a generalized interval natural extension. Reliable Computing = Nadezhnye vychisleniia, 16: 130–183, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Goldsztejn:2012:MIRb

[753] A. Goldsztejn. Modal intervals revisited, Part 2: a generalized interval mean value extension. Reliable Computing = Nadezhnye vychisleniia, 16: 184–209, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Mayer:2012:EMR

[754] G. Mayer. On an expression for the midpoint and the radius of the product of two intervals. Reliable Computing = Nadezhnye vychisleniia, 16: 210–224, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Milosevic:2012:OLM

[755] M. R. Milosevic and M. S. Petković. Ostrowski-like method for the inclusion of a single complex polynomial zero. Reliable Computing = Nadezhnye vychisleniia, 16:225–238, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Herrero:2012:EIS

[756] P. Herrero, P. Georgiou, C. Toumazou, B. Delaunay, and L. Jaulin. An efficient implementation of the SIVIA algorithm in a high-level numerical programming language. Reliable Computing = Nadezhnye vychisleniia, 16:239–251, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Harlow:2012:MRP

[757] J. Harlow, R. Sainudiin, and W. Tucker. Mapped regular pavings. Reliable Computing = Nadezhnye vychisleniia, 16:252–282, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

deWeerdt:2012:PIF

[758] E. de Weerdt, E. van Kampen, Q. P. Chu, and J. A. Mulder. Polynomial inclusion functions. Reliable Computing = Nadezhnye vychisleniia, 16: 283–307, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Wang:2012:NUB

[759] C. Wang. A new uncertainty-bearing floating-point arithmetic. Reliable Computing = Nadezhnye vychisleniia, 16:308–361, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Delgado:2012:OBR

[760] J. Delgado and J. M. Pena. Optimality of Bernstein representations for computational purposes. Reliable Computing = Nadezhnye vychisleniia, 17(1):1–10, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Leroy:2012:CUS

[761] R. Leroy. Convergence under subdivision and complexity of polynomial minimization in the simplicial Bernstein basis. Reliable Computing = Nadezhnye vychisleniia, 17(1):11–21, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Michelucci:2012:CBC

[762] D. Michelucci, S. Foufou, and A. Kubicki. On the complexity of the Bernstein combinatorial problem. Reliable Computing = Nadezhnye vychisleniia, 17(1):22–33, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Narkawicz:2012:BRR

[763] A. Narkawicz, J. Garloff, A. P. Smith, and C. A. Muñoz. Bounding the range of a rational function over a box. Reliable Computing = Nadezhnye vychisleniia, 17(1):34–39, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Ray:2012:MME

[764] S. Ray and P. S. V. Nataraj. A matrix method for efficient computation of Bernstein coefficients. Reliable Computing = Nadezhnye vychisleniia, 17 (1):40–71, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Bela:2012:RRA

[765] S. Béla and B. Jüttler. Real root approximation using fat spheres. Reliable Computing = Nadezhnye vychisleniia, 17(2):72–96, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Crespo:2012:UFU

[766] L. G. Crespo, D. P. Giesy, and S. P. Kenny. A unifying framework to uncertainty quantification of polynomial systems subject to aleatory and epistemic uncertainty. Reliable Computing = Nadezhnye vychisleniia, 17 (2):97–127, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Dang:2012:RAP

[767] T. Dang and R. Testylier. Reachability analysis for polynomial dynamical systems using the Bernstein expansion. Reliable Computing = Nadezhnye vychisleniia, 17(2):128–152, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

deVisser:2012:INA

[768] C. C. de Visser, E.-J. van Kampen, Q. P. Chu, and J. A. Mulder. Intersplines: a new approach to globally optimal multivariate splines using interval analysis. Reliable Computing = Nadezhnye vychisleniia, 17(2): 153–191, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Foufou:2012:BBA

[769] S. Foufou and D. Michelucci. The Bernstein basis and its applications in solving geometric constraint systems. Reliable Computing = Nadezhnye vychisleniia, 17(2):192–208, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Narkawicz:2012:FVC

[770] A. Narkawicz and C. A. Muñoz. Formal verification of conflict detection algorithms for arbitrary trajectories. Reliable Computing = Nadezhnye vychisleniia, 17(2):209–237, 2012. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic).

Dobronets:1990:TSN

[771] B. S. Dobronets and V. V. Shaydurov. Two-sided Numerical Methods. Nauka (Siberian Department), Novosibirsk, Russia, 1990. ISBN ???? 208 pp. LCCN ????

Ludyk:1990:CDS

[772] Günter Ludyk. CAE von Dynamischen Systemen. Analyse, Simulation, Entwurf von Regelungssystemen. (German) [CAE of Dynamic Systems. Analysis, Simulation, Design of control systems]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-51676-X, 0-387-51676-X. xi + 335 pp. LCCN ???? URL http://www.

zentralblatt-math.org/zmath/en/search/?an=0808.68020.

Neumaier:1990:IMS

[773] Arnold Neumaier. Interval Methods for Systems of Equations, volume 37 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-33196-X. xvi + 255 pp. LCCN QA297.75 .N49 1990.

Nesterov:1991:IC

[774] Vyacheslav M. Nesterov, editor. Interval Computations, 1991. ISSN 1385-3139 (print), 1573-1340 (electronic). URL mailto:nest@nit.spb.su. Box 52, St.Petersburg 195256, Russia. A refereed international journal, Interval Computations is the only periodical in the world devoted specifically to various aspects of reliable numerical computations based on the interval approach. It is managed by an international editorial board from Bulgaria, Germany, Japan, Russia, and the United States, and printed in Russia. The journal includes various items in the fields of theoretical research, computer tools, applications, interdisciplinary research and other relevant areas. The IC has been prepared with TEX since 1992. Recently a new style has been designed based on standard L^{A}TEX article style with changes in fonts, headings layout, running heads etc. It uses 10.8pt Baskerville for body text (the unusual font size is selected for better wedding with CM math) and horizontally-squeezed Palatino for headings. For mathematical formulae, standard Computer Modern is preserved. Proofs are made using emTEX DVIHPLJ on a Hewlett–Packard LaserJet IIIp, and final pages are printed at 600dpi. Using VTEX is now under consideration. Papers are accepted in any TEX variety; L^{A}TEX article.sty preferred.

Anonymous:1994:S

[775] Anonymous. SCAN-93. Moscow Institute for New Technologies in Education, Moscow, 1994. ISSN 1385-3139 (print), 1573-1340 (electronic). 1–133 pp. Papers from the IMACS/GAMM International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics held in Vienna, September 26–29, 1993, Interval Comput./Interval. Vychisl. 1994, no. 4.

Aberth:1998:PNM

[776] Oliver Aberth. Precise numerical methods using C++. Academic Press, New York, NY, USA, 1998. ISBN 0-12-041750-2. xiv + 238 pp. LCCN QA76.73.C153 A32 1998. URL http://

www.loc.gov/catdir/description/els033/98117803.html;   http:
//www.loc.gov/catdir/toc/els032/98117803.html.

Kreinovich:1998:CCF

[777] Vladik Kreinovich, Anatoly Lakeyev, Jiři Rohn, and Patrick Kahl. Computational Complexity and Feasibility of Data Processing and Interval Computations, volume 10 of Applied optimization. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-4865-6 (hardcover). xii + 459 pp. LCCN QA267.7 .C68 1998.

URL http://www.loc.gov/catdir/enhancements/fy0814/97043085-
d.html;   http://www.loc.gov/catdir/enhancements/fy0814/97043085-
t.html.

Fortnow:1999:BRB

[778] Lance Fortnow. Book review: Bounded Queries in Recursion Theory, by William A. Gasarch and Georgia A. Martin (Birkhäuser. Boston, Basel, Berlin, 1999). ACM SIGACT News, 30(3):13–15, September 1999. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [779].

Gasarch:1999:BQR

[779] William I. Gasarch and Georgia A. Martin. Bounded Queries in Recursion Theory, volume 16 of Progress in computer science and applied logic. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-3966-7 (hardcover). xiii + 353 pp. LCCN QA267.7 .G38 1999.

Bauch:1987:ITA

[780] Hartmut Bauch, K. U. Jahn, D. Oelschägel, H. Süsse, and V. Wiebigke, editors. Intervallmathematik — Theorie Und Anwendungen. Number 72 in Mathematisch-naturwissenschaftliche Bibliothek. Bsb B. G. Teubner Verlagsges, Leipzig, East Germany, 1987. ISBN 3-322-00384-1. ISSN 0465-3769. 260 pp. LCCN QA297.75 .I6 1987. DM28.00.

Voshchinin:1992:IIC

[781] A. P. Voshchinin and J. W. von Gudenberg, editors. Interval ’92: International conference on interval and stochastic methods in science and engineering — September 22–25, 1992, Moscow, number 3–4 in Interval Computations (6). Moscow Institute for New Technologies in Education, Moscow, Russia, 1992. ISSN 0135-4868. Interval Comput./Interval. Vychisl. 1992, no. 3; Interval Comput./Interval. Vychisl. 1992, no. 4.

Corliss:1993:AIC

[782] G. F. Corliss and R. B. Kearfott, editors. Abstracts for an International Conference on Numerical Analysis with Automatic Result Verification: Mathematics, Application and Software, February 25–March 1, 1993, Lafayette, LA, 1993, volume 3(3–4) of Interval Computations = Interval’nye vychisleniia. ????, ????, 1993. ISBN ???? ISSN 0135-4868. LCCN ????

Nesterov:1993:PIC

[783] V. M. Nesterov, editor. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification. Moscow Institute for New Technologies in Education, Moscow, Russia, 1993. ISSN 0135-4868. Held in Lafayette, Louisiana, February 25–March 1, 1993, Interval Comput. /Interval. Vychisl. 1993.

Kearfott:1994:NAA

[784] R. B. Kearfott and G. F. Corliss, editors. Numerical analysis with automatic result verification: International conference — February 1993, Lafayette, LA, number 3 (or 4??) in INTERVAL COMPUTATIONS 1993. Institute for new technologies, ????, 1994. ISSN 0135-4868.